### v1

# Enhancing variational quantum algorithms by balancing training on classical and quantum hardware

Abstract
Quantum computers offer a promising route to tackling problems that are classically intractable such as in prime-factorization, solving large-scale linear algebra and simulating complex quantum systems, but require fault-tolerant quantum hardware. On the other hand, variational quantum algorithms (VQAs) have the potential to provide a near-term route to quantum utility or advantage, and is usually constructed by using parametrized quantum circuits (PQCs) in combination with a classical optimizer for training. Although VQAs have been proposed for a multitude of tasks such as ground-state estimation, combinatorial optimization and unitary compilation, there remain major challenges in its trainability and resource costs on quantum hardware. Here we address these challenges by adopting Hardware Efficient and dynamical LIe algebra Supported Ansatz (HELIA), and propose two training schemes that combine an existing $\mathfrak{g}$-sim method (that uses the underlying group structure of the operators) and the Parameter-Shift Rule (PSR). Our improvement comes from distributing the resources required for gradient estimation and training to both classical and quantum hardware. We numerically test our proposal for ground-state estimation using Variational Quantum Eigensolver (VQE) and classification of quantum phases using quantum neural networks. Our methods show better accuracy and success of trials, and also need fewer calls to the quantum hardware on an average than using only PSR (upto 60% reduction), that runs exclusively on quantum hardware. We also numerically demonstrate the capability of HELIA in mitigating barren plateaus, paving the way for training large-scale quantum models.
I Introduction
Quantum computing holds great promise for tackling problems that are intractable for classical computers or would take them years to solve, such as simulating natural systems [1], prime factorization [2], solving linear equations [3], machine learning tasks [4, 5, 6], optimization [7] and quantum chemistry [8, 9]. Despite recent breakthroughs in implementing quantum error correction [10], it may still take many years [11, 12, 13] to develop fault-tolerant quantum hardware. Current quantum devices face significant challenges, including number of qubits, qubit coherence time and gates with limited fidelity [10, 14, 15, 16, 17, 18, 19].
Limitations of current quantum computing hardware has spurred focus towards variational quantum algorithms (VQA). These are hybrid quantum-classical approaches that utilize parameterized quantum circuits (PQC) to address challenges across diverse fields, including machine learning, optimization, and ground-state energy calculations. VQAs are especially suited for noisy intermediate-scale quantum (NISQ) devices, which have a limited qubit count and noisy (unprotected) operations. The core idea is to train quantum circuits with tunable parameters and optimizing these parameters using a classical optimizer for an objective function related to the problem being solved.
In the training process of VQAs, the parameters of PQC are updated via either gradient-free or gradient-based optimization techniques. The gradient-free methods [20, 21, 22, 23, 24, 25] are preferred when calculation of gradient is challenging or the loss function is not differentiable, but their performance might not scale well with problem size and system noise [26, 27]. In contrast, the literature on gradient-based training of quantum machine learning (QML) primarily emphasizes the efficient computation of gradients of parameters of trainable quantum circuits on quantum hardware. The most commonly employed techniques include the parameter-shift rule (PSR) [28] and its various generalization or extensions [29, 30]. PSR requires running two or more quantum circuit evaluations with shifted parameter values, for each trainable parameter in the circuit, and it run exclusively on quantum hardware.
Classical simulation methods can be an effective alternative for advancing quantum technologies by emulating quantum systems using classical resources. Goh et al. proposed an efficient circuit simulation and gradient computation method called $\mathfrak{g}$-sim , reducing the resource burden leveraging the Lie-algebraic structure of variational quantum circuits [31]. Clifford Perturbation theory has been recently put forward to approximate operator expectation values in near-Clifford circuits [32]. These techniques are crucial for adapting classical backpropagation methods in a quantum context [33].
Optimization in quantum computing is often challenging due to several factors. A major hurdle being the barren plateau (BP) phenomenon [34], where gradients vanish exponentially as qubit count increases  , making it difficult to find numerically meaningful updates for large circuits. Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization. Several BP-free models have been proposed in literature like quantum convolutional neural networks [35, 36] and permutation-equivariant quantum neural networks [37]. Recent research also points towards the possibility that all BP-free quantum models might indeed be efficiently classically simulable [38], raising doubts about achieving any quantum advantage in the future.
Classical algorithms for simulating general quantum circuits do exist, but they suffer from exponential computational and memory overhead [39]. Over the years, several methods have been developed for accurate and efficient classical simulation of certain quantum circuits. These methods often approximate the circuits in specific regimes or restrict the types of gates and measurements allowed, leveraging special structures in the problems. These include MPS-based tensor networks methods that can simulate shallow noisy circuits for hundreds of qubits with limited gate fidelity [40]. $\mathfrak{g}$-sim is another method that offers an efficient simulation algorithm. This method is effective when both the generators of the ansatz and the measurement operator are within a dynamical Lie algebra (DLA) [31], whose dimension scales polynomially with the number of qubits. We will explain more about the $\mathfrak{g}$-sim method in a later section as it is central to our ansatz construction.
Recent advancements in classical simulation methods have significantly enhanced our ability to estimate and optimize quantum circuits, even in the presence of noise. Fontana et.al proposed low weight efficient simulation algorithm (LOWESA) to classically estimate expectation values of parametrized quantum circuits in noisy hardware [41]. The authors recreate a classical surrogate of cost landscape in the presence of device noise by ignoring Pauli terms beyond a certain frequency threshold. Rudolph et.al extended this to noiseless quantum simulation and classically simulated the 127-qubit quantum utility experiment [42, 43] while Angrisani et. al further demonstrated provable guarantees for most noiseless quantum circuits [44]. Recently, Lerch et.al showed it is always possible to identify patches for which similar classical surrogate can be generated [45]. These techniques should be accounted for when constructing variational models with potential for quantum utility.
The optimization landscape in quantum circuits is also highly non-convex, with numerous local minima and saddle points that can trap algorithms and complicate the search for global minima [46, 47, 48, 49]. Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization. Even for models that are trainable, computing gradients for quantum circuits is often quantum resource-intensive as methods such as the parameter-shift rule or finite differences necessitate multiple circuit executions, which scale linearly [33] with the number of parameters and shots. These factors, coupled with hardware limitations such as qubit connectivity and decoherence [10, 14, 15, 16, 17, 18, 19], make effective gradient-based optimization in quantum computing a complex task, frequently driving the need for alternative or hybrid methods.
To mitigate the high quantum resource costs associated with training VQAs, we propose a hybrid method that integrates $\mathfrak{g}$-sim and the PSR. The gradient estimation task is delegated to both classical and quantum hardware. This approach aims to reduce the number of Quantum Processing Unit (QPU) calls, thereby saving resources. Our research focuses on training quantum models, emphasizing the use of specific ansätze and customized training schemes to enhance the training process and improve accuracy in tasks like ground state estimation using VQE and classification using quantum neural networks (QNNs). While this study does not evaluate the practical utility of these models or their performance relative to classical counterparts, it aims to explore the potential of scalable and accurate trainable quantum models. Our hope is that with increase in the size and accuracy of the trainable models we can have enough overlap with useful quantum models and move towards quantum utility.
We will expand on the main results in the coming sections. The paper is divided as follows. We introduce the relevant background literature in Sec. II. We then discuss our contributions in terms of the choice of quantum circuits and two training schemes namely Alternate and Simultaneous in Sec. III. In Sec. IV we demonstrate the improvement in various metrics through extensive numerics. Finally, we end with the Conclusion and Outlook in Sec. V.
II Overview of Literature
II.1 Variational Quantum Algorithms
VQAs constitute a fundamental framework within hybrid quantum-classical computation, providing a viable approach to exploit the capabilities of quantum systems in the NISQ era. By integrating PQCs with classical optimization techniques, VQAs are specifically tailored to tackle computationally demanding problems across domains such as quantum chemistry, combinatorial optimization, and machine learning [50, 51, 52, 53, 54, 55].
VQAs involves an $n$-qubit state $\rho$ (usually encoding problem data), acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ which undergoes evolution through a parameterized quantum unitary. The unitary transformation is expressed as a sequence of $L$ parameterised unitary gates
$U(\boldsymbol{\theta})=\prod_{i=1}^{L}U_{i}(\theta_{i})W_{i}$ (1)
where $\boldsymbol{\theta}=[\theta_{1},\theta_{2},\ldots,\theta_{L}]$ denotes a set of trainable real-valued parameters and $W_{i}$ are non-trainable gates. A slightly more general scenario can be when each $U_{i}(\theta_{i})$ is composed of multiple trainable parametrized unitaries itself. However, we ignore that for ease of notation, and the same ideas can be extended to the general unitaries as well.
The expectation value of a Hermitian observable $\hat{O}$ is subsequently measured on the evolved quantum state, using a quantum device, given as
$l_{\boldsymbol{\theta}}(\rho,\hat{O})=Tr[U^{\dagger}(\boldsymbol{\theta)}\rho U(\boldsymbol{\theta})\hat{O}].$ (2)
Classical optimizers are utilized to perform the optimization task
$\underset{\boldsymbol{\theta}}{argmin}\>\>{F(l_{\boldsymbol{\theta}}(\rho,\hat{O}))},$ (3)
where we minimize some loss function $F$ of the expectation value depending on the task. For example, when finding ground state of a Hamiltonian $H$ using VQE, the loss function is the expectation value of $H$ on the trainable quantum circuit. For classification problem, the loss function can be the mean squared error (MSE) between predicted and true labels. The algorithm iteratively optimizes a cost function using the output of a quantum circuit, enabling the exploration of complex solution spaces that are potentially intractable for classical approaches.
However, despite their promise, VQA face several critical challenges, including barren plateaus in the optimization landscape, limited scalability with increasing problem size, and the detrimental effects of noise and decoherence on circuit fidelity [34, 50, 56]. Overcoming these obstacles necessitates advancements in algorithmic design, optimization techniques, and the development of robust quantum hardware.
Algorithms considered in this work
In this study, we focus on two applications of VQAs, illustrating their potential to address complex quantum problems. The first application investigates the Variational Quantum Eigensolver (VQE), a hybrid quantum-classical algorithm widely used for determining the ground state of quantum systems and calculating their corresponding energies [57]. VQE has become a cornerstone in quantum chemistry and materials science, enabling the efficient simulation of molecular systems and facilitating the discovery of new materials. Recent research has also incorporated VQE into Quantum Error Detection pipeline [58, 59].
The second application delves into quantum phase classification, where VQAs are utilized to identify and distinguish different quantum phases of matter [35, 60, 61]. This application leverages the unique capabilities of quantum circuits to encode and process quantum states, enabling the precise detection of phase transitions and the systematic classification of quantum phases.
II.2 Existing PQCs
[…]
×L
Figure 1: Quantum circuit block for YZ linear ansatz. The parametrized $R_{y}$ and $R_{z}$ rotations are applied on each qubit followed by $CNOT$ between neighboring qubits in a linear fashion. The blocks are often repeated to increase entanglement and expressivity
The available options for PQC is quite vast and it is not a priori clear which circuit structure is ideal for the task at hand. We will consider three particular ansatz choices and motivate their use for our ansatz construction.
One often considered ansatz choice is made based on qubit connectivity, to reduce circuit depth, by using the class of Hardware Efficient Ansatz (HEA) [62]. Blocks of single-qubit rotations followed by entangling two-qubit gates are repeated for a chosen number of layers. An example within this class is $YZ$ linear ansatz shown in Fig.1, where each block consists of a $Y$-rotation and a $Z$-rotation applied on each qubit, followed by CNOT gates between neighbouring qubits in a linear fashion.
Commonly selected problem-motivated ansatz include the Unitary Coupled Cluster (UCC) [63], which is widely utilized for investigating ground states of fermionic molecular Hamiltonians, and the Quantum Alternating Operator Ansatz (QAOA) [64, 65, 66], often applied to combinatorial optimization problems.
Another interesting problem-motivated example is the Hamiltonian Variational Ansatz (HVA) [67]. Given the task of finding ground state of a Hamiltonian $H=\sum_{i}H_{i}$, the ansatz is composed of unitary blocks of the form
$U(\boldsymbol{\theta})=\prod_{k}e^{-i\theta_{k}H_{k}}.$ (4)
Building on the insights from [31], this approach can be extended to include terms from the DLA of $\{H_{k}\}_{H_{k}\in H}$, provided the DLA has a dimension that scales polynomially with the number of qubits.
In our examples, we will combine the HVA ansatz with the HEA ansatz. For a detailed overview of the different ansatz options explored in the literature, refer [50].
After selecting an appropriate ansatz based on the architecture and the task, the subsequent step involves determining the optimal parameters of the ansatz to address the problem being solved. This process, referred to as the training method, is elaborated upon in the following section.
II.3 Training Methods
The training process aims to identify the optimal parameters of the selected ansatz for a given architecture. This is an iterative procedure, with each iteration comprising a loss function evaluation followed by a parameter update step. Although several gradient-free methods exist for parameter updation [20, 21, 22, 23], our primary focus here is on gradient-based techniques which require efficient and accurate algorithms to compute the partial derivatives of the cost function for effective training of variational quantum models.
Contemporary classical machine learning models, which often contain billions of parameters, efficiently compute gradients using the backpropagation algorithm [68]. This method requires only a single forward and backward pass to compute the gradients for all parameters simultaneously. Since classical backpropagation cannot be directly applied to quantum circuits [33], the Parameter-Shift Rule [28] is typically employed to compute circuit gradients. Alternatively, when the quantum circuit can be efficiently simulated classically, automatic differentiation frameworks, such as PyTorch [69] and Tensorflow [70], can be utilized for gradient evaluation. A notable case within this second scenario is $\mathfrak{g}$-sim , where placing certain restriction on quantum gates and measurement operators enables efficient classical simulation of the circuit. These two methods are discussed in detail in the following sections.
II.3.1 Parameter-shift Rule
Parameter-shift Rule (PSR) is commonly used for accurately estimating gradients of a PQC. For a loss function of the form given in Eq.(2), where each unitary is defined as $U_{i}(\theta_{i})=e^{-i\theta_{i}G_{i}}$ for some hermitian operator $G_{i}$, the partial derivatives are given by
$\frac{\partial{\ell}}{\partial{\theta_{k}}}$ $=i\bra{0}{U}^{\dagger}_{\geq k}G_{k}({U}^{\dagger}_{<k}{\hat{O}}U_{<k})U_{\geq k}\ket{0}$
$-i\bra{0}U^{\dagger}_{\geq k}(U^{\dagger}_{<k}\hat{O}{U}_{<k})G_{k}{U}_{\geq k}\ket{0}$ (5)
where we defined $U_{<k}=U_{1}W_{1}U_{2}W_{2}\ldots U_{k-1}W_{k-1}$ and $U_{\geq k}=U_{k}W_{k}U_{k+1}W_{k+1}\ldots U_{n}W_{n}$. PSR can be employed to obtain a closed-form expression of Eq.(II.3.1) when the eigenspectrum of the generator $G_{k}$ either has only two unique eigenvalues or is symmetric and evenly spaced [71, 28, 72, 73, 30].
Considering the case where $G_{k}$ with only two unique eigenvalues ±$\pm r$ i.e., $G_{k}^{2}=r^{2}I$ the corresponding unitary can be rewritten as
$U_{k}(\theta)=\cos{r\theta}I-ir^{-1}G_{k}\sin{r\theta}$ ) (6)
Following [28], through some simple algebra the above equation can be turned into
$\frac{\partial{\ell}}{\partial{\theta_{k}}}$ $=\frac{r}{2}\Big{(}\langle\psi|(I-ir^{-1}G)^{\dagger}Q(I-ir^{-1}G)|\psi\rangle$
$\langle\psi|(I+ir^{-1}G)^{\dagger}Q(I+ir^{-1}G)|\psi\rangle\Big{)}$ (7)
where we have defined $|\psi\rangle=U_{\geq k}|0\rangle$ and $Q={U}^{\dagger}_{<k}{\hat{O}}U_{<k}$. Finally using Eq.(6) and substituting $\theta=\pi/4r$, the partial derivative simplifies to
$\frac{\partial{\ell}}{\partial{\theta_{k}}}=r\Big{(}\ell(\boldsymbol{\theta}+\frac{\pi}{4r}\hat{e}_{k})-\ell(\boldsymbol{\theta}-\frac{\pi}{4r}\hat{e}_{k})\Big{)},$ (8)
a difference between the loss functions evaluated on either side of the parameter with a specific eigenvalue-related shift.
By measuring the loss function at two shifted values of the parameter $\theta_{k}$, the partial derivative can be estimated accurately upto shot noise. For more general eigenspectrum with equispaced eigenvalues, the parameter shift-rule can be generalized as shown in [30, 73]. Markovich et.al. have further extended this procedure to irregeular eigenspectrum as well [74].
The computation of a gradient for a single parameter in a quantum circuit requires executing the circuit at least twice, apart from the measurement overhead to obtain expectation value accurately. In quantum machine learning (QML), backpropagation is not directly applicable because it relies on the ability to store or reuse information [33], which poses scalability challenges. This limitation is illustrated through a specific example. Assuming each circuit iteration takes 1.48 milliseconds [75], equivalent to the latest coherence time of superconducting qubits, one iteration for a circuit with a billion parameters would take approximately $2\times 10^{9}\times 1.48$ ms, or 296 million seconds (3 days , 10 hours and 19 minutes). This demonstrates the poor scalability of PSR for large-scale systems. For certain examples of PQC construction, the training may scale more favorably. One such example is proposed in Bowles et. al to address the scalability issue by introducing structured circuits consisting of blocks of commuting parametrized quantum gates [76]. However, the utility of such structured PQCs is still not fully understood.
II.3.2 $\mathfrak{g}$-sim Method
In this section, we briefly explain the details of $\mathfrak{g}$-sim method proposed by Goh et al. [31], that is fully classical in nature. $\mathfrak{g}$-sim method is based on the study of the Lie group and the associated Lie algebra $g$, which is generated by the parametrized quantum circuit. Note that the foundational techniques used in [31] for this gradient estimation method are well-established by Somma et al. [77], Goh et al. primary contributed to reframe these techniques in a modern context tailored to the quantum computing community [31].
Figure 2: Quantum circuit for $\mathfrak{g}$-sim algorithm corresponding to the unitary in Eq. (9)
Consider a problem in which an $n$-qubit state $\rho$, acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ is sent through a PQC $U(\boldsymbol{\theta})$ of the form
$U(\boldsymbol{\theta})=\prod_{\ell=1}^{L}\prod_{k=1}^{K}e^{i\theta_{\ell k}H_{k}},$ (9)
where $\boldsymbol{\theta}=[\theta_{11},\theta_{12},\ldots,\theta_{LK}]$ is a set of trainable real-valued parameters, and $\{H_{1},H_{2},\ldots,H_{K}\}$ are $K$ Hermitian operators that are the gate generators of the circuit, as shown in Fig 2.
The vector space spanned by all possible nested commutators of $\{iH_{1},\dots,iH_{K}\}$, obtained by repeatedly taking the commutator between all elements until no new linearly independent element emerge is known as the DLA. For the ansatz of the form in Eq. (9), this DLA, denoted as $\mathfrak{g}$, is used to describe the system’s structure.
The dynamical Lie group $\mathcal{G}$ of a circuit of the form Eq. (9) is defined as
$\mathcal{G}=\{e^{iA}:iA\in\mathfrak{g}\},$
that is, the dynamical Lie group is obtained via the exponentiation of DLA.
The significance of the DLA lies in the fact that all unitaries of the form in Eq. (9) belong to a Lie group $\mathcal{G}$. Specifically, for any $V\in\mathcal{G}$, there exists (at least) one choice of parameter values $\boldsymbol{\theta}$ such that for a sufficiently large, but finite, number of layers $L$, we have $U(\boldsymbol{\theta})=V$[78]. That is, the dynamical Lie group $\mathcal{G}$ determines all possible unitaries that can be implemented by circuits of the form in Eq. (9).
$\mathfrak{g}$-sim method is applicable only under two assumptions [31]:
  • either the measurement operator or the initial state of a variational quantum circuit belong to the DLA of the ansatz
  • the dimension of DLA scales polynomially with the number of qubits.
In order to understand, how $\mathfrak{g}$-sim method works, we need to know how the elements of the DLA $\mathfrak{g}$ transform under commutation with elements of the DLA $\mathfrak{g}$, and under conjugation with elements of the dynamical Lie group $\mathcal{G}$. As $\mathfrak{g}$ is closed under commutation, the commutator between any two elements in $\mathfrak{g}$ is an element of $\mathfrak{g}$ and can be decomposed as a linear combination of a Schmidt-orthonormal basis $\{iG_{\alpha}:\alpha\in\{1,2,\ldots,\text{dim}(g)\}\}$ of $\mathfrak{g}$. In particular,
[$\commutator{iG_{\alpha}}{iG_{\beta}}=\sum_{\gamma=1}^{\text{dim}(\mathfrak{g})}f_{\alpha\beta}^{\gamma}iG_{\gamma}$ (10)
where $f_{\alpha\beta}^{\gamma}=\text{Tr}[iG_{\gamma}[iG_{\alpha},iG_{\beta}]]\in\mathbb{R}$ is computed via the standard projection in vector space $\mathfrak{g}$. The coefficients $f_{\alpha\beta}^{\gamma}$ are called the structure constants of the DLA $\mathfrak{g}.$
Input: $T_{max},\alpha,H=\sum_{i}^{g_{max}}a_{i}G_{i},\linebreak U(\vec{\theta})=e^{-iG_{1}\theta_{1}}\ldots e^{-iG_{n}\theta_{n}}$
Output: Optimized $\vec{\theta^{*}}$
Initialize: $\vec{\theta}\sim Norm(0,1)$
Define: $C(\vec{\theta})=\langle 0^{\otimes q}|U^{\dagger}(\vec{\theta})HU(\vec{\theta})|0^{\otimes q}\rangle$
Set: $H^{n}=H;a^{n}_{k}=a_{k},H^{n}_{k}=a_{k}^{n}G_{k}$
begin
       $i=0$
       while $i\leq T_{max}$ do
             $m=n$
             while $m\geq 1$ do
                   $k=0$
                   for $k\leq g_{max}$ do
1                         $H_{k}^{m-1}=e^{iG_{m}\theta_{m}}(a_{k}^{m}G_{k})e^{-iG_{m}\theta_{m}}$
                         $\quad\quad=\sum_{j}b^{m-1}_{j}G_{j}$
                         $k\leftarrow k+1$
2
3                  $H^{m-1}=\sum_{i}^{g_{max}}H_{i}^{m-1}=\sum_{i}^{g_{max}}a_{i}^{m-1}G_{i}$
4
                  $m\leftarrow m-1$
            $l=0$
             while $l\leq g_{max}$ do
5                   $g_{l}=\langle 0^{\otimes q}|G_{l}|0^{\otimes q}\rangle$
                   $l\leftarrow l+1$
6            $C\leftarrow\sum_{i}^{g_{max}}a_{i}^{0}g_{i}$
7             Obtain gradients $\gradient_{\vec{\theta}}C$ classically using an automatic differentiation framework
8             $\vec{\theta}\leftarrow\vec{\theta}-\alpha\gradient_{\vec{\theta}}C$
9
10            $i\leftarrow i+1$
11
12
13
Algorithm 1 $\mathfrak{g}$-sim Implementation for VQE
Using the above DLA structure we demonstrate an implementation of the full $\mathfrak{g}$-sim sub-routine in action for VQE in Alg.1. We choose a Hamiltonian with a DLA of polynomial dimension. The ansatz is chosen as $U(\boldsymbol{\theta})=e^{-i\theta_{1}G_{1}}\ldots e^{-i\theta_{n}G_{n}}$, where the generators are from the same DLA, allowing us to efficiently use the DLA structure in our computation.
The crucial factor facilitating the use of DLA structure is the Baker-Campbell-Hausdorff (BCH) formula [79] in Step. 1, which is given by
[…]
For operators $G_{i},G_{j}$ belonging to a chosen DLA, the expression for $e^{i\theta G_{i}}G_{j}e^{-iG_{i}}$ consists of only nested commutators. Since the DLA is closed under commutation operation, all of the terms in the infinite series in Eq.(11) belong to the DLA itself. By choosing a DLA scaling polynomially in the number of qubits, we ensure the expansion only has polynomially many unique operators, whose co-efficients are functions of the structure constants $f_{\alpha\beta}^{\gamma}$ as shown in Eq.(10).
However, it might still be difficult to obtain a closed form expression for Eq.(11). Hence, we further simplify it to the scenario when all the operators are n-qubit Pauli strings. This gives a very simple expression
[…]
where $P_{i},P_{j}$ are n-qubit Pauli strings. We further make our numerics computationally efficient by leveraging the $2n\times 2n$ binary symplectic matrix representation of Pauli, rather than their $2^{n}\times 2^{n}$ unitary matrix representation (Section 2.1 of [80]). As a result, we can replace matrix multiplication by bit-wise addition, which is much faster.
Fast application of the BCH formula due to choosing Pauli strings as well a DLA of polynomial dimension allows to calculate the cost function efficiently with polynomial compute resources. The gradients can then be efficiently calculated using an automatic-differentiation framework such as Pytorch [69], Tensorflow [70].
The limitations of the PSR and the efficiency of $\mathfrak{g}$-sim provide motivation for combining them into a hybrid approach to achieve efficient gradient evaluation. Apart from scalability and resource challenges in gradient evaluation, the training of quantum models remains challenging, primarily due to the vanishing gradient phenomenon, also known as the BP problem, where gradients decay exponentially as the number of qubits increases, making larger models effectively untrainable. We expand on this point in the following subsection.
II.4 Barren Plateau
Optimization in quantum algorithms encounters significant challenges due to the intrinsic limitations imposed by quantum measurements. Classical optimization methods rely on precise loss evaluations or accurate gradient information, both of which are compromised by the finite precision and stochastic nature of quantum measurements.
Barren plateaus (BPs) in quantum variational algorithms are characterized by regions in the parameter space where the gradient of the loss function becomes exponentially small, severely impeding the efficient optimization of quantum circuits [81, 82, 83]. The presence of a BP landscape implies that, for most parameter configurations, an exponentially large number of measurement samples is required to reliably estimate the loss gradient, which is essential for effective optimization.
The Variational Quantum Algorithm (VQA) is believed to exhibit a BP when variations in the model parameters $\theta$ in the unitary operation $U(\theta)$ lead to only exponentially small changes in the loss function $l_{\theta}(\rho,\hat{O})$, or magnitude of gradients $\partial l_{\theta}(\rho,\hat{O})/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\theta$) [34, 82, 83]. Cerezo et al., in their investigation of cost function locality, established that the inclusion of global observables $(\hat{O})$ in the cost function results in the emergence of exponentially vanishing gradients, irrespective of the circuit’s shallow depth [84]. Additional factors contributing to this phenomenon include the degree of entanglement [85, 86] and the presence of noise [87]. An unified theoretical framework, employing group theoretical principles, covering above factors were presented in [82, 83].
Beyond theoretical insights, substantial efforts have been devoted to developing practical strategies for mitigating barren plateaus [88]. These include advanced initialization techniques, such as layer-wise training [89], Bayesian-inspired parameter settings [90], meta-learning for initializing parameters [91], initializing circuits with sequences of identity blocks [92], and transfer-learning-based initialization approaches [93]. Alternative optimization frameworks have also been introduced, including adaptive learning rate strategies [94] and novel line search methods designed to facilitate navigation through flat loss landscapes [95]. Mhiri et.al. provided an unified analysis on warm-starting approaches in BP-affected loss landscape [96]. Classical neural network-based methodologies have also been explored, such as generating parameters for parameterized quantum circuits (PQCs) using neural networks [97, 98].
Empirical studies, such as those conducted in Ref. [99], have validated the effectiveness of these approaches in reducing the prevalence of barren plateaus in practical applications, including quantum neural networks. Additionally, targeted strategies for specific ansatz designs, such as Hamiltonian variational ansatz and parameterized matchgate circuits, have demonstrated their potential in mitigating or entirely avoiding barren plateaus [100, 101].
However, recent research has shed light on the classical simulability of BP-free models, offering valuable insights into their underlying structure. Cerezo et al. highlights that BP-free models are classically simulable when classical data is gathered during an initial data acquisition phase from the quantum device. This conclusion stems from the fact that barren plateaus arise from the curse of dimensionality, with mitigation strategies effectively reducing the problem to subspaces that are classically tractable [38].
Finally, using classical processing to improve the results from quantum device is not unique to our work, and has been used in [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112]. In fact, a recent paper also deals with the idea of processing different parts of the quantum circuit using classical and quantum hardware with Clifford Perturbation Theory [113]. However, our approach of combining $\mathfrak{g}$-sim and the PSR in a hybrid iterative fashion offers a balanced method to address the barren plateau problem, improve resource efficiency, and maintain non-classical simulability. This makes it a promising area for further research and development in quantum computing.
III Our Contributions
Gathering insights from current literature, we find extensive challenges across all building blocks of a variational quantum algorithm trained on today’s hardware. In this paper, we address issues related to choice of PQC, resource-inefficiency of gradient evaluation and trainability of large models in regards to scaling the number of parameters and the concurrent barren plateau issues that appear. We go through each of the improvements in a sequential manner highlighting clear instances of where and how the improvements are obtained using detailed numerical analysis.
III.1 PQC Selection
Our choice of PQC is motivated from the standpoint of reducing quantum resources for training as well as avoiding Barren Plateaus. In general, simulating the full density matrix requires keeping track of all the $n$-qubit Pauli operators which scales as $14^{n}-$ 1. However, to effectively utilize $\mathfrak{g}$-sim without incurring exponential computational and memory overhead, it is essential to restrict the training ansatz and measurement operator associated with a QML task to a poly-DLA. This can be a challenge, for example in VQE if the ground state lies far outside the polynomial DLA(poly-DLA) being explored (as shown in Table 3 in Appendix A), or a QML task whose accuracy gets limited by the polynomial search space (as shown in Fig.1).
Figure 3: The HELIA comprises two blocks of quantum gates: the orange block or Block Q ($U_{q}$) consists of layers of HEA (non-DLA part), followed by the blue block or Block G ($U_{g}$), which includes multi-qubit Pauli rotation gates with generators forming the basis of a task-specific DLA. The final measurement operator also belongs to the same DLA
To elevate $\mathfrak{g}$-sim to a higher expressibility, we propose a hybrid ansatz consisting of a parameterized non-DLA block $U_{q}$, which cannot be classically simulated, followed by a parameterized $\mathfrak{g}$-sim block $U_{g}$ that can be classically simulated within the $\mathfrak{g}$-sim framework. A specific ansatz falling within this proposal is the Hardware Efficient and dynamical LIe algebra Supported Ansatz (HELIA), shown in Fig. 3, where $U_{q}$ represents the hardware efficient ansatz (HEA). $U_{g}$ is composed of multi-qubit Pauli rotations $\exp{-i\theta P}$ ), where the Paulis $P$ form an orthonormal basis of a chosen poly-DLA. Both HEA [62] and Lie algebra Supported Ansatz [83] are known in literature, however our proposal aims at combining them into a single ansatz, which we will show has several important advantages.
This specific ordering of gates is important for our protocol, as we require the gates and measurement operator chosen from the poly-DLA to be adjacent in the quantum circuit for $\mathfrak{g}$-sim to be used efficiently. If the order is reversed, the Heisenberg-evolved operator obtained after applying the HEA gates on the measurement operator will have terms from outside the poly-DLA, making $\mathfrak{g}$-sim quite ineffective. In our examples of VQE and Classification, we use a YZ linear ansatz, shown in Fig. 1 (which belongs to the HEA class), with varying layers, and the DLA is motivated by the problem being solved.
Figure 4: VQE performed for a 6-qubit TFIM Hamiltonian. Only using HEA using PSR (shown in orange) or multi-qubit Pauli rotation gates using $\mathfrak{g}$-sim (shown in blue) is not enough to reach exact ground state (shown in red). But combining both methods into an Alternate training procedure (in green) allows us to reach the correct solution.
The primary advantage of HELIA lies in its ability to use PSR for computing gradients in the orange block of gates, while utilizing $\mathfrak{g}$-sim for the blue block in Fig 3. Additionally, we introduce two novel training schemes and demonstrate their superior performance compared to standard PSR or $\mathfrak{g}$-sim independently. This improvement is illustrated in Fig.4 for a 6-qubit Transverse Field Ising Model (TFIM) Hamiltonian, where using only the orange block (HEA via PSR) or the blue block (multi-qubit Pauli gates via $\mathfrak{g}$-sim ) fails to reach the target energy in 500 iterations. However, combining the two blocks using an Alternate training scheme (discussed in Sec.III.2) successfully converges to the exact ground state energy. Finally, we investigate the barren plateau phenomenon and empirically demonstrate that our ansatz exhibits a slower decay of gradients with increasing qubit count, compared to a deep hardware-efficient circuit, thereby enabling the effective training of larger qubit models.
III.2 Training Scheme
For the selected PQC, two distinct training schemes are employed: Alternate and Simultaneous. We evaluate the proposed protocols on two different VQA applications: ground state estimation using VQE and the Quantum Phase Classification. The core idea in both the schemes is to exploit the ansatz structure and distribute the resources required for gradient evaluation to both classical and quantum hardware. This delegation not only significantly reduces the number of calls to the quantum hardware, but also increases the accuracy as shown in the results of VQE in Table 2 and 3 of Appendix A. In the subsequent text we will interchangeably refer to the orange block of gates as Block Q or PSR Block, and the blue block as Block G or $\mathfrak{g}$-sim Block, based on the protocol used in obtaining the associated gradients.
III.2.1 Alternate Training
This section outlines a training protocol that integrates both the PSR and $\mathfrak{g}$-sim to develop a hybrid algorithm for gradient evaluation and quantum model training. As described in Algorithm 2 and depicted in Fig. 5, this protocol alternates the gradient evaluation process between classical and quantum hardware.
In each optimization step we update the parameters of Block Q (from $\theta_{i}$ to $\theta_{i+1}$) in Fig.3 using standard PSR. This is followed by a step of updating the parameters of Block G (from $\phi_{i}$ to $\phi_{i+1}$) using $\mathfrak{g}$-sim based on the updated parameters $\theta_{i+1}$. The optimization process is repeated iteratively until convergence.
Figure 5: Workflow of Alternate training protocol based on Alg.2
Let’s consider the case where the Block Q is composed of YZ linear ansatz (forming an unitary $U_{q}(\theta_{i})$) with $p$ parameters and the Block G has $g$ parametrized gates corresponding to a DLA of size $g$ (forming an unitary $U_{g}(\phi_{i})$), and the initial state is $|\psi\rangle$. Obtaining the gradients of the Block Q will require atleast $2p$ different runs of the quantum circuit for PSR. Using the gradients, we update Block Q to $U_{q}(\theta_{i+1})$.
1
Input: $T_{max},\alpha,U_{q}(\mathbf{\theta_{0}}),U_{g}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
Initialize: $\mathbf{\phi}_{0},\mathbf{\theta}_{0}\sim Norm(0,1);i=0$
Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{g}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
begin
       while $i\leq T_{s}$ do
             PSR Step
2             Obtain gradients $\gradient_{\theta_{i}}C(\mathbf{\theta_{i}},\mathbf{\phi_{i}})$, w.r.t $\theta_{i}$ via PSR
3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\alpha\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
             $\mathfrak{g}$-sim Step
4             Prepare $|\psi(\theta_{i+1})\rangle=U_{q}(\theta_{i+1})|0^{\otimes n}\rangle$
5             for $\hat{O}_{m}\in\mathfrak{g}_{DLA}$ do
6                   Measure and store $o_{m}=\langle\psi(\theta_{i+1})|\hat{O}_{m}|\psi(\theta_{i+1})\rangle$
7            Obtain gradients $\gradient_{\phi_{i}}C(\theta_{i+1},\phi_{i})$, w.r.t $\phi_{i}$ via $\mathfrak{g}$-sim using $\{o_{m}\}_{\hat{O}_{m}\in\mathfrak{g}_{DLA}}$
8             $\mathbf{\phi}_{i+1}\leftarrow\phi_{i}-\alpha\gradient_{\phi_{i}}C(\theta_{i+1},\phi_{i})$
9             $i\leftarrow i+1$
10
11
12
Algorithm 2 Alternate Training for VQE
To evaluate gradients of the Block G, we measure the expectation value of all $g$ Pauli operators on the state $|\psi(\theta_{i+1})\rangle=U_{q}(\theta_{i+1})|\psi\rangle$ on quantum hardware. This data then feeds into the $\mathfrak{g}$-sim algorithm which can evaluate the gradients entirely on classical hardware using an automatic differentiation framework [69, 70]. Finally, the parameters of Block G are updated from $\phi_{i}$ to $\phi_{i+1}$, finishing one iteration. Hence, each of these iterations, costs us $2p+g$ circuit evaluations (ignoring the factor of number of shots required for estimating the expectation values). This is in contrast with the $2p+2g$ circuit evaluations required per iteration to train the entire circuit via PSR alone.
III.2.2 Simultaneous Training
In this section, we propose a modification to the Alternate training protocol that enables the simultaneous updating of parameters in each iteration.
The pseudo-algorithm is shown in Algorithm 3. The main difference with the Alternate training scheme is, here we update the Block G gates based on $\theta_{i}$ (not the updated $\theta_{i+1}$). This is essentially equivalent to obtaining gradients for the full ansatz by PSR, and updating all parameters simultaneously. However, we significantly reduce the resource requirements by delegating a portion of the gradient estimation task to classical hardware via $\mathfrak{g}$-sim . The required circuit evaluations per iteration is also the same as the above Alternate scheme, as the only difference comes in by using the non-updated unitary $U_{q}(\theta_{i})$ in the measurement step before $\mathfrak{g}$-sim .
1
Input: $T_{max},\alpha,U_{q}(\mathbf{\theta_{0}}),U_{g}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
Initialize: $\mathbf{\phi_{0}},\mathbf{\theta_{0}}\sim Norm(0,1);i=0$
Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{g}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
begin
       while $i\leq T_{s}$ do
             PSR Step
2             Obtain gradients $\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$, w.r.t $\theta_{i}$ via PSR
3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\alpha\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
             $\mathfrak{g}$-sim Step
4             Prepare $|\psi(\theta_{i})\rangle=U_{q}(\theta_{i})|0^{\otimes n}\rangle$
5             for $\hat{O}_{m}\in\mathfrak{g}_{DLA}$ do
6                   Measure and store $o_{m}=\langle\psi(\theta_{i})|\hat{O}|\psi(\theta_{i})\rangle$
7            Obtain gradients $\gradient_{\phi_{i}}C(\theta_{i},\phi_{i})$, w.r.t $\phi_{i}$ via $\mathfrak{g}$-sim using $\{o_{m}\}_{\hat{O}_{m}\in\mathfrak{g}_{DLA}}$
8             $\mathbf{\phi}_{i+1}\leftarrow\phi_{i}-\alpha\gradient_{\phi_{i}}C(\theta_{i},\phi_{i})$
9             $i\leftarrow i+1$
10
11
12
Algorithm 3 Simultaneous Training for VQE
In practice, we find the Alternate scheme for a fixed number of iterations followed by several iterations of the Simultaneous training provides the most optimal results. We will expand on this point in the subsequent sections.
IV Results
To evaluate the effectiveness of our method we employ several metrics,
Error:
$\text{Error}=\frac{E_{g}^{t}-E_{g}^{*}}{E_{g}^{*}}$ (13)
where we define $E_{g}^{t}$ as the lowest possible energy achieved in the current trial and $E_{g}^{*}$ is an estimate of the true ground state energy of the Hamiltonian. For large Hamiltonians (beyond 14 qubits) it is computationally impractical to numerically diagonalize it and obtain the true ground state energy. Instead, taking inspiration from [31], we use the lowest energy reached across all trials for the corresponding Hamiltonian as an estimate of the true ground state energy $E_{g}^{*}$.
Figure 6: XY Hamiltonian VQE (18 Qubits): The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G for 18 qubits. For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared. On the left, we show that both Alternate and Alternate+Simultaneous protocol has better success rates than PSR. For error (Column 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black). The Alternate method only performs well for Layer 1 configuration, which is improved by Alternate+Simultaneous, giving lower error than PSR for all layer configurations. Most importantly, on the right QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction). The spread in standard deviation is shown in black. We consistently see reduction in QPU calls using our methods, with frequently lower error values and higher success metrics as compared to PSR.
Figure 7: XY Hamiltonian VQE: The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G in each row, for qubit sizes ranging from 6 to 16. For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared. We consistently see reduction in QPU calls using our methods at high qubits, with frequently lower error values and higher success metrics as compared to PSR.
Figure 8: Gradients for XY Hamiltonian VQE: Variance of the first parameter from PSR and $\mathfrak{g}$-sim block are plotted with increasing qubits for both Full-PSR (blue) and our Alternate method (orange) for finding ground state of XY Hamiltonian using VQE. The diagrams also show the plots for varying number of PSR block layers (1,3,6,9 from top to bottom). For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green. The green curve is the same in all of the plots and provided as a reference for our improved gradients. In both $\mathfrak{g}$-sim and PSR block, we notice a slower decay of gradients as compared to HEA50, with the PSR block maintaining slightly higher magnitude than the $\mathfrak{g}$-sim block. Overall the plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
Success: Fraction of the trials that are able to reach below a error threshold of $10^{-3}$, which has been chosen ad-hoc. This gives an estimate of how often can the method get close to the true solution. Going forward, we report the relevant statistics only for the successful trials, except for $\mathfrak{g}$-sim as the error for this case in the mentioned examples is always higher than the threshold.
QPU Calls Reduction: Fraction of quantum circuit evaluations, referred to as Quantum Processing Unit (QPU) calls, reduced in current method as compared to standard training with PSR. We use the number of QPU calls required to reach the point of lowest energy in the corresponding trial as our estimate.
IV.1 Ground State Estimation using VQE
The proposed protocols can be efficiently used for VQE when the Hamiltonian spans a poly-DLA as shown in Sec. IV.1.1. However, for more general cases, where the Hamiltonian spans an exponential DLA, we can still use our method as a pre-training procedure. We discuss this in detail for Longitudinal Transverse Field Ising Model (LTFIM) in Sec. IV.1.2.
IV.1.1 Improved VQE for XY Spin Hamiltonian (poly-DLA)
We apply our chosen ansatz and proposed training schemes on VQE, and demonstrate its effectiveness for this task in comparison to PSR and $\mathfrak{g}$-sim . For this, we consider a XY Spin Hamiltonian given by,
$H_{XY}=\sum_{i}^{N-1}\alpha_{i}\hat{X}_{i}\hat{X}_{i+1}+\beta_{i}\hat{Y}_{i}\hat{Y}_{i+1}$ (14)
XY Spin Hamiltonian has a poly-DLA with the dimension given by $dim(\mathfrak{g})=n^{2}-n$. This is necessary to use $\mathfrak{g}$-sim effectively in our algorithm.
While running the $\mathfrak{g}$-sim -only experiments, all of the DLA elements of the XY spin Hamiltonian produce zero expectation value with the zero initial state $|0^{\otimes n}\rangle$, we added a layer of Hadamard gates $H^{\otimes n}$ to facilitate training . In order to maintain consistency we have added the $H^{\otimes n}$ layer before Block G in all of the examples for the Hamiltonian. Hence, via $\mathfrak{g}$-sim we optimize the following cost function,
$C(\phi)=\langle 0|H^{\otimes n}U_{g}^{\dagger}(\phi)H_{XY}U_{g}(\phi)H^{\otimes n}|0\rangle$ (15)
For the Full-PSR (all parameters of Block Q & G optimized using PSR), Alternate and Alternate + Simultaneous columns we optimize using a larger ansatz by adding the Block Q, and obtain the following cost function,
$C(\theta,\phi)=\langle 0|U_{q}^{\dagger}(\theta)H^{\otimes n}U_{g}^{\dagger}(\phi)H_{XY}U_{g}(\phi)H^{\otimes n}U_{q}(\theta)|0\rangle$ (16)
Now we are not restricted to only searching in the same polynomial space but can search a much larger state space, by virtue of the gates in Block Q. For visualization, we plot the success, error and QPU reduction metrics comparing Full-PSR, Alternate and Alternate + Simultaneous in Fig. 6 and 7, while the actual numbers along with comparison with $\mathfrak{g}$-sim is provided in Table 4 and 3 in Appendix A. The error for $\mathfrak{g}$-sim is significantly higher in the chosen examples, as compared to other methods. Hence, we ignore it in the plots, and only mention the values in Table 3. Later in IV.1.2, we also discuss how this method can be used as a starting point for finding ground states in more general Hamiltonians without the restriction of having a poly-DLA.
Although, we are exploring a larger state space, often the ground state might infact lie close to the space searched by $\mathfrak{g}$-sim , rendering a quantum method unnecessary for the task. This infact happens in the odd qubit examples for XY Hamiltonian, where $\mathfrak{g}$-sim is able get very low errors without additional gates. We mention this here, but in the numerics we only focus on even qubit cases to demonstrate where our method can be effective and standard $\mathfrak{g}$-sim fails. Some of the odd qubit examples are demonstrated in Appendix. B.
Experiment Details: We present the results for 4 different configurations for each qubit, with the number of YZ linear layers: 1,3,6 and 9, changing between the configurations. The results are given in Table 2 and 3 of Appendix A. The Full-PSR column refers to training the full ansatz using standard PSR, Alternate refers to training it by the Alternate training scheme till convergence and Alternate + Simultaneous trains the ansatz using Alternate Scheme for 500 iterations followed by Simultaneous training till convergence. In order to keep the comparison fair, we use an Adam optimizer with initial learning rate $0.01$ and keep it consistent across all experiments.
In Fig. 6 and 7, we have plotted the mean and standard deviation (in black) over the successful trials when measuring the QPU calls reduction. For error, we find the data is often asymmetrically spread around the mean, owing to being lower-bounded by 00 (by definition Eq.(13)) and upper bounded by $10^{-3}$ (success threshold). Hence, for this we resort to the median, and show the spread from $25\%$ to $75\%$ quantiles in black instead of mean and standard deviation. This is due to the fact that median is a more robust metric when the data is distributed asymmetrically. This unequal distribution is also explicit from the $25\%$ to $75\%$ quantiles which extend unequally from the median.
Mitigating Barren Plateaus
We also analyze the gradients of our ansatz for the given VQE task with varying number of qubits. In each case, we track the gradient of the first parameter from each of the Block Q (referred as PSR block) and Block G (referred as $\mathfrak{g}$-sim block) and plots its variance during the training (Fig.8). Each datapoint is obtained after evaluating the variance from the full training of 64 independent trials.
For comparison, we also perform the same VQE task using a 50 layer HEA (referred as HEA50), which is known to exhibit BP [81], and plot the variance of the first parameter of this circuit. In other words, while it is possible that our method will also run into the BP issue at a certain qubit count, atleast for the system sizes we have tested so far, the method holds up well against BP issues in previous methods. Secondly, we believe the reason the direct BP condition in [82, 83] is not applicable here because we choose gates from a poly-DLA and a shallow HEA circuit preventing us from forming a 2-design. However, we do not provide proof of this statement, and propose this as only a potential reason.
HELIA shows a clear improvement in terms of slow decay in gradients in Fig 8. The gradients from $\mathfrak{g}$-sim block demonstrate an exponential decrease, but with much less slope as compared to the HEA50 circuit, while the PSR block shows a sub-exponential decay. Overall, this allows for larger qubit models to be trained efficiently without running into vanishingly small gradients.
IV.1.2 Pre-training in LTFIM Hamiltonian (exp-DLA)
The ability to combine $\mathfrak{g}$-sim and PSR efficiently in our protocol depends on the poly-DLA of the $\mathfrak{g}$-sim block as well as the associated measurement operator. This limits the kind of Hamiltonians on which we can perform VQE task. However, we can still use our method as a starting point for general Hamiltonians with exponential DLA as demonstrated in Algorithm 4.
Although similar to the pre-training approach used in Ref. [31], our method is unique in its use of the full ansatz over the complete training procedure, with only the Hamiltonian changing between the two phases of training. The first phase can be thought of as a warm-start method, while the second phase trains the circuit to find ground-state of the full Hamiltonian starting from the parameters of the first phase. We expand this point further in the following paragraphs and highlight this important difference through a numerical example.
To start the protocol, we need a step of choosing a subset of operators that form a poly-DLA. In poly-DLA Phase, we perform Alternate, followed by Simultaneous training on the reduced Hamiltonian. During the Exponential DLA Phase of training in Alg.4, we train only the PSR block for some iterations before training the full block. This is done with the aim of keeping the overall quantum resources low, as the PSR block has fewer trainable parameters and hence needs less circuit evaluations to obtain gradients during training. Finally, all of the parameters are trained together using standard PSR.
We demonstrate the results of this protocol for longitudinal-transverse field Ising model (LTFIM),
$H_{LTFIM}=\sum_{j}\alpha_{j}X_{j}X_{j+1}+\sum_{j}(\beta_{j}Z_{j}+\gamma_{j}X_{j})$ (17)
which has an exponential DLA. A poly-DLA can be constructed by dropping the longitudinal terms
$i\mathfrak{g}_{TFIM}=\langle\{iX_{j}X_{j+1},iZ_{j}\}\rangle_{Lie}$ (18)
(a) LTFIM Hamiltonian with 10 qubits using PSR
(b) LTFIM Hamiltonian with 10 qubits using our proposed method
(c) LTFIM Hamiltonian with 12 qubits using PSR
(d) LTFIM Hamiltonian with 12 qubits using proposed method
Figure 9: Demonstration of PSR (9(a) and 9(c)) and our method (9(b) and 9(d)) to run VQE for LTFIM Hamiltonian (with an exponential DLA) in 10 & 12 qubit cases. The purple and red regions show Alternate and Simultaneous training respectively for the restricted Hamiltonian, followed by training the PSR block and full circuit in green and yellow regions respectively for the full LTFIM Hamiltonian. In both 10 & 12 qubit examples with 3 & 6 YZ linear configurations in the orange block, we notice that our training methods in the initial phase allows VQE to reach much closer to the target ground state during the final training, as compared to just using PSR for the full training, for similar number of of total training iterations. For the overall training, we also slightly reduce the QPU calls ($\sim 6-10\%$)
Experiment Details: The PSR block is composed of varying layers of YZ linear ansatz (3 and 6 layers), while the $\mathfrak{g}$-sim block is generated out of gates from $\mathfrak{g}_{TFIM}$. In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of the PSR block and 1000 iterations of full circuit training. Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods. However, these choices are made ad-hoc and can be modified based on other stopping conditions, such as Early Stopping, that can further reduce QPU calls.
Input: $T_{alt},T_{sim},T_{psr},T_{full},U_{q}(\mathbf{\theta_{0}}),H=\sum_{i}\beta_{i}P_{i}$
Output: $\theta^{*},\phi^{*}$
Initialize: $\vec{\theta}_{0}\sim\mathcal{N}(0,1);i=0$
begin
       poly-DLA Phase
1       Obtain a subset of operators $\{P_{j}\}_{j\in\mathcal{A}}$ from $H$ such that it forms a poly-DLA $\mathfrak{g}$.
2       Define $H_{poly}=\sum_{P_{i}\in\mathcal{A}}\beta_{i}P_{i}$, and unitary $U_{g}(\vec{\phi_{0}})=e^{-iP_{1}\phi_{1}^{0}}\ldots e^{-iP_{n}\phi_{n}^{0}}$
3       Randomly initialize $\vec{\phi}_{0}\sim\mathcal{N}(0,1)$
4       Define the cost function: $C_{poly}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})H_{poly}U_{g}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$
5       Run Alternate optimization given in Alg.2 for $T_{alt}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
6       Run Simultaneous optimization given in Alg.3 for $T_{sim}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
7       Extract the optimized parameters $(\vec{\theta}^{*},\vec{\phi}^{*})$
       Exponential DLA Phase
8       Define the full cost function, $C_{full}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})HU_{g}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$ initialized at $(\vec{\theta}^{*},\vec{\phi}^{*})$
9       Optimize only parameters $\vec{\theta}$ of $U_{q}(\vec{\theta})$ in $C_{full}(\vec{\theta},\vec{\phi})$ for $T_{psr}$ iterations using PSR
10       Optimize full cost function $C_{full}(\vec{\theta},\vec{\phi})$ for $T_{full}$ using PSR
11       Extract optimized parameters at the end of training $(\vec{\theta}^{*},\vec{\phi}^{*})$
12
Algorithm 4 VQE for General Hamiltonian
We plot the error metric in Eq.(13) for the LTFIM Hamiltonian with 10 & 12 qubits in Fig.9. Importantly, although the initial training phase is based on the reduced Hamiltonian, the plotted energy values correspond to expectation values of the full Hamiltonian. As shown in Fig.9, the algorithm is indeed able to converge to good solutions by using the our protocols in the first phase of the training.
Compared with PSR, our protocol shows lower error value on an average in Fig. IV.1.2 for both 10 & 12 qubits. More importantly, the QPU calls required for the 10-qubit Hamiltonian VQE using PSR is $54.5e+$ 5 and $56.3e+$ 5 for Layer 3 and 5 respectively, while our proposed method requires $54.065e+$ 5 ($\sim 9.6\%$ reduction) and $55.925e+$ 5 ($\sim 6.5\%$ reduction) QPU calls respectively. Similarly, for 12 qubit, PSR uses $56.12e+$ 5 and $58.28e+$ 5 for Layer 3 and 5 respectively while our method requires $55.466e+$ 5 ($\sim 10.7\%$ reduction) and $57.698e+$ 5 ($\sim 7\%$ reduction) respectively. Overall, our method shows lower error while reducing the required quantum resources.
IV.2 Classification of Quantum Phases
Figure 10: Classification Workflow: $\rho_{i}$ represents the initial state obtained from a training dataset, which is processed using $U_{q}(\theta)$ and $U_{g}(\phi)$ as described before. The measurement operator belongs to a chosen DLA, which also generates the gates for $U_{g}(\phi)$
Our training protocols are useful for QML tasks such as classification. To illustrate this we choose the bond-alternating spin-$1/2$ Heisenberg chain [114],
$H=J\sum_{i}\vec{S}_{2i-1}\cdot\vec{S}_{2i}+J^{\prime}\sum_{i}\vec{S}_{2i}\cdot\vec{S}_{2i+1}$ (19)
where $\vec{S}=(X,Y,Z)$ are the Pauli matrices. The model undergoes a second-order phase transition at $J/J^{\prime}=1$[114].
We pose the task of learning the quantum phase of the ground state as a classification problem for 10 & 12 qubit examples. Since, quantum phase transition only happens in the limit $N\rightarrow\infty$, it is not accurate to refer to these states as different phases. Rather, we just assign two separate labels ±$\pm 1$ to the ground states for $J<J^{\prime}$ and $J>J^{\prime}$.
Since the qubit counts are small enough, the training and test data (100 train + 100 test dataset) are generated by sampling $J,J^{\prime}$ uniformly from $[-1,1]$ numerically diagonalizing the Hamiltonian in Eq.(19) and obtaining the ground state. This ground state is used as an input into the PQC and the ansatz is trained to generate the correct labels after measuring a Pauli string, that is randomly chosen and fixed at the beginning of the experiment.
For designing the PQC, we encounter the choice of gates for PSR and $\mathfrak{g}$-sim blocks. For the PSR block, we stick to the YZ linear ansatz choice with 9 layers. For $\mathfrak{g}$-sim block however, we do not have a unique choice like we had for previous examples. Hence, we explore 3 different DLA choices to construct the gates:
$i\mathfrak{g}_{XY}$ =$=\langle\{iX_{j}X_{j+1},iY_{j}Y_{j+1}\}\rangle_{Lie}$ (20)
$i\mathfrak{g}_{YZ}$ =$=\langle\{iZ_{j}Z_{j+1},iY_{j}Y_{j+1}\}\rangle_{Lie}$ (21)
$i\mathfrak{g}_{ZX}$ =$=\langle\{iX_{j}X_{j+1},iZ_{j}Z_{j+1}\}\rangle_{Lie}$ (22)
For each of these configurations of DLA, we run 10 independent trials and compare test accuracy. The results are plotted in Fig.1.
We observe different peak accuracies for each of the DLA choices with $i\mathfrak{g}_{YZ}$ providing the highest accuracy in $\mathfrak{g}$-sim for 10 qubits ($0.854\pm 0.008$) and $i\mathfrak{g}_{ZX}$ providing the highest accuracy in $\mathfrak{g}$-sim for 12 qubits ($0.864\pm 0.013$) . In each case, the Alternate and Full-PSR method provide a higher peak accuracy on average, although only by $\sim 1-6\%$. However, the Alternate Method always requires much fewer QPU calls to reach peak accuracy as compared to Full-PSR as shown in Table 1.
In other words, almost same accuracy is reached with our Alternate Method but using much lesser QPU calls compared to only the PSR method, and with better accuracy compared to the $\mathfrak{g}$-sim method.
DLA Full PSR g-sim Alternate
XY 0.898 $\pm$ 0.009 0.860 $\pm$ 0.004 0.897 $\pm$ 0.018
YZ 0.911 $\pm$ 0.008 0.853 $\pm$ 0.005 0.907 $\pm$ 0.013
ZX 0.905 $\pm$ 0.007 0.864 $\pm$ 0.013 0.908 $\pm$ 0.007
(a) Test Accuracy across different methods
DLA Full PSR g-sim Alternate
XY 36736 $\pm$ 30481 - 13381 $\pm$ 6143
YZ 32800 $\pm$ 28216 - 7080 $\pm$ 3540
ZX 25256 $\pm$ 17172 - 11894 $\pm$ 9386
(b) QPU Calls across different methods
Table 1: Classification task using $\mathfrak{g}$-sim , Alternate and Full-PSR Training (with 9 layers in PSR block) for 12 qubits. Three different choices of poly-DLA (referred to as $\mathfrak{g}_{XY},\mathfrak{g}_{YZ}$ and $\mathfrak{g}_{ZX}$) are explored for $\mathfrak{g}$-sim , and also to construct the Block G for Alternate and Full-PSR Method. Average peak test accuracy and number of QPU calls required to reach peak accuracy is reported for each of the configuration. For all examples, the Alternate improves over the accuracy of $\mathfrak{g}$-sim , and is similar to the Full-PSR training with dramatically reduced QPU calls
V Conclusion and Outlook
In order to reach quantum advantage, we need to deploy larger and more complex algorithms on quantum computers while ensuring classical intractability. The reason this is critical for variational algorithms is that, since these are algorithms are heuristic in nature, identifying the thresholds for quantum utility or advantage is usually hard. Therefore, the rather seemingly pedestrian approach of deploying larger algorithms in practice and testing them rigorously, may actually turn out to produce valuable insights.
In this work, we provide a step in that direction by an informed ansatz design and combining $\mathfrak{g}$-sim with PSR, thus delegating the gradient estimation task to both classical and quantum hardware carefully. We extensively test our novel protocols on several practical tasks like VQE and Classification for qubit counts ranging from 6 to 18. In our examples, we notice significant reduction in QPU calls (even upto 60% for the case of VQE in 16 qubit XY Hamiltonian with 1 YZ linear layer), especially at higher number of qubits, alongside improvement in error and success metrics for VQE and accuracy for Classification. Importantly, we also notice a slow decay in gradients as we scale up in qubits, providing a potential strategy in mitigating the ominous Barren Plateau phenomenon. These improvements highlight the benefit of our hybrid gradient-estimation protocol in bringing larger and more complex quantum models closer to us eventually the way towards quantum utility.
We now discuss several open questions for further research that can improve our proposed methods :
  • In the numerical examples, we have restricted ourselves to only Pauli gates as the BCH formula takes a convenient expression in this case. Since in general, this step requires calculating an infinite series, efficient protocols for calculating or approximating it will allow for more interesting gate choices in $\mathfrak{g}$-sim .
  • The PSR method, although invaluable for training quantum models, has a high quantum overhead. Our proposed method is aimed at reducing the necessity of using PSR to only a limited number of parameters, but leaves open the possibility of improving PSR itself.
  • True quantum utility lies at much higher qubit counts and in using actual quantum hardware. We believe our methods will continue to show improvements beyond 20 qubits, and have more noise-robustness owing to fewer quantum calls required. However, a thorough analysis of the scaling behaviors are necessary.
  • Another promising direction to explore is extending the $\mathfrak{g}$-sim method [31], which could further enhance its scalability and efficiency, enabling it to handle a broader range of quantum systems with complex symmetries. The $\mathfrak{g}$-sim method [31], which efficiently simulates quantum systems by leveraging Lie algebraic properties, can potentially be extended to handle a broader class of systems by incorporating representation theory. By decomposing the space of linear operators acting on n-qubits into a direct sum of invariant subspaces under a Lie group’s action, the simulation could exploit symmetry properties to reduce computational complexity. This approach would enable selective simulation within invariant subspaces, reducing overhead and enhancing scalability. Representation theory provides a natural framework to identify these invariant subspaces and compute their contributions efficiently, facilitating the simulation of highly symmetric quantum systems.
    However, finding this decomposition into invariant subspaces is a challenging problem in representation theory. Advanced concepts such as Cartan decomposition, which separates the Lie algebra into solvable and semisimple parts, and the theory of highest weight vectors are crucial for identifying irreducible representations corresponding to invariant subspaces. These tasks involve intricate techniques for determining weight vectors, understanding their structure, and classifying modules. An example of such decomposition into invariant subspaces is demonstrated in Diaz et al. [100], where Majorana fermionic operators are used as basis sets for invariant subspaces. Readers interested in exploring these advanced representation theory concepts required to find such decomposition can refer to Ragone et al. [115].
  • Another direction would be to move away from the standard PQC structure to include non-unitary or general Completely Positive Trace Preserving (CPTP) maps such as in [116].
The authors thank quantum team at Fujitsu Research India Pvt.Ltd., especially Naipunnya Raj, Ruchira Bhat and Rajiv Sangle for their discussions and inputs. The authors extend their immense gratitude to Yasuhiro Endo, Hirotaka Oshima and Shintaro Sato as well as the entire Robust Quantum Computing Department at Fujitsu Limited for their strategic and technical support. The authors also thank Masayoshi Hashima for his valuable support and inputs regarding usage of the state of the art Fujitsu Quantum Simulator. The code used in this paper will be made available at a later date. A preliminary version of this paper has been presented at Fujitsu-IISc Quantum Workshop at Bangalore, India.
Appendix A Numerical Experiments on XY Hamiltonian
Here we tabulate the results of the numerical experiments comparing $\mathfrak{g}$-sim , standard PSR, Alternate and Alternate+Simultaneous method of training on the XY Hamiltonian for various qubit counts in Table 2 and 3, as described in Sec. IV.1.1 and visualized in Fig. 6 and 7. We run VQE on XY Hamiltonian for qubits ranging from 6 to 18, for the configurations with 1,3,6 and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G.
Overall, our method consistently shows reduction in error of estimating the ground state (upto an order of magnitude reduction), improved success rates of reaching correct solution (upto 39% increase) and reduction in QPU calls (upto $\sim 60\%$ reduction), hence reducing the overall quantum resources compared to standard PSR training.
Qubits Success(%) Full-PSR Alternate Alt+Sim 6 81.25 70.31 90.62 8 70.31 59.38 75.00 10 29.69 48.44 50.00 12 65.62 70.31 90.62 14 51.56 51.56 51.56 16 21.88 43.75 48.44 18 35.94 50.00 54.69 (a) Success rate for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits Success(%) Full-PSR Alternate Alt+Sim 6 75.00 81.25 75.00 8 68.75 65.62 68.75 10 37.50 40.62 48.44 12 59.38 70.31 73.44 14 40.62 40.62 43.75 16 29.69 34.38 34.38 18 25.00 34.38 42.19 (b) Success rate for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits Success(%) Full-PSR Alternate Alt+Sim 6 82.81 79.69 81.25 8 79.69 81.25 79.69 10 43.75 45.31 54.69 12 64.06 75.00 81.25 14 40.62 35.94 46.88 16 23.44 32.81 37.50 18 28.12 37.50 37.50 (c) Success rate for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits Success(%) Full-PSR Alternate Alt+Sim 6 82.81 81.25 98.44 8 54.69 57.81 60.94 10 43.75 43.75 54.69 12 71.88 85.94 90.62 14 57.81 48.44 53.12 16 42.19 43.75 32.81 18 28.12 29.69 54.69 (d) Success rate for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 4.27 $\pm$ 2.98 -33.27 $\pm$ 17.07 8 16.98 $\pm$ 10.62 27.53 $\pm$ 13.20 10 18.73 $\pm$ 10.87 41.88 $\pm$ 21.81 12 24.19 $\pm$ 12.65 45.25 $\pm$ 21.95 14 23.66 $\pm$ 13.55 32.07 $\pm$ 15.24 16 34.59 $\pm$ 16.09 60.00 $\pm$ 25.11 18 25.95 $\pm$ 12.14 46.81 $\pm$ 26.52 (e) QPU Reduction for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 7.71 $\pm$ 3.98 22.62 $\pm$ 12.21 8 20.36 $\pm$ 7.64 36.58 $\pm$ 15.15 10 22.57 $\pm$ 5.90 49.32 $\pm$ 18.98 12 25.48 $\pm$ 8.67 43.24 $\pm$ 16.37 14 39.11 $\pm$ 14.60 47.99 $\pm$ 20.32 16 39.50 $\pm$ 12.94 56.12 $\pm$ 21.31 18 39.02 $\pm$ 13.58 52.63 $\pm$ 22.83 (f) QPU Reduction for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 19.51 $\pm$ 6.45 7.93 $\pm$ 3.25 8 16.61 $\pm$ 4.52 22.87 $\pm$ 8.77 10 25.22 $\pm$ 7.24 32.82 $\pm$ 14.19 12 31.20 $\pm$ 10.25 22.63 $\pm$ 8.37 14 31.66 $\pm$ 11.57 46.14 $\pm$ 17.34 16 46.86 $\pm$ 20.55 50.88 $\pm$ 19.61 18 40.39 $\pm$ 16.54 45.45 $\pm$ 20.12 (g) QPU Reduction for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 -4.23 $\pm$ 1.25 -4.70 $\pm$ 2.13 8 14.00 $\pm$ 3.55 17.64 $\pm$ 7.73 10 18.27 $\pm$ 4.36 22.70 $\pm$ 10.14 12 27.63 $\pm$ 8.91 13.74 $\pm$ 3.93 14 25.34 $\pm$ 8.07 38.17 $\pm$ 14.61 16 30.14 $\pm$ 10.76 44.47 $\pm$ 19.66 18 27.78 $\pm$ 9.02 39.07 $\pm$ 15.71 (h) QPU Reduction for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Table 2: XY Hamiltonian (Success and QPU Reduction): The success and QPU calls reduction are shown for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G using Full-PSR, Alternate and Alternate + Simultaneous training (referred as Alt+Sim). The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR. The metrics show that our methods are generally better at reaching higher success rates and lower error(in Table 3), while the QPU calls are reduced significantly
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
6 12862.78 12856.2 12889.49 0.1272 0.1272 0.1272 0.1187 0.1102 0.1187 0.1187 0.1102 0.1187
8 7111.32 7110.25 7114.96 0.0454 0.0454 0.0454 0.0340 0.0227 0.0340 0.0340 0.0340 0.0340
10 359.87 358.91 362.73 0.0515 0.0515 0.1949 0.0441 0.0294 0.0809 0.0441 0.0441 0.0441
12 315.23 314.98 319.3 3.1575 0.6097 3.6670 3.1457 0.0353 3.7230 3.1457 0.0353 3.1457
14 940.51 933.67 956.65 0.0532 0.0532 0.0621 0.0355 0.0266 0.0443 0.0443 0.0443 0.0532
16 426.07 423.47 433.41 0.8264 0.0590 4.3119 0.0472 0.0236 0.0708 0.0472 0.0472 0.0590
18 118.38 117.59 119.3 0.0758 0.0758 0.1137 0.0716 0.0590 0.0842 0.0674 0.0632 0.0758
(a) Error for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
6 12862.78 12856.2 12889.49 0.1272 0.1272 0.1653 0.1187 0.1102 0.1780 0.1187 0.1187 0.1272
8 7111.32 7110.25 7114.96 0.0454 0.0454 0.0936 0.0340 0.0227 0.0993 0.0340 0.0340 0.0369
10 359.87 358.91 362.73 0.1802 0.0515 0.5828 0.3015 0.0827 1.7409 0.0441 0.0441 0.0515
12 315.23 314.98 319.3 3.1634 0.5950 4.0234 3.1575 0.6244 4.7009 1.5905 0.0353 3.1516
14 940.51 933.67 956.65 0.0621 0.0532 0.1685 0.1286 0.0532 0.7404 0.0532 0.0532 0.0532
16 426.07 423.47 433.41 0.1299 0.0649 0.6021 1.0507 0.1033 3.0192 0.0472 0.0472 0.0590
18 118.38 117.59 119.3 0.1390 0.0821 0.3032 3.7105 0.7981 9.5521 0.0674 0.0674 0.0758
(b) Error for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
6 12862.78 12856.2 12889.49 0.5676 0.1611 1.4060 0.5760 0.1441 1.4949 0.1272 0.1272 0.1611
8 7111.32 7110.25 7114.96 0.1589 0.0681 0.7207 0.5845 0.1475 1.7165 0.0454 0.0340 0.0567
10 359.87 358.91 362.73 0.4706 0.1893 0.9890 1.4413 0.3750 3.7428 0.0588 0.0515 0.0735
12 315.23 314.98 319.3 1.6377 0.1885 3.2753 3.6346 1.3578 5.5020 1.4786 0.0353 3.1575
14 940.51 933.67 956.65 0.1330 0.0709 0.2416 2.9881 1.1970 3.9945 0.0532 0.0532 0.0621
16 426.07 423.47 433.41 0.1417 0.1181 0.4014 5.3951 2.3257 8.6062 0.0590 0.0472 0.0708
18 118.38 117.59 119.3 0.1137 0.0927 0.1853 4.9487 2.4343 15.1262 0.0842 0.0674 0.1032
(c) Error for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
6 12862.78 12856.2 12889.49 0.5337 0.1272 4.7936 2.3969 0.4342 5.1028 0.1272 0.1187 0.3220
8 7111.32 7110.25 7114.96 0.5107 0.1759 1.7534 2.7351 0.6242 6.0377 0.0681 0.0454 0.0965
10 359.87 358.91 362.73 0.4559 0.2390 0.9798 2.8788 1.1104 5.4323 0.0735 0.0552 0.1177
12 315.23 314.98 319.3 2.0206 0.2121 3.5522 4.7009 3.1811 7.9762 1.5905 0.0589 3.1693
14 940.51 933.67 956.65 0.1064 0.0709 0.1862 4.2472 2.5581 5.7191 0.0709 0.0621 0.1131
16 426.07 423.47 433.41 0.1062 0.0590 0.2774 5.2593 3.0399 12.2541 0.0826 0.0708 0.1062
18 118.38 117.59 119.3 0.1390 0.0863 0.2843 6.6966 4.5486 13.9533 0.1011 0.0842 0.1137
(d) Error for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
Table 3: XY Hamiltonian VQE (Error) : The error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G. The error for $\mathfrak{g}$-sim with only XY Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods. Alternate and Alternate + Simultaneous method shows generally lower errors compared to the Full-PSR training
Appendix B Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
Although our aim has been to elevate $\mathfrak{g}$-sim to a hybrid scheme with a more expressive circuit, often it might not be necessary for a given task. In fact, the limited search space might be enough to reach a target solution as can be seen from the below examples. We use $\mathfrak{g}$-sim for finding the ground state of the XY Hamiltonian (Eq. 14) for 13 and 17 qubit cases, and the results are shown in Fig.11. Clearly, $\mathfrak{g}$-sim is able to converge to good solutions (Error $\sim 10^{-3}-10^{-5}$) without adding any additional gates.
(a) XY Hamiltonian with 13 qubits
(b) XY Hamiltonian with 17 qubits
Figure 11: Error vs Iteration curve for finding ground state of an XY Hamiltonian using only $\mathfrak{g}$-sim . For 13 (11(a)) and 17 (11(b)) qubits, $\mathfrak{g}$-sim is enough to obtain the ground state and does not require our protocol.
Appendix C VQE for TFIM Hamiltonian
We also implement our training schemes for VQE using a different Hamiltonian example, in order to demonstrate that the improvements are not restricted to a specific choice. For this, we use the Transverse Field Ising Models (TFIM) Hamiltonian,
$H_{TFIM}=\sum_{i}\alpha_{i}X_{i}X_{i+1}+\sum_{j}\beta_{j}Z_{j}$ (23)
which has a poly-DLA scaling as $2n^{2}-n$ for $n$ qubits. We measure the improvement in terms of Error, Success and QPU calls reduction as defined in IV.
We tabulate the numerical results comparing $\mathfrak{g}$-sim , Full-PSR, Alternate and Alternate+Simultaneous in Table 4 and 5 and provide plots for visualization in Fig.12. Except the 12 qubit case, where $\mathfrak{g}$-sim is able to produce accurate solutions (similar to the scenario mentioned in App.B), our proposed Alternate+Simultaneous method generally achieves a lower error and higher success rate compared to the other protocols, while almost always reducing the QPU calls necessary compared to Full-PSR.
Figure 12: TFIM Hamiltonian VQE: The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G, for various qubit counts in each row. For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared. For error (Column 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black). For QPU calls reduction only Alternate and Alternate+Simultaneous is plotted since we compare it relative to PSR (which will coincide with 0% QPU calls reduction). The spread in standard deviation is shown in black. We consistently see reduction in QPU calls using our methods, with frequently lower error values and higher success metrics as compared to PSR
Figure 13: Gradients for TFIM Hamiltonian VQE: Variance of the first parameter from PSR and $\mathfrak{g}$-sim block are plotted with increasing qubits for both Full-PSR (blue) and our Alternate method (orange) for finding ground state of TFIM Hamiltonian using VQE. The diagrams also show the plots for varying number of PSR block layers (1,3,6,9 from top to bottom). For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green. The green curve is the same in all of the plots and provided as a reference for our improved gradients. The plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
Qubits Success (%) Full-PSR Alternate Alt+Sim 8 48.44 50.00 45.31 10 18.75 9.38 26.56 12 25.00 28.12 32.81 14 100.00 100.00 100.00 (a) Success rate for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits Success (%) Full-PSR Alternate Alt+Sim 8 84.38 82.81 89.06 10 20.31 37.50 59.38 12 53.12 57.81 70.31 14 98.44 98.44 98.44 (b) Success rate for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits Success (%) Full-PSR Alternate Alt+Sim 8 93.75 95.31 100 10 43.75 45.31 64.06 12 75.00 76.56 76.56 14 100.00 100.00 100.00 (c) Success rate for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits Success (%) Full-PSR Alternate Alt+Sim 8 57.81 60.94 73.44 10 31.25 43.75 70.31 12 82.81 85.94 79.69 14 98.44 100.00 100.00 (d) Success rate for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 8 29.78 +/- 15.08 48.09 +/- 22.71 10 11.52 +/- 4.17 52.11 +/- 26.46 12 29.46 +/- 16.81 51.80 +/- 25.37 14 40.91 +/- 15.93 50.66 +/- 17.60 (e) QPU Reduction for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 8 17.45 +/- 5.79 30.08 +/- 13.40 10 17.97 +/- 1.97 38.72 +/- 16.46 12 25.66 +/- 6.49 35.29 +/- 15.03 14 48.79 +/- 15.07 25.09 +/- 8.19 (f) QPU Reduction for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 8 24.52 +/- 5.93 0.64 +/- 0.15 10 24.27 +/- 5.04 21.82 +/- 9.20 12 32.43 +/- 9.54 25.35 +/- 8.84 14 49.64 +/- 21.50 9.29 +/- 2.21 (g) QPU Reduction for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 8 0.68 +/- 0.11 1.39 +/- 0.54 10 9.33 +/- 1.70 6.55 +/- 2.57 12 27.27 +/- 8.83 14.90 +/- 5.16 14 45.79 +/- 22.98 -1.98 +/- 0.37 (h) QPU Reduction for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Table 4: TFIM Hamiltonian (Success and QPU Reduction): The success and QPU calls reduction are shown for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G using Full-PSR, Alternate and Alternate + Simultaneous training (referred as Alt+Sim). The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR. The metrics show that our methods are generally better at reaching higher success rates and lower error(in Table 5), while the QPU calls are reduced significantly
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
8 4034.16 4034.16 4034.17 0.0409 0.0327 0.0409 0.0409 0.0327 0.0409 0.0491 0.0409 0.0491
10 284.92 287.65 306.97 0.2309 0.0181 38.2376 0.2717 0.0430 14.3759 0.0272 0.0272 0.0362
12 0.04 0.04 0.04 0.0424 0.0424 0.0795 0.0424 0.0344 0.0609 0.0424 0.0424 0.0424
14 24.8 24.8 24.8 12.3975 12.3975 12.4076 12.3975 12.3975 12.4304 12.4178 12.4178 12.4178
(a) Error for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
8 4034.16 4034.16 4034.17 0.0450 0.0266 0.2270 0.0573 0.0327 0.1636 0.0491 0.0491 0.0491
10 284.92 287.65 306.97 3.3506 2.1643 17.4322 4.9852 0.5818 32.6050 0.0362 0.0294 0.0453
12 0.04 0.04 0.04 0.0636 0.0318 0.2358 0.5723 0.0530 1.7380 0.0530 0.0424 0.0636
14 24.8 24.8 24.8 12.4178 12.3874 12.5901 12.4989 12.3975 12.8790 12.4279 12.3975 12.4381
(b) Error for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
8 4034.16 4034.16 4034.17 0.5031 0.1432 1.2640 1.0553 0.2372 2.1598 0.0736 0.0491 0.0982
10 284.92 287.65 306.97 4.1430 1.0731 32.6571 3.1785 0.9961 12.7414 0.0634 0.0453 0.0996
12 0.04 0.04 0.04 0.2278 0.0715 0.4292 1.2929 0.2331 3.5501 0.0636 0.0530 0.0742
14 24.8 24.8 24.8 12.4279 11.3585 12.5420 13.3352 11.9059 15.0483 12.4381 11.2799 12.4786
(c) Error for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Qubits Error (1e-5)
$\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
Median Q25 Q75 Median Q25 Q75 Median Q25 Q75 Median Q25 Q75
8 4034.16 4034.16 4034.17 9.7762 0.6708 45.0443 9.9071 3.4442 40.1194 0.1473 0.0736 4.1968
10 284.92 287.65 306.97 7.0227 1.3991 22.5464 21.5390 10.7876 51.6152 0.3441 0.0815 38.5591
12 0.04 0.04 0.04 0.1484 0.0636 0.4875 2.8719 0.9644 7.3757 0.0848 0.0636 0.1325
14 24.8 24.8 24.8 12.4685 7.7801 12.6256 17.4254 13.2718 23.5684 12.4381 9.5845 12.4887
(d) Error for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
Table 5: TFIM Hamiltonian (Error): The error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G. The error for $\mathfrak{g}$-sim with only TFIM Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods. Alternate and Alternate + Simultaneous method shows generally lower errors compared to the Full-PSR training
We also provide plots of variance of gradients for the first parameter from each of $\mathfrak{g}$-sim and PSR block in Fig. 13. Consistent with our previous example, the gradients decay slower using our choice of ansatz as compared to a deep Hardware Efficient Ansatz with 50 layers.

### v2

@@
 Abstract
-Quantum computers offer a promising route to tackling problems that are classically intractable such as in prime-factorization, solving large-scale linear algebra and simulating complex quantum systems, but require fault-tolerant quantum hardware.
-On the other hand, variational quantum algorithms (VQAs) have the potential to provide a near-term route to quantum utility or advantage, and is usually constructed by using parametrized quantum circuits (PQCs) in combination with a classical optimizer for training.
-Although VQAs have been proposed for a multitude of tasks such as ground-state estimation, combinatorial optimization and unitary compilation, there remain major challenges in its trainability and resource costs on quantum hardware.
-Here we address these challenges by adopting Hardware Efficient and dynamical LIe algebra Supported Ansatz (HELIA), and propose two training schemes that combine an existing $\mathfrak{g}$-sim method (that uses the underlying group structure of the operators) and the Parameter-Shift Rule (PSR).
+Quantum computers offer a promising route to tackling problems that are classically intractable such as in prime-factorization, solving large-scale linear algebra and simulating complex quantum systems, but potentially require fault-tolerant quantum hardware.
+On the other hand, variational quantum algorithms (VQAs) are a promising approach for leveraging near-term quantum computers to solve complex problems.
+However, there remain major challenges in their trainability and resource costs on quantum hardware.
+Here we address these challenges by adopting Hardware Efficient and dynamical LIe algebra supported Ansatz (HELIA), and propose two training methods that combine an existing classical-enhanced $\mathfrak{g}$-sim method and the quantum-based Parameter-Shift Rule (PSR).
 Our improvement comes from distributing the resources required for gradient estimation and training to both classical and quantum hardware.
-We numerically test our proposal for ground-state estimation using Variational Quantum Eigensolver (VQE) and classification of quantum phases using quantum neural networks.
-Our methods show better accuracy and success of trials, and also need fewer calls to the quantum hardware on an average than using only PSR (upto 60% reduction), that runs exclusively on quantum hardware.
+We numerically evaluate our approach for ground-state estimation of 6 to 18-qubit Hamiltonians using the Variational Quantum Eigensolver (VQE) and quantum phase classification for up to 12-qubit Hamiltonians using quantum neural networks.
+For VQE, our method achieves higher accuracy and success rates, with an average reduction in quantum hardware calls of up to 60% compared to purely quantum-based PSR.
+For classification, we observe test accuracy improvements of up to 2.8%.
 We also numerically demonstrate the capability of HELIA in mitigating barren plateaus, paving the way for training large-scale quantum models.
@@
 Despite recent breakthroughs in implementing quantum error correction [10], it may still take many years [11, 12, 13] to develop fault-tolerant quantum hardware.
-Current quantum devices face significant challenges, including number of qubits, qubit coherence time and gates with limited fidelity [10, 14, 15, 16, 17, 18, 19].
-Limitations of current quantum computing hardware has spurred focus towards variational quantum algorithms (VQA).
+Current quantum devices face significant challenges, including low number of qubits, qubit coherence time and gates with limited fidelity [10, 14, 15, 16, 17, 18, 19].
+Limitations of current quantum computing hardware have spurred focus towards variational quantum algorithms (VQA).
 These are hybrid quantum-classical approaches that utilize parameterized quantum circuits (PQC) to address challenges across diverse fields, including machine learning, optimization, and ground-state energy calculations.
@@
 The core idea is to train quantum circuits with tunable parameters and optimizing these parameters using a classical optimizer for an objective function related to the problem being solved.
-In the training process of VQAs, the parameters of PQC are updated via either gradient-free or gradient-based optimization techniques.
+In VQAs, the parameters of PQC are updated via either gradient-free or gradient-based optimization techniques.
 The gradient-free methods [20, 21, 22, 23, 24, 25] are preferred when calculation of gradient is challenging or the loss function is not differentiable, but their performance might not scale well with problem size and system noise [26, 27].
 In contrast, the literature on gradient-based training of quantum machine learning (QML) primarily emphasizes the efficient computation of gradients of parameters of trainable quantum circuits on quantum hardware.
-The most commonly employed techniques include the parameter-shift rule (PSR) [28] and its various generalization or extensions [29, 30].
-PSR requires running two or more quantum circuit evaluations with shifted parameter values, for each trainable parameter in the circuit, and it run exclusively on quantum hardware.
-Classical simulation methods can be an effective alternative for advancing quantum technologies by emulating quantum systems using classical resources.
-Goh et al.
-proposed an efficient circuit simulation and gradient computation method called $\mathfrak{g}$-sim , reducing the resource burden leveraging the Lie-algebraic structure of variational quantum circuits [31].
-Clifford Perturbation theory has been recently put forward to approximate operator expectation values in near-Clifford circuits [32].
-These techniques are crucial for adapting classical backpropagation methods in a quantum context [33].
-Optimization in quantum computing is often challenging due to several factors.
+The commonly employed techniques include the parameter-shift rule (PSR) [28] and its various generalization or extensions [29, 30].
+PSR requires running two or more quantum circuit evaluations with shifted parameter values for each trainable parameter in the circuit, and it runs exclusively on quantum hardware.
+VQAs are often challenging due to several factors.
 A major hurdle being the barren plateau (BP) phenomenon [31], where gradients vanish exponentially as qubit count increases, making it difficult to find numerically meaningful updates for large circuits.
-Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization.
 Several BP-free models have been proposed in literature like quantum convolutional neural networks [32, 33] and permutation-equivariant quantum neural networks [34].
-Recent research also points towards the possibility that all BP-free quantum models might indeed be efficiently classically simulable [38], raising doubts about achieving any quantum advantage in the future.
-Classical algorithms for simulating general quantum circuits do exist, but they suffer from exponential computational and memory overhead [39].
-Over the years, several methods have been developed for accurate and efficient classical simulation of certain quantum circuits.
-These methods often approximate the circuits in specific regimes or restrict the types of gates and measurements allowed, leveraging special structures in the problems.
-These include MPS-based tensor networks methods that can simulate shallow noisy circuits for hundreds of qubits with limited gate fidelity [40].
-$\mathfrak{g}$-sim is another method that offers an efficient simulation algorithm.
-This method is effective when both the generators of the ansatz and the measurement operator are within a dynamical Lie algebra (DLA) [31], whose dimension scales polynomially with the number of qubits.
-We will explain more about the $\mathfrak{g}$-sim method in a later section as it is central to our ansatz construction.
-Recent advancements in classical simulation methods have significantly enhanced our ability to estimate and optimize quantum circuits, even in the presence of noise.
-Fontana et.al proposed low weight efficient simulation algorithm (LOWESA) to classically estimate expectation values of parametrized quantum circuits in noisy hardware [41].
-The authors recreate a classical surrogate of cost landscape in the presence of device noise by ignoring Pauli terms beyond a certain frequency threshold.
-Rudolph et.al extended this to noiseless quantum simulation and classically simulated the 127-qubit quantum utility experiment [42, 43] while Angrisani et.
-al further demonstrated provable guarantees for most noiseless quantum circuits [44].
-Recently, Lerch et.al showed it is always possible to identify patches for which similar classical surrogate can be generated [45].
-These techniques should be accounted for when constructing variational models with potential for quantum utility.
+Recent research also points towards the possibility that BP-free quantum models might be efficiently classically simulable [35].
+These issues need to be taken into consideration when designing new VQA, in order to avoid such pitfalls.
 The optimization landscape in quantum circuits is also highly non-convex, with numerous local minima and saddle points that can trap algorithms and complicate the search for global minima [36, 37, 38, 39].
 Additionally, quantum measurements are inherently stochastic requiring many runs to obtain accurate estimates and thus potentially slowing down optimization.
-Even for models that are trainable, computing gradients for quantum circuits is often quantum resource-intensive as methods such as the parameter-shift rule or finite differences necessitate multiple circuit executions, which scale linearly [33] with the number of parameters and shots.
+Even for models that are trainable, computing gradients for quantum circuits is often quantum resource-intensive as methods such as the PSR or finite differences necessitate multiple circuit executions, which scale linearly [40] with the number of parameters and shots.
 These factors, coupled with hardware limitations such as qubit connectivity and decoherence [10, 14, 15, 16, 17, 18, 19], make effective gradient-based optimization in quantum computing a complex task, frequently driving the need for alternative or hybrid methods.
-To mitigate the high quantum resource costs associated with training VQAs, we propose a hybrid method that integrates $\mathfrak{g}$-sim and the PSR.
-The gradient estimation task is delegated to both classical and quantum hardware.
-This approach aims to reduce the number of Quantum Processing Unit (QPU) calls, thereby saving resources.
-Our research focuses on training quantum models, emphasizing the use of specific ansätze and customized training schemes to enhance the training process and improve accuracy in tasks like ground state estimation using VQE and classification using quantum neural networks (QNNs).
+Classical algorithms for simulating general quantum circuits do exist, but they suffer from exponential computational and memory overhead [41].
+Several methods have been developed for efficient classical simulation of certain classes of quantum circuits, leveraging special structures in the problems.
+These include Matrix Product State-based tensor networks methods that can simulate shallow noisy circuits for hundreds of qubits with limited gate fidelity [42].
+Clifford Perturbation theory has been recently put forward to approximate operator expectation values in near-Clifford circuits [43].
+$\mathfrak{g}$-sim is another method that offers an efficient simulation algorithm, based on the study of the Lie group and the associated Lie algebra $\mathfrak{g}$, which is generated by the PQC.
+This method is effective when both the generators of the ansatz and the measurement operator are within a dynamical Lie algebra (DLA) [44], whose dimension scales polynomially with the number of qubits.
+Initially porposed by Somma et al.
+[45], the techniques were reframed in a modern context tailored to the quantum computing community by Goh et al.
+[44].
+We will explain more about the $\mathfrak{g}$-sim method in a later section as it is central to our ansatz construction.
+Although simulating arbitrary quantum circuits is hard, recent advancements in classical simulation methods have significantly enhanced our ability to estimate quantum circuits in specific parameter regions and in the presence of noise.
+Fontana et al.
+proposed low weight efficient simulation algorithm (LOWESA) to classically estimate expectation values of PQCs in noisy hardware [46].
+The authors recreate a classical surrogate of the cost landscape in the presence of device noise by ignoring Pauli terms beyond a certain frequency threshold.
+Rudolph et al.
+extended this to noiseless quantum simulation and classically simulated the 127-qubit quantum utility experiment [47, 48] while Angrisani et.
+al further demonstrated provable guarantees for most noiseless quantum circuits [49].
+Recently, Lerch et al.
+showed it is always possible to identify patches for which similar classical surrogate can be generated [50].
+These techniques should be accounted for when constructing variational models with potential for quantum utility.
+Figure 1: Overview of our main contributions (a) We propose Hardware Efficient and dynamical LIe algebra supported Ansatz (HELIA) composed of two blocks of gates $U_{q}$ and $U_{\mathfrak{g}}$ whose gradients can be obtained using parameter-shift rule and $\mathfrak{g}$-sim respectively.
+The resources required for gradient evaluation of the full ansatz can hence be delegated to both quantum and classical hardware.
+We further propose two training methods: Alternate and Simultaneous which benefit from the hybrid gradient of HELIA.
+(b) To evaluate the gradients of $U_{\mathfrak{g}}$, we use a $\mathfrak{g}$-sim based method.
+The operators of a dynamical Lie algebra are measured after applying $U_{q}$ (using QPU in the leftmost block), which is then passed on to classical hardware that evaluates the cost function and gradients using $\mathfrak{g}$-sim (using CPU in the middle block).
+Further details of our contribution are elaborated in Sec.
+III.
+Using our proposed ansatz and training methods we are able to reduce quantum hardware usage, improved accuracy and mitigate Barren Plateaus as elaborated in Sec.
+IV.
+We focus on enhancing quantum model training through tailored ansätze and customized training methods, improving accuracy in tasks such as ground-state estimation with the VQE and quantum phase classification using quantum neural networks (QNNs).
+To mitigate the high quantum resource costs associated with training VQAs, we propose Hardware Efficient and dynamical LIe algebra supported Ansatz (HELIA) and two hybrid methodsAlternate and Simultaneousthat integrate $\mathfrak{g}$-sim and PSR (Fig.
+1).
+Our gradient estimation task is delegated to both classical and quantum hardware to reduce the number of Quantum Processing Unit (QPU) calls, thereby saving resources.
 While this study does not evaluate the practical utility of these models or their performance relative to classical counterparts, it aims to explore the potential of scalable and accurate trainable quantum models.
-Our hope is that with increase in the size and accuracy of the trainable models we can have enough overlap with useful quantum models and move towards quantum utility.
-We will expand on the main results in the coming sections.
-The paper is divided as follows.
+Our findings suggest that increasing the scale and accuracy of trainable models could bridge the gap toward practical quantum utility.
+The remaining sections are divided as follows.
 We introduce the relevant background literature in Sec.
 II.
-We then discuss our contributions in terms of the choice of quantum circuits and two training schemes namely Alternate and Simultaneous in Sec.
+We then discuss our contributions in terms of the choice of quantum circuits and two training methods namely Alternate and Simultaneous in Sec.
 III.
 In Sec.
-IV we demonstrate the improvement in various metrics through extensive numerics.
+IV we demonstrate the improvement in various metrics through extensive numerical simulations.
 Finally, we end with the Conclusion and Outlook in Sec.
@@
 By integrating PQCs with classical optimization techniques, VQAs are specifically tailored to tackle computationally demanding problems across domains such as quantum chemistry, combinatorial optimization, and machine learning [51, 52, 53, 54, 55, 56].
-VQAs involves an $n$-qubit state $\rho$ (usually encoding problem data), acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ which undergoes evolution through a parameterized quantum unitary.
+VQAs involve an $n$-qubit state $\rho$ (usually encoding problem data), acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ which undergoes evolution through a parameterized quantum unitary.
 The unitary transformation is expressed as a sequence of $L$ parameterised unitary gates
-$U(\boldsymbol{\theta})=\prod_{i=1}^{L}U_{i}(\theta_{i})W_{i}$ (1)
-where $\boldsymbol{\theta}=[\theta_{1},\theta_{2},\ldots,\theta_{L}]$ denotes a set of trainable real-valued parameters and $W_{i}$ are non-trainable gates.
-A slightly more general scenario can be when each $U_{i}(\theta_{i})$ is composed of multiple trainable parametrized unitaries itself.
-However, we ignore that for ease of notation, and the same ideas can be extended to the general unitaries as well.
+$U(\boldsymbol{\theta})=\prod_{i=1}^{L}U_{i}(\boldsymbol{\theta}_{i})W_{i}$ (1)
+where $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots,\boldsymbol{\theta}_{L}]$ denotes a set of trainable real-valued parameters and $W_{i}$ are non-trainable gates.
+An example of such case is shown in Fig.
+2, where $U_{i}(\boldsymbol{\theta}_{i})$ is be composed of multiple trainable single-qubit gates, while the non-trainable block contains entangling CNOT gates.
 The expectation value of a Hermitian observable $\hat{O}$ is subsequently measured on the evolved quantum state, using a quantum device, given as
-$l_{\boldsymbol{\theta}}(\rho,\hat{O})=Tr[U^{\dagger}(\boldsymbol{\theta)}\rho U(\boldsymbol{\theta})\hat{O}].$ (2)
+[…]
 Classical optimizers are utilized to perform the optimization task
@@
 where we minimize some loss function $F$ of the expectation value depending on the task.
-For example, when finding ground state of a Hamiltonian $H$ using VQE, the loss function is the expectation value of $H$ on the trainable quantum circuit.
-For classification problem, the loss function can be the mean squared error (MSE) between predicted and true labels.
-The algorithm iteratively optimizes a cost function using the output of a quantum circuit, enabling the exploration of complex solution spaces that are potentially intractable for classical approaches.
-However, despite their promise, VQA face several critical challenges, including barren plateaus in the optimization landscape, limited scalability with increasing problem size, and the detrimental effects of noise and decoherence on circuit fidelity [34, 50, 56].
+The form of the loss function in Eq.
+(3) can vary depending on the task.
+For example, when estimating ground states the loss function can be the expectation value of a Hamiltonian, and finding the ground state is equivalent to minimizing this expectation value.
+For classification task, the loss function can be mean squared error between predicted labels from a quantum model and the true labels obtained from a dataset.
+In both cases, the algorithm iteratively optimizes a cost function using the output of a quantum circuit, enabling the exploration of complex solution spaces that are potentially intractable for classical approaches.
+Despite their promise, VQA face several critical challenges, including barren plateaus in the optimization landscape, limited scalability with increasing problem size, and the detrimental effects of noise and decoherence on circuit fidelity [31, 51, 57].
 Overcoming these obstacles necessitates advancements in algorithmic design, optimization techniques, and the development of robust quantum hardware.
-Algorithms considered in this work
-In this study, we focus on two applications of VQAs, illustrating their potential to address complex quantum problems.
-The first application investigates the Variational Quantum Eigensolver (VQE), a hybrid quantum-classical algorithm widely used for determining the ground state of quantum systems and calculating their corresponding energies [57].
-VQE has become a cornerstone in quantum chemistry and materials science, enabling the efficient simulation of molecular systems and facilitating the discovery of new materials.
-Recent research has also incorporated VQE into Quantum Error Detection pipeline [58, 59].
-The second application delves into quantum phase classification, where VQAs are utilized to identify and distinguish different quantum phases of matter [35, 60, 61].
-This application leverages the unique capabilities of quantum circuits to encode and process quantum states, enabling the precise detection of phase transitions and the systematic classification of quantum phases.
-II.2 Existing PQCs
 […]
@@
 […]
-×L
-Figure 1: Quantum circuit block for YZ linear ansatz.
+[…]
+Figure 2: Quantum circuit block for YZ linear (HEA) ansatz.
 The parametrized $R_{y}$ and $R_{z}$ rotations are applied on each qubit followed by $CNOT$ between neighboring qubits in a linear fashion.
-The blocks are often repeated to increase entanglement and expressivity
-The available options for PQC is quite vast and it is not a priori clear which circuit structure is ideal for the task at hand.
-We will consider three particular ansatz choices and motivate their use for our ansatz construction.
-One often considered ansatz choice is made based on qubit connectivity, to reduce circuit depth, by using the class of Hardware Efficient Ansatz (HEA) [62].
-Blocks of single-qubit rotations followed by entangling two-qubit gates are repeated for a chosen number of layers.
-An example within this class is $YZ$ linear ansatz shown in Fig.1, where each block consists of a $Y$-rotation and a $Z$-rotation applied on each qubit, followed by CNOT gates between neighbouring qubits in a linear fashion.
+The blocks are often repeated to increase entanglement and expressivity.
+The available options for PQC in VQAs is quite vast and it is not a priori clear which circuit structure is ideal for the task at hand.
+One common choice is Hardware Efficient Ansatz (HEA), which consists of blocks of single-qubit rotations followed by entangling two-qubit gates repeated for a chosen number of layers [58].
+An example within this class is the $YZ$ linear ansatz shown in Fig.
+2, where each block consists of a $Y$-rotation and a $Z$-rotation applied on each qubit, followed by CNOT gates between neighbouring qubits in a linear fashion.
 Commonly selected problem-motivated ansatz include the Unitary Coupled Cluster (UCC) [59], which is widely utilized for investigating ground states of fermionic molecular Hamiltonians, and the Quantum Alternating Operator Ansatz (QAOA) [60, 61, 62], often applied to combinatorial optimization problems.
@@
 Given the task of finding ground state of a Hamiltonian $H=\sum_{i}H_{i}$, the ansatz is composed of unitary blocks of the form
-$U(\boldsymbol{\theta})=\prod_{k}e^{-i\theta_{k}H_{k}}.$ (4)
-Building on the insights from [31], this approach can be extended to include terms from the DLA of $\{H_{k}\}_{H_{k}\in H}$, provided the DLA has a dimension that scales polynomially with the number of qubits.
-In our examples, we will combine the HVA ansatz with the HEA ansatz. For a detailed overview of the different ansatz options explored in the literature, refer [50].
-After selecting an appropriate ansatz based on the architecture and the task, the subsequent step involves determining the optimal parameters of the ansatz to address the problem being solved.
-This process, referred to as the training method, is elaborated upon in the following section.
-II.3 Training Methods
-The training process aims to identify the optimal parameters of the selected ansatz for a given architecture.
+$U(\boldsymbol{\theta})=\prod_{i}e^{-i\theta_{i}H_{i}},$ (4)
+and can repeated for a chosen number of layers.
+Building on the insights from Ref.
+[44], this approach can be extended to include terms from the dynamical Lie Algebra (DLA) of $\{iH_{i}\}_{H_{i}\in H}$, provided the DLA has a dimension that scales polynomially with the number of qubits.
+We elaborate on the details of DLA in Sec.
+II.2.2.
+In this study, we will combine the HVA ansatz with the HEA ansatz.
+For a detailed overview of the ansatz options explored in the literature, we refer to Ref.
+[51].
+II.2 Training Methods
+After selecting an appropriate ansatz, the subsequent step involves determining the optimal parameters of the ansatz to address the problem being solved.
 This is an iterative procedure, with each iteration comprising a loss function evaluation followed by a parameter update step.
-Although several gradient-free methods exist for parameter updation [20, 21, 22, 23], our primary focus here is on gradient-based techniques which require efficient and accurate algorithms to compute the partial derivatives of the cost function for effective training of variational quantum models.
+Although several gradient-free methods exist for parameter update [20, 21, 22, 23], here we focus on gradient-based techniques, which require efficient and accurate algorithms to compute the partial derivatives of the cost function.
 Contemporary classical machine learning models, which often contain billions of parameters, efficiently compute gradients using the backpropagation algorithm [64].
 This method requires only a single forward and backward pass to compute the gradients for all parameters simultaneously.
-Since classical backpropagation cannot be directly applied to quantum circuits [33], the Parameter-Shift Rule [28] is typically employed to compute circuit gradients.
+Since classical backpropagation cannot be directly applied to quantum circuits [40], the PSR [65, 28] is typically employed to compute circuit gradients.
 Alternatively, when the quantum circuit can be efficiently simulated classically, automatic differentiation frameworks, such as PyTorch [66] and Tensorflow [67], can be utilized for gradient evaluation.
 A notable case within this second scenario is $\mathfrak{g}$-sim , where placing certain restriction on quantum gates and measurement operators enables efficient classical simulation of the circuit.
-These two methods are discussed in detail in the following sections.
-II.3.1 Parameter-shift Rule
-Parameter-shift Rule (PSR) is commonly used for accurately estimating gradients of a PQC.
-For a loss function of the form given in Eq.(2), where each unitary is defined as $U_{i}(\theta_{i})=e^{-i\theta_{i}G_{i}}$ for some hermitian operator $G_{i}$, the partial derivatives are given by
-$\frac{\partial{\ell}}{\partial{\theta_{k}}}$ $=i\bra{0}{U}^{\dagger}_{\geq k}G_{k}({U}^{\dagger}_{<k}{\hat{O}}U_{<k})U_{\geq k}\ket{0}$
-$-i\bra{0}U^{\dagger}_{\geq k}(U^{\dagger}_{<k}\hat{O}{U}_{<k})G_{k}{U}_{\geq k}\ket{0}$ (5)
-where we defined $U_{<k}=U_{1}W_{1}U_{2}W_{2}\ldots U_{k-1}W_{k-1}$ and $U_{\geq k}=U_{k}W_{k}U_{k+1}W_{k+1}\ldots U_{n}W_{n}$.
-PSR can be employed to obtain a closed-form expression of Eq.(II.3.1) when the eigenspectrum of the generator $G_{k}$ either has only two unique eigenvalues or is symmetric and evenly spaced [71, 28, 72, 73, 30].
-Considering the case where $G_{k}$ with only two unique eigenvalues ±$\pm r$ i.e., $G_{k}^{2}=r^{2}I$ the corresponding unitary can be rewritten as
-$U_{k}(\theta)=\cos{r\theta}I-ir^{-1}G_{k}\sin{r\theta}$ ) (6)
-Following [28], through some simple algebra the above equation can be turned into
-$\frac{\partial{\ell}}{\partial{\theta_{k}}}$ $=\frac{r}{2}\Big{(}\langle\psi|(I-ir^{-1}G)^{\dagger}Q(I-ir^{-1}G)|\psi\rangle$
-$\langle\psi|(I+ir^{-1}G)^{\dagger}Q(I+ir^{-1}G)|\psi\rangle\Big{)}$ (7)
-where we have defined $|\psi\rangle=U_{\geq k}|0\rangle$ and $Q={U}^{\dagger}_{<k}{\hat{O}}U_{<k}$. Finally using Eq.(6) and substituting $\theta=\pi/4r$, the partial derivative simplifies to
-$\frac{\partial{\ell}}{\partial{\theta_{k}}}=r\Big{(}\ell(\boldsymbol{\theta}+\frac{\pi}{4r}\hat{e}_{k})-\ell(\boldsymbol{\theta}-\frac{\pi}{4r}\hat{e}_{k})\Big{)},$ (8)
-a difference between the loss functions evaluated on either side of the parameter with a specific eigenvalue-related shift.
-By measuring the loss function at two shifted values of the parameter $\theta_{k}$, the partial derivative can be estimated accurately upto shot noise.
-For more general eigenspectrum with equispaced eigenvalues, the parameter shift-rule can be generalized as shown in [30, 73].
+II.2.1 Parameter-shift Rule (PSR)
+PSR is commonly used for accurately estimating gradients in a PQC [28] using quantum hardware.
+Consider the loss function $\ell$ in Eq.
+(2), where the unitary $U_{i}(\theta_{i})$ is of the form $U_{i}(\theta_{i})=\exp{-i\theta_{i}\hat{G}_{i}}$ ) with $G_{i}$ having eigenvalues ±$\pm r$.
+The derivative of the loss function with respect to $\theta_{i}$ can be expressed as (Appendix A for a detailed derivation):
+$\frac{\partial{\ell}}{\partial{\theta_{i}}}=r\Big{(}\ell(\boldsymbol{\theta}+\frac{\pi}{4r}\hat{e}_{i})-\ell(\boldsymbol{\theta}-\frac{\pi}{4r}\hat{e}_{i})\Big{)}.$ (5)
+By measuring the loss function at two shifted values of the parameter $\theta_{i}$, the partial derivative can be estimated accurately up to shot noise [28].
+For more general eigenspectrum with equispaced eigenvalues, the parameter shift-rule can be generalized as shown in [68, 29, 30].
 Markovich et.al.
-have further extended this procedure to irregeular eigenspectrum as well [74].
+have further extended this procedure to irregular eigenspectrum as well [69].
 The computation of a gradient for a single parameter in a quantum circuit requires executing the circuit at least twice, apart from the measurement overhead to obtain expectation value accurately.
-In quantum machine learning (QML), backpropagation is not directly applicable because it relies on the ability to store or reuse information [33], which poses scalability challenges.
 This limitation is illustrated through a specific example.
@@
 For certain examples of PQC construction, the training may scale more favorably.
-One such example is proposed in Bowles et.
-al to address the scalability issue by introducing structured circuits consisting of blocks of commuting parametrized quantum gates [76].
-However, the utility of such structured PQCs is still not fully understood.
-II.3.2 $\mathfrak{g}$-sim Method
+One such example is proposed by Bowles et al.
+to address the scalability issue by introducing structured circuits consisting of blocks of commuting parametrized quantum gates [71, 72].
+Brnović et al.
+further extended this approach using layerwise-commuting PQC [73].
+However, the utility of such structured PQCs still need to be fully understood.
+II.2.2 $\mathfrak{g}$-sim Method
 In this section, we briefly explain the details of $\mathfrak{g}$-sim method proposed by Goh et al.
 [44], that is fully classical in nature.
-$\mathfrak{g}$-sim method is based on the study of the Lie group and the associated Lie algebra $g$, which is generated by the parametrized quantum circuit.
-Note that the foundational techniques used in [31] for this gradient estimation method are well-established by Somma et al.
-[77], Goh et al.
-primary contributed to reframe these techniques in a modern context tailored to the quantum computing community [31].
-Figure 2: Quantum circuit for $\mathfrak{g}$-sim algorithm corresponding to the unitary in Eq. (9)
 Consider a problem in which an $n$-qubit state $\rho$, acting on a Hilbert space $\mathcal{H}={(\mathbb{C}^{2})}^{\bigotimes n},$ is sent through a PQC $U(\boldsymbol{\theta})$ of the form
 $U(\boldsymbol{\theta})=\prod_{\ell=1}^{L}\prod_{k=1}^{K}e^{i\theta_{\ell k}H_{k}},$ (6)
-where $\boldsymbol{\theta}=[\theta_{11},\theta_{12},\ldots,\theta_{LK}]$ is a set of trainable real-valued parameters, and $\{H_{1},H_{2},\ldots,H_{K}\}$ are $K$ Hermitian operators that are the gate generators of the circuit, as shown in Fig 2.
+where $\boldsymbol{\theta}=[\theta_{11},\theta_{12},\ldots,\theta_{LK}]$ is a set of trainable real-valued parameters, and $\{H_{1},H_{2},\ldots,H_{K}\}$ are $K$ Hermitian operators that are the gate generators of the circuit.
 The vector space spanned by all possible nested commutators of $\{iH_{1},\dots,iH_{K}\}$, obtained by repeatedly taking the commutator between all elements until no new linearly independent element emerge is known as the DLA.
@@
 (6).
-$\mathfrak{g}$-sim method is applicable only under two assumptions [31]:
-  • either the measurement operator or the initial state of a variational quantum circuit belong to the DLA of the ansatz
+The $\mathfrak{g}$-sim method is applicable only under two assumptions [44]:
+  • either the measurement operator or the initial state of a variational quantum circuit belong to the DLA of the ansatz,
   • the dimension of DLA scales polynomially with the number of qubits.
-In order to understand, how $\mathfrak{g}$-sim method works, we need to know how the elements of the DLA $\mathfrak{g}$ transform under commutation with elements of the DLA $\mathfrak{g}$, and under conjugation with elements of the dynamical Lie group $\mathcal{G}$.
-As $\mathfrak{g}$ is closed under commutation, the commutator between any two elements in $\mathfrak{g}$ is an element of $\mathfrak{g}$ and can be decomposed as a linear combination of a Schmidt-orthonormal basis $\{iG_{\alpha}:\alpha\in\{1,2,\ldots,\text{dim}(g)\}\}$ of $\mathfrak{g}$.
-In particular,
-[$\commutator{iG_{\alpha}}{iG_{\beta}}=\sum_{\gamma=1}^{\text{dim}(\mathfrak{g})}f_{\alpha\beta}^{\gamma}iG_{\gamma}$ (10)
-where $f_{\alpha\beta}^{\gamma}=\text{Tr}[iG_{\gamma}[iG_{\alpha},iG_{\beta}]]\in\mathbb{R}$ is computed via the standard projection in vector space $\mathfrak{g}$.
-The coefficients $f_{\alpha\beta}^{\gamma}$ are called the structure constants of the DLA $\mathfrak{g}.$
-Input: $T_{max},\alpha,H=\sum_{i}^{g_{max}}a_{i}G_{i},\linebreak U(\vec{\theta})=e^{-iG_{1}\theta_{1}}\ldots e^{-iG_{n}\theta_{n}}$
-Output: Optimized $\vec{\theta^{*}}$
-Initialize: $\vec{\theta}\sim Norm(0,1)$
-Define: $C(\vec{\theta})=\langle 0^{\otimes q}|U^{\dagger}(\vec{\theta})HU(\vec{\theta})|0^{\otimes q}\rangle$
-Set: $H^{n}=H;a^{n}_{k}=a_{k},H^{n}_{k}=a_{k}^{n}G_{k}$
-begin
-       $i=0$
-       while $i\leq T_{max}$ do
-             $m=n$
-             while $m\geq 1$ do
-                   $k=0$
-                   for $k\leq g_{max}$ do
-1                         $H_{k}^{m-1}=e^{iG_{m}\theta_{m}}(a_{k}^{m}G_{k})e^{-iG_{m}\theta_{m}}$
-                         $\quad\quad=\sum_{j}b^{m-1}_{j}G_{j}$
-                         $k\leftarrow k+1$
-2
-3                  $H^{m-1}=\sum_{i}^{g_{max}}H_{i}^{m-1}=\sum_{i}^{g_{max}}a_{i}^{m-1}G_{i}$
-4
-                  $m\leftarrow m-1$
-            $l=0$
-             while $l\leq g_{max}$ do
-5                   $g_{l}=\langle 0^{\otimes q}|G_{l}|0^{\otimes q}\rangle$
-                   $l\leftarrow l+1$
-6            $C\leftarrow\sum_{i}^{g_{max}}a_{i}^{0}g_{i}$
-7             Obtain gradients $\gradient_{\vec{\theta}}C$ classically using an automatic differentiation framework
-8             $\vec{\theta}\leftarrow\vec{\theta}-\alpha\gradient_{\vec{\theta}}C$
-9
-10            $i\leftarrow i+1$
-11
-12
-13
-Algorithm 1 $\mathfrak{g}$-sim Implementation for VQE
-Using the above DLA structure we demonstrate an implementation of the full $\mathfrak{g}$-sim sub-routine in action for VQE in Alg.1.
-We choose a Hamiltonian with a DLA of polynomial dimension.
-The ansatz is chosen as $U(\boldsymbol{\theta})=e^{-i\theta_{1}G_{1}}\ldots e^{-i\theta_{n}G_{n}}$, where the generators are from the same DLA, allowing us to efficiently use the DLA structure in our computation.
-The crucial factor facilitating the use of DLA structure is the Baker-Campbell-Hausdorff (BCH) formula [79] in Step. 1, which is given by
-[…]
-For operators $G_{i},G_{j}$ belonging to a chosen DLA, the expression for $e^{i\theta G_{i}}G_{j}e^{-iG_{i}}$ consists of only nested commutators.
-Since the DLA is closed under commutation operation, all of the terms in the infinite series in Eq.(11) belong to the DLA itself.
-By choosing a DLA scaling polynomially in the number of qubits, we ensure the expansion only has polynomially many unique operators, whose co-efficients are functions of the structure constants $f_{\alpha\beta}^{\gamma}$ as shown in Eq.(10).
-However, it might still be difficult to obtain a closed form expression for Eq.(11).
-Hence, we further simplify it to the scenario when all the operators are n-qubit Pauli strings.
-This gives a very simple expression
-[…]
-where $P_{i},P_{j}$ are n-qubit Pauli strings.
-We further make our numerics computationally efficient by leveraging the $2n\times 2n$ binary symplectic matrix representation of Pauli, rather than their $2^{n}\times 2^{n}$ unitary matrix representation (Section 2.1 of [80]).
-As a result, we can replace matrix multiplication by bit-wise addition, which is much faster.
-Fast application of the BCH formula due to choosing Pauli strings as well a DLA of polynomial dimension allows to calculate the cost function efficiently with polynomial compute resources.
-The gradients can then be efficiently calculated using an automatic-differentiation framework such as Pytorch [69], Tensorflow [70].
+For more details on implementation and theory behind $\mathfrak{g}$-sim we refer the reader to Appendix B.
 The limitations of the PSR and the efficiency of $\mathfrak{g}$-sim provide motivation for combining them into a hybrid approach to achieve efficient gradient evaluation.
 Apart from scalability and resource challenges in gradient evaluation, the training of quantum models remains challenging, primarily due to the vanishing gradient phenomenon, also known as the BP problem, where gradients decay exponentially as the number of qubits increases, making larger models effectively untrainable.
-We expand on this point in the following subsection.
-II.4 Barren Plateau
-Optimization in quantum algorithms encounters significant challenges due to the intrinsic limitations imposed by quantum measurements.
-Classical optimization methods rely on precise loss evaluations or accurate gradient information, both of which are compromised by the finite precision and stochastic nature of quantum measurements.
-Barren plateaus (BPs) in quantum variational algorithms are characterized by regions in the parameter space where the gradient of the loss function becomes exponentially small, severely impeding the efficient optimization of quantum circuits [81, 82, 83].
-The presence of a BP landscape implies that, for most parameter configurations, an exponentially large number of measurement samples is required to reliably estimate the loss gradient, which is essential for effective optimization.
-The Variational Quantum Algorithm (VQA) is believed to exhibit a BP when variations in the model parameters $\theta$ in the unitary operation $U(\theta)$ lead to only exponentially small changes in the loss function $l_{\theta}(\rho,\hat{O})$, or magnitude of gradients $\partial l_{\theta}(\rho,\hat{O})/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\theta$) [34, 82, 83].
-Cerezo et al., in their investigation of cost function locality, established that the inclusion of global observables $(\hat{O})$ in the cost function results in the emergence of exponentially vanishing gradients, irrespective of the circuit’s shallow depth [84].
-Additional factors contributing to this phenomenon include the degree of entanglement [85, 86] and the presence of noise [87].
-An unified theoretical framework, employing group theoretical principles, covering above factors were presented in [82, 83].
-Beyond theoretical insights, substantial efforts have been devoted to developing practical strategies for mitigating barren plateaus [88].
-These include advanced initialization techniques, such as layer-wise training [89], Bayesian-inspired parameter settings [90], meta-learning for initializing parameters [91], initializing circuits with sequences of identity blocks [92], and transfer-learning-based initialization approaches [93].
-Alternative optimization frameworks have also been introduced, including adaptive learning rate strategies [94] and novel line search methods designed to facilitate navigation through flat loss landscapes [95].
-Mhiri et.al.
-provided an unified analysis on warm-starting approaches in BP-affected loss landscape [96].
-Classical neural network-based methodologies have also been explored, such as generating parameters for parameterized quantum circuits (PQCs) using neural networks [97, 98].
-Empirical studies, such as those conducted in Ref.
-[99], have validated the effectiveness of these approaches in reducing the prevalence of barren plateaus in practical applications, including quantum neural networks.
-Additionally, targeted strategies for specific ansatz designs, such as Hamiltonian variational ansatz and parameterized matchgate circuits, have demonstrated their potential in mitigating or entirely avoiding barren plateaus [100, 101].
-However, recent research has shed light on the classical simulability of BP-free models, offering valuable insights into their underlying structure.
-Cerezo et al.
-highlights that BP-free models are classically simulable when classical data is gathered during an initial data acquisition phase from the quantum device.
-This conclusion stems from the fact that barren plateaus arise from the curse of dimensionality, with mitigation strategies effectively reducing the problem to subspaces that are classically tractable [38].
-Finally, using classical processing to improve the results from quantum device is not unique to our work, and has been used in [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112].
+For a survey of BP phenomenon and proposed mitigation techniques, refer to Appendix C.
+Finally, using classical processing to improve the results from quantum device is not unique to our work, and has been used in Refs.
+[75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86].
 In fact, a recent paper also deals with the idea of processing different parts of the quantum circuit using classical and quantum hardware with Clifford Perturbation Theory [87].
-However, our approach of combining $\mathfrak{g}$-sim and the PSR in a hybrid iterative fashion offers a balanced method to address the barren plateau problem, improve resource efficiency, and maintain non-classical simulability.
+However, our approach of combining $\mathfrak{g}$-sim and the PSR in a hybrid iterative fashion offers a balanced method to address the BP problem, improve resource efficiency, and maintain non-classical simulability.
 This makes it a promising area for further research and development in quantum computing.
 III Our Contributions
-Gathering insights from current literature, we find extensive challenges across all building blocks of a variational quantum algorithm trained on today’s hardware.
-In this paper, we address issues related to choice of PQC, resource-inefficiency of gradient evaluation and trainability of large models in regards to scaling the number of parameters and the concurrent barren plateau issues that appear.
-We go through each of the improvements in a sequential manner highlighting clear instances of where and how the improvements are obtained using detailed numerical analysis.
+We address issues related to :
+  • choice of PQC,
+  • resource-inefficiency of gradient evaluation,
+  • trainability of large models in regards to scaling the number of parameters, and the concurrent BP issues that appear.
+We go through each of the improvements, highlighting clear instances of where and how the improvements are obtained using detailed numerical analysis.
+Algorithms considered in this work
+In this study, we focus on two applications of VQAs, illustrating their potential to address complex quantum problems.
+The first application investigates the Variational Quantum Eigensolver (VQE), a hybrid quantum-classical algorithm widely used for determining the ground state of quantum systems and calculating their corresponding energies [88].
+VQE has become a cornerstone in quantum chemistry and materials science, enabling the efficient simulation of molecular systems and facilitating the discovery of new materials.
+Recent research has also incorporated VQE into Quantum Error Detection pipeline [89, 90].
+In this work, we emply VQE to find ground states of XY, Transverse Field Ising Model (TFIM), Longitudinal-Transverse Field Ising Model (LTFIM) and LiH Hamiltonian.
+The second application delves into quantum phase classification, where VQAs are utilized to identify and distinguish different quantum phases of matter [32, 91, 92] of the bond-alternating spin-$1/2$ Heisenberg chain [93].
+This application leverages the unique capabilities of quantum circuits to encode and process quantum states, enabling the precise detection of phase transitions and the systematic classification of quantum phases.
 III.1 PQC Selection
-Our choice of PQC is motivated from the standpoint of reducing quantum resources for training as well as avoiding Barren Plateaus.
+Our choice of PQC is motivated from the standpoint of reducing quantum resources for training as well as avoiding BPs.
 In general, simulating the full density matrix requires keeping track of all the $n$-qubit Pauli operators which scales as $14^{n}-$ 1.
-However, to effectively utilize $\mathfrak{g}$-sim without incurring exponential computational and memory overhead, it is essential to restrict the training ansatz and measurement operator associated with a QML task to a poly-DLA.
-This can be a challenge, for example in VQE if the ground state lies far outside the polynomial DLA(poly-DLA) being explored (as shown in Table 3 in Appendix A), or a QML task whose accuracy gets limited by the polynomial search space (as shown in Fig.1).
-Figure 3: The HELIA comprises two blocks of quantum gates: the orange block or Block Q ($U_{q}$) consists of layers of HEA (non-DLA part), followed by the blue block or Block G ($U_{g}$), which includes multi-qubit Pauli rotation gates with generators forming the basis of a task-specific DLA.
-The final measurement operator also belongs to the same DLA
-To elevate $\mathfrak{g}$-sim to a higher expressibility, we propose a hybrid ansatz consisting of a parameterized non-DLA block $U_{q}$, which cannot be classically simulated, followed by a parameterized $\mathfrak{g}$-sim block $U_{g}$ that can be classically simulated within the $\mathfrak{g}$-sim framework.
+However, to effectively utilize $\mathfrak{g}$-sim without incurring exponential computational and memory overhead, it is essential to restrict the training ansatz and measurement operator for a QML task to a poly-DLA.
+This can be a challenge, for example in VQE if the ground state lies far outside the polynomial DLA (poly-DLA) being explored (as shown in Table 4 in Appendix D), or a QML task whose accuracy gets limited by the polynomial search space (as shown in Table 2).
+Choice of unitary blocks : To elevate $\mathfrak{g}$-sim to a higher expressibility, we propose a hybrid ansatz consisting of a parameterized non-DLA block of gates $U_{q}$, which cannot be classically simulated, followed by a parameterized DLA block of gates $U_{\mathfrak{g}}$ that can be classically simulated within the $\mathfrak{g}$-sim framework.
 A specific ansatz falling within this proposal is the Hardware Efficient and dynamical LIe algebra Supported Ansatz (HELIA), shown in Fig.
-3, where $U_{q}$ represents the hardware efficient ansatz (HEA).
-$U_{g}$ is composed of multi-qubit Pauli rotations $\exp{-i\theta P}$ ), where the Paulis $P$ form an orthonormal basis of a chosen poly-DLA.
-Both HEA [62] and Lie algebra Supported Ansatz [83] are known in literature, however our proposal aims at combining them into a single ansatz, which we will show has several important advantages.
-This specific ordering of gates is important for our protocol, as we require the gates and measurement operator chosen from the poly-DLA to be adjacent in the quantum circuit for $\mathfrak{g}$-sim to be used efficiently.
-If the order is reversed, the Heisenberg-evolved operator obtained after applying the HEA gates on the measurement operator will have terms from outside the poly-DLA, making $\mathfrak{g}$-sim quite ineffective.
-In our examples of VQE and Classification, we use a YZ linear ansatz, shown in Fig.
-1 (which belongs to the HEA class), with varying layers, and the DLA is motivated by the problem being solved.
-Figure 4: VQE performed for a 6-qubit TFIM Hamiltonian.
-Only using HEA using PSR (shown in orange) or multi-qubit Pauli rotation gates using $\mathfrak{g}$-sim (shown in blue) is not enough to reach exact ground state (shown in red).
-But combining both methods into an Alternate training procedure (in green) allows us to reach the correct solution.
-The primary advantage of HELIA lies in its ability to use PSR for computing gradients in the orange block of gates, while utilizing $\mathfrak{g}$-sim for the blue block in Fig 3.
-Additionally, we introduce two novel training schemes and demonstrate their superior performance compared to standard PSR or $\mathfrak{g}$-sim independently.
-This improvement is illustrated in Fig.4 for a 6-qubit Transverse Field Ising Model (TFIM) Hamiltonian, where using only the orange block (HEA via PSR) or the blue block (multi-qubit Pauli gates via $\mathfrak{g}$-sim ) fails to reach the target energy in 500 iterations.
-However, combining the two blocks using an Alternate training scheme (discussed in Sec.III.2) successfully converges to the exact ground state energy.
-Finally, we investigate the barren plateau phenomenon and empirically demonstrate that our ansatz exhibits a slower decay of gradients with increasing qubit count, compared to a deep hardware-efficient circuit, thereby enabling the effective training of larger qubit models.
-III.2 Training Scheme
-For the selected PQC, two distinct training schemes are employed: Alternate and Simultaneous.
-We evaluate the proposed protocols on two different VQA applications: ground state estimation using VQE and the Quantum Phase Classification.
-The core idea in both the schemes is to exploit the ansatz structure and distribute the resources required for gradient evaluation to both classical and quantum hardware.
-This delegation not only significantly reduces the number of calls to the quantum hardware, but also increases the accuracy as shown in the results of VQE in Table 2 and 3 of Appendix A.
-In the subsequent text we will interchangeably refer to the orange block of gates as Block Q or PSR Block, and the blue block as Block G or $\mathfrak{g}$-sim Block, based on the protocol used in obtaining the associated gradients.
+1, where $U_{q}$ represents the hardware efficient ansatz (HEA).
+$U_{\mathfrak{g}}$ is composed of multi-qubit Pauli rotations $\exp{-i\theta P}$ ), where the Paulis $P$ form an orthonormal basis of a chosen poly-DLA.
+In our numerical experiments, we use the YZ linear ansatz, shown in Fig.
+2 (which belongs to the HEA class) for the non-DLA circuits, and the Lie algebra Supported Ansatz is motivated by the problem being solved.
+Ordering of unitary blocks : The specific ordering of gates is important for $\mathfrak{g}$-sim to be used efficiently in our proposal.
+This is because we require the gates and measurement operator chosen from the poly-DLA to be adjacent in the quantum circuit.
+To elaborate, let’s consider the operator $\hat{H}$, chosen from a poly-DLA.
+If we look at the Heisenberg-evolved operator created by applying $U$, we obtain the following $H^{\prime}=U^{\dagger}HU$.
+When the generators of $U$ are chosen from the poly-DLA the operator $H^{\prime}$ is efficiently computable using $\mathfrak{g}$-sim as we only need to track polynomially many operators.
+This is the case when we consider unitary of type $U_{g}(\phi)$.
+On the other hand, if $U$ is composed of a general set of gates similar to $U_{q}(\theta)$, the expression for $H^{\prime}$ may not be classically tractable as it can contain exponentially many operators.
+Both HEA [58] and Lie algebra Supported Ansatz [94] are studied in literature.
+Our proposal, however, aims at combining them into a single ansatz, which we will show has several important advantages.
+The primary advantage of HELIA lies in its ability to use PSR for computing gradients in $U_{q}$, while utilizing $\mathfrak{g}$-sim for the $U_{\mathfrak{g}}$.
+Additionally, we introduce two training methods and demonstrate their superior performance compared to standard PSR or $\mathfrak{g}$-sim independently.
+Finally, we investigate the BP phenomenon and demonstrate that our ansatz exhibits a slower decay of gradients with increasing qubit count, compared to a deep hardware-efficient circuit, thereby enabling the effective training of larger qubit models.
+III.2 Training Method
+For the selected PQC, two distinct training methods are employed: Alternate and Simultaneous.
+We evaluate the proposed methods on two different VQA applications: ground state estimation using VQE and Quantum Phase Classification.
+The core idea in both the methods is to exploit the ansatz structure and distribute the resources required for gradient evaluation to both classical and quantum hardware.
+This delegation not only significantly reduces the number of calls to the quantum hardware, but also increases the accuracy as shown in the results of VQE in Table 3 and 4 of Appendix D.
 III.2.1 Alternate Training
-This section outlines a training protocol that integrates both the PSR and $\mathfrak{g}$-sim to develop a hybrid algorithm for gradient evaluation and quantum model training.
+This section outlines a training method that integrates both the PSR and $\mathfrak{g}$-sim to develop a hybrid algorithm for gradient evaluation and quantum model training.
 As described in Algorithm 1 and depicted in Fig.
-5, this protocol alternates the gradient evaluation process between classical and quantum hardware.
-In each optimization step we update the parameters of Block Q (from $\theta_{i}$ to $\theta_{i+1}$) in Fig.3 using standard PSR.
-This is followed by a step of updating the parameters of Block G (from $\phi_{i}$ to $\phi_{i+1}$) using $\mathfrak{g}$-sim based on the updated parameters $\theta_{i+1}$.
+3, this method alternates the gradient evaluation process between classical and quantum hardware.
+Although, we use VQE to explain our training methods, it can be easily extended to Classification task without changing the core idea.
+In each optimization step we update the parameters of $U_{q}$ (from $\theta_{i}$ to $\theta_{i+1}$) using standard PSR.
+This is followed by a step of updating the parameters of $U_{\mathfrak{g}}$ (from $\phi_{i}$ to $\phi_{i+1}$) using $\mathfrak{g}$-sim based on the updated parameters $\theta_{i+1}$.
 The optimization process is repeated iteratively until convergence.
-Figure 5: Workflow of Alternate training protocol based on Alg.2
-Let’s consider the case where the Block Q is composed of YZ linear ansatz (forming an unitary $U_{q}(\theta_{i})$) with $p$ parameters and the Block G has $g$ parametrized gates corresponding to a DLA of size $g$ (forming an unitary $U_{g}(\phi_{i})$), and the initial state is $|\psi\rangle$.
-Obtaining the gradients of the Block Q will require atleast $2p$ different runs of the quantum circuit for PSR.
-Using the gradients, we update Block Q to $U_{q}(\theta_{i+1})$.
+Figure 3: Workflow of Alternate training method based on Alg.1
+Let’s consider the case where $U_{q}$ is composed of YZ linear ansatz (forming an unitary $U_{q}(\theta_{i})$) with $p$ parameters and $U_{\mathfrak{g}}$ has $g$ parametrized gates corresponding to a DLA of size $g$ (forming an unitary $U_{g}(\phi_{i})$), and the initial state is $|\psi\rangle$.
+Obtaining the gradients of $U_{q}$ will require atleast $2p$ different runs of the quantum circuit for PSR.
+Using the gradients, we update $U_{q}$ to $U_{q}(\theta_{i+1})$.
 1
-Input: $T_{max},\alpha,U_{q}(\mathbf{\theta_{0}}),U_{g}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
+Input: $T_{max},\eta,U_{q}(\mathbf{\theta_{0}}),U_{\mathfrak{g}}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
 Initialize: $\mathbf{\phi}_{0},\mathbf{\theta}_{0}\sim Norm(0,1);i=0$
-Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{g}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
+Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{\mathfrak{g}}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
 begin
-       while $i\leq T_{s}$ do
+       while $i\leq T_{max}$ do
              PSR Step
 2             Obtain gradients $\gradient_{\theta_{i}}C(\mathbf{\theta_{i}},\mathbf{\phi_{i}})$, w.r.t $\theta_{i}$ via PSR
-3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\alpha\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
+3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\eta\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
              $\mathfrak{g}$-sim Step
@@
 Algorithm 1 Alternate Training for VQE
-To evaluate gradients of the Block G, we measure the expectation value of all $g$ Pauli operators on the state $|\psi(\theta_{i+1})\rangle=U_{q}(\theta_{i+1})|\psi\rangle$ on quantum hardware.
-This data then feeds into the $\mathfrak{g}$-sim algorithm which can evaluate the gradients entirely on classical hardware using an automatic differentiation framework [69, 70].
-Finally, the parameters of Block G are updated from $\phi_{i}$ to $\phi_{i+1}$, finishing one iteration.
-Hence, each of these iterations, costs us $2p+g$ circuit evaluations (ignoring the factor of number of shots required for estimating the expectation values).
+To evaluate gradients of $U_{\mathfrak{g}}$, we measure the expectation value of all $g$ Pauli operators on the state $|\psi(\theta_{i+1})\rangle=U_{q}(\theta_{i+1})|\psi\rangle$ on quantum hardware.
+The measured values are fed into the $\mathfrak{g}$-sim algorithm to evaluate the gradients entirely on classical hardware using an automatic differentiation framework [66, 67].
+Finally, the parameters of $U_{\mathfrak{g}}$ are updated from $\phi_{i}$ to $\phi_{i+1}$, finishing one iteration.
+Hence, each of these iterations, costs us $2p+g$ circuit evaluations (ignoring the factor of number of shots required for estimating the expectation values.
+Since, the measurement outcome is processed differently for gradient evaluation of $U_{q}$ and $U_{\mathfrak{g}}$, the number of shots required for the respective case might be different).
 This is in contrast with the $2p+2g$ circuit evaluations required per iteration to train the entire circuit via PSR alone.
 III.2.2 Simultaneous Training
-In this section, we propose a modification to the Alternate training protocol that enables the simultaneous updating of parameters in each iteration.
+In this section, we propose a modification to the Alternate training method that enables the simultaneous updating of parameters in each iteration.
 The pseudo-algorithm is shown in Algorithm 2.
-The main difference with the Alternate training scheme is, here we update the Block G gates based on $\theta_{i}$ (not the updated $\theta_{i+1}$).
+The main difference with the Alternate training method is, here we update $U_{\mathfrak{g}}$ based on $\theta_{i}$ (not the updated $\theta_{i+1}$).
 This is essentially equivalent to obtaining gradients for the full ansatz by PSR, and updating all parameters simultaneously.
 However, we significantly reduce the resource requirements by delegating a portion of the gradient estimation task to classical hardware via $\mathfrak{g}$-sim .
-The required circuit evaluations per iteration is also the same as the above Alternate scheme, as the only difference comes in by using the non-updated unitary $U_{q}(\theta_{i})$ in the measurement step before $\mathfrak{g}$-sim .
+The required circuit evaluations per iteration is also the same as the above Alternate method, as the only difference comes in by using the non-updated unitary $U_{q}(\theta_{i})$ in the measurement step before $\mathfrak{g}$-sim .
+In practice, we find that using the Alternate method for a fixed number of iterations followed by Simultaneous training till convergence provides the most optimal results.
 1
-Input: $T_{max},\alpha,U_{q}(\mathbf{\theta_{0}}),U_{g}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
+Input: $T_{max},\eta,U_{q}(\mathbf{\theta_{0}}),U_{\mathfrak{g}}(\mathbf{\phi_{0}}),\mathfrak{g}_{DLA},H$
 Initialize: $\mathbf{\phi_{0}},\mathbf{\theta_{0}}\sim Norm(0,1);i=0$
-Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{g}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
+Define: $C(\theta,\phi)=\langle 0^{\otimes n}|U_{q}^{\dagger}(\theta)U_{g}^{\dagger}(\phi)HU_{\mathfrak{g}}(\phi)U_{q}(\theta)|0^{\otimes n}\rangle$
 begin
-       while $i\leq T_{s}$ do
+       while $i\leq T_{max}$ do
              PSR Step
 2             Obtain gradients $\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$, w.r.t $\theta_{i}$ via PSR
-3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\alpha\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
+3             $\mathbf{\theta}_{i+1}\leftarrow\theta_{i}-\eta\gradient_{\theta_{i}}C(\theta_{i},\phi_{i})$
              $\mathfrak{g}$-sim Step
@@
 Algorithm 2 Simultaneous Training for VQE
-In practice, we find the Alternate scheme for a fixed number of iterations followed by several iterations of the Simultaneous training provides the most optimal results.
-We will expand on this point in the subsequent sections.
+To evaluate the gradient for $U_{q}$, our approach is to implement the full circuit on quantum hardware for the shifted parameter values, as done in standard PSR.
+This requires $2p$ unique circuit evaluations.
+In another approach for gradient evaluation, one can prepare the states $|\psi(\theta_{i})\rangle=U_{q}(\theta_{i})|\psi\rangle$, as described in Sec.III.2.1, for each of the shifted values of the parameters $\theta_{i}$ followed by measuring the operators of the DLA.
+This will incur a total of $2p\times dim(\mathfrak{g})$ unique circuit evaluations, for $U_{q}$ with $p$ parameters and $U_{\mathfrak{g}}$ with $dim(\mathfrak{g})$ operators.
+Based on the lower unique circuit evaluations for the full circuit implementation, we use it for our numerical simulations.
+However in practice, one might have to make a choice between these two methodologies based on circuit depth and number of unique circuit evaluations.
 IV Results
-To evaluate the effectiveness of our method we employ several metrics,
-Error:
-$\text{Error}=\frac{E_{g}^{t}-E_{g}^{*}}{E_{g}^{*}}$ (13)
-where we define $E_{g}^{t}$ as the lowest possible energy achieved in the current trial and $E_{g}^{*}$ is an estimate of the true ground state energy of the Hamiltonian.
-For large Hamiltonians (beyond 14 qubits) it is computationally impractical to numerically diagonalize it and obtain the true ground state energy.
-Instead, taking inspiration from [31], we use the lowest energy reached across all trials for the corresponding Hamiltonian as an estimate of the true ground state energy $E_{g}^{*}$.
-Figure 6: XY Hamiltonian VQE (18 Qubits): The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G for 18 qubits.
-For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared.
-On the left, we show that both Alternate and Alternate+Simultaneous protocol has better success rates than PSR.
-For error (Column 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black).
-The Alternate method only performs well for Layer 1 configuration, which is improved by Alternate+Simultaneous, giving lower error than PSR for all layer configurations.
-Most importantly, on the right QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction).
+We demonstrate the effectiveness of our proposals on two main tasks without having any particular restriction on the DLA sizes of the operators involved: VQE and Classification using QNNs.
+In VQE task where the Hamiltonian has poly-DLA, we use HELIA and our proposed training methods.
+If one wishes to apply our methods, for the case of exponential-DLA Hamiltonian, our method cannot be applied directly owing to limitations of $\mathfrak{g}$-sim.
+Instead, we propose a pre-training procedure to leverage the benefits of our method.
+For Classification of quantum phases using QNNs, we consider the bond-alternating spin-$1/2$ Heisenberg chain which spans an exponential-DLA.
+To construct HELIA out of the given Hamiltonian, we choose multiple poly-DLA options for $U_{\mathfrak{g}}$ of the quantum circuit (as described in Sec.
+III.1).
+In these numerical experiments, our proposed methods shows improvement in accuracy.
+IV.1 Ground State Estimation
+We demonstrate the benefit of using HELIA as compared to $U_{q}$ and $U_{\mathfrak{g}}$ seperately through a simple example, illustrated in Fig.
+4.
+To find ground state of a 6-qubit Transverse Field Ising Model (TFIM) Hamiltonian using VQE, a quantum circuit with only $U_{q}$ (HEA via PSR) or $U_{\mathfrak{g}}$ (multi-qubit Pauli gates via $\mathfrak{g}$-sim ) fails to reach the target energy in 500 iterations.
+However, combining the two blocks using an Alternate training method (discussed in Sec.
+III.2) successfully converges to the exact ground state energy.
+Although the combined ansatz has more trainable parameters than either of $U_{q}$ and $U_{\mathfrak{g}}$ individually, the quantum resources required to train the full ansatz is reduced significantly, compared to standard PSR, by using our training methods.
+Figure 4: VQE performed for a 6-qubit TFIM Hamiltonian.
+Only using HEA with PSR (shown in orange) or multi-qubit Pauli rotation gates with $\mathfrak{g}$-sim (shown in blue) is not enough to reach exact ground state (shown in red).
+But combining both methods into an Alternate training procedure (in green) allows us to reach the correct solution.
+To rigorously test the performance of our proposal, we employ several metrics.
+Relative Error:
+[…]
+Here, we define $E_{g}^{t}$ as the lowest possible energy achieved in the current trial and $E_{g}^{*}$ is an estimate of the true ground state energy of the Hamiltonian.
+For large Hamiltonians, it is computationally impractical to numerically diagonalize it and obtain the true ground state energy.
+Instead, taking inspiration from Ref.
+[44], we use the lowest energy reached across all trials for the corresponding Hamiltonian as an estimate of the true ground state energy $E_{g}^{*}$.
+Figure 5: XY Hamiltonian VQE (1 YZ linear layer): The success (Row 1), relative error (Row 2) and QPU calls reduction (Row 3) are plotted for configurations with 1 layer of YZ linear ansatz in $U_{q}$ and Hamiltonian DLA gates in $U_{\mathfrak{g}}$ for 6 to 18 qubits.
+For each qubit count and metric, PSR (orange), Alternate (Alt, blue) and Alternate+Simultaneous (Alt+Sim, purple) training methods are compared.
+On the first row, we show that both Alt and Alt+Sim methods have better success rates than PSR.
+For relative error (Row 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black).
+The Alt method performs well for Layer 1 configuration, which is improved by Alt+Sim, giving lower relative error than PSR for all qubits.
+Most importantly, on the lowermost row we plot the percentage of reduction in QPU calls for Alt and Alt+Sim compared to PSR.
 The spread in standard deviation is shown in black.
-We consistently see reduction in QPU calls using our methods, with frequently lower error values and higher success metrics as compared to PSR.
-Figure 7: XY Hamiltonian VQE: The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G in each row, for qubit sizes ranging from 6 to 16.
-For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared.
-We consistently see reduction in QPU calls using our methods at high qubits, with frequently lower error values and higher success metrics as compared to PSR.
-Figure 8: Gradients for XY Hamiltonian VQE: Variance of the first parameter from PSR and $\mathfrak{g}$-sim block are plotted with increasing qubits for both Full-PSR (blue) and our Alternate method (orange) for finding ground state of XY Hamiltonian using VQE.
-The diagrams also show the plots for varying number of PSR block layers (1,3,6,9 from top to bottom).
+Except for 6 qubit case, we consistently see reduction in QPU calls using our methods, with lower relative error values and higher success metrics as compared to PSR.
+Figure 6: Gradients for XY Hamiltonian VQE: Variance of the first parameter from $U_{q}$ (red) and $U_{g}$ (blue) are plotted with increasing qubits for both Full-PSR and our Alternate method for finding ground state of XY Hamiltonian using VQE.
+The diagrams show the plots for 1 YZ linear layer in $U_{q}$.
 For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green.
-The green curve is the same in all of the plots and provided as a reference for our improved gradients.
-In both $\mathfrak{g}$-sim and PSR block, we notice a slower decay of gradients as compared to HEA50, with the PSR block maintaining slightly higher magnitude than the $\mathfrak{g}$-sim block.
-Overall the plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
-Success: Fraction of the trials that are able to reach below a error threshold of $10^{-3}$, which has been chosen ad-hoc.
+In both $U_{q}$ and $U_{\mathfrak{g}}$, we notice a slower decay of gradients as compared to HEA50.
+Overall the plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even for higher number of qubits in both of the blocks.
+Success: Fraction of the trials that are able to reach below a relative error threshold of $10^{-3}$, which has been chosen ad-hoc.
 This gives an estimate of how often can the method get close to the true solution.
-Going forward, we report the relevant statistics only for the successful trials, except for $\mathfrak{g}$-sim as the error for this case in the mentioned examples is always higher than the threshold.
-QPU Calls Reduction: Fraction of quantum circuit evaluations, referred to as Quantum Processing Unit (QPU) calls, reduced in current method as compared to standard training with PSR.
+We report the relevant statistics only for the successful trials, except for $\mathfrak{g}$-sim as the relative error for this case in the mentioned examples is always higher than the threshold.
+QPU Calls Reduction: Fraction of quantum circuit evaluations, referred to as QPU calls, reduced in current method as compared to standard training with PSR.
 We use the number of QPU calls required to reach the point of lowest energy in the corresponding trial as our estimate.
-IV.1 Ground State Estimation using VQE
-The proposed protocols can be efficiently used for VQE when the Hamiltonian spans a poly-DLA as shown in Sec.
-IV.1.1.
-However, for more general cases, where the Hamiltonian spans an exponential DLA, we can still use our method as a pre-training procedure.
-We discuss this in detail for Longitudinal Transverse Field Ising Model (LTFIM) in Sec.
-IV.1.2.
 IV.1.1 Improved VQE for XY Spin Hamiltonian (poly-DLA)
-We apply our chosen ansatz and proposed training schemes on VQE, and demonstrate its effectiveness for this task in comparison to PSR and $\mathfrak{g}$-sim .
-For this, we consider a XY Spin Hamiltonian given by,
-$H_{XY}=\sum_{i}^{N-1}\alpha_{i}\hat{X}_{i}\hat{X}_{i+1}+\beta_{i}\hat{Y}_{i}\hat{Y}_{i+1}$ (14)
-XY Spin Hamiltonian has a poly-DLA with the dimension given by $dim(\mathfrak{g})=n^{2}-n$. This is necessary to use $\mathfrak{g}$-sim effectively in our algorithm.
-While running the $\mathfrak{g}$-sim -only experiments, all of the DLA elements of the XY spin Hamiltonian produce zero expectation value with the zero initial state $|0^{\otimes n}\rangle$, we added a layer of Hadamard gates $H^{\otimes n}$ to facilitate training .
-In order to maintain consistency we have added the $H^{\otimes n}$ layer before Block G in all of the examples for the Hamiltonian.
+We apply our chosen ansatz and proposed training methods on VQE, and demonstrate its effectiveness for this task in comparison to PSR and $\mathfrak{g}$-sim.
+We consider a XY Spin Hamiltonian given by:
+$H_{XY}=\sum_{i=0}^{N-1}\alpha_{i}\hat{X}_{i}\hat{X}_{i+1}+\beta_{i}\hat{Y}_{i}\hat{Y}_{i+1},$ (8)
+which has a poly-DLA with the dimension given by $dim(\mathfrak{g})=n^{2}-n$.
+Since all of DLA elements of the XY spin Hamiltonian produce zero expectation value with the zero initial state $|0^{\otimes n}\rangle$, we added a layer of Hadamard gates $H^{\otimes n}$ to facilitate training .
+In order to maintain consistency we have added the $H^{\otimes n}$ layer before $U_{\mathfrak{g}}$ in all of the examples for the Hamiltonian.
 Hence, via $\mathfrak{g}$-sim we optimize the following cost function,
-$C(\phi)=\langle 0|H^{\otimes n}U_{g}^{\dagger}(\phi)H_{XY}U_{g}(\phi)H^{\otimes n}|0\rangle$ (15)
-For the Full-PSR (all parameters of Block Q & G optimized using PSR), Alternate and Alternate + Simultaneous columns we optimize using a larger ansatz by adding the Block Q, and obtain the following cost function,
-$C(\theta,\phi)=\langle 0|U_{q}^{\dagger}(\theta)H^{\otimes n}U_{g}^{\dagger}(\phi)H_{XY}U_{g}(\phi)H^{\otimes n}U_{q}(\theta)|0\rangle$ (16)
-Now we are not restricted to only searching in the same polynomial space but can search a much larger state space, by virtue of the gates in Block Q.
-For visualization, we plot the success, error and QPU reduction metrics comparing Full-PSR, Alternate and Alternate + Simultaneous in Fig.
-6 and 7, while the actual numbers along with comparison with $\mathfrak{g}$-sim is provided in Table 4 and 3 in Appendix A.
-The error for $\mathfrak{g}$-sim is significantly higher in the chosen examples, as compared to other methods.
+$C(\phi)=\langle 0|H^{\otimes n}U_{\mathfrak{g}}^{\dagger}(\phi)H_{XY}U_{\mathfrak{g}}(\phi)H^{\otimes n}|0\rangle.$ (9)
+For the Full-PSR (all parameters of $U_{q}$ and $U_{\mathfrak{g}}$ are optimized using PSR), Alternate (Alt) and Alternate + Simultaneous (Alt+Sim, using Alternate training followed by Simultaneous training) methods, we consider a larger ansatz by adding $U_{q}$, and minimize the following cost function:
+$C(\theta,\phi)=\langle 0|U_{q}^{\dagger}(\theta)H^{\otimes n}U_{\mathfrak{g}}^{\dagger}(\phi)H_{XY}U_{\mathfrak{g}}(\phi)H^{\otimes n}U_{q}(\theta)|0\rangle.$ (10)
+Now we are not restricted to only searching in the same polynomial space but can search a much larger state space, by virtue of the gates in $U_{q}$.
+For visualization, we plot the success, relative error and QPU reduction metrics comparing Full-PSR, Alt and Alt+Sim using 1 YZ linear layer in $U_{q}$ in Fig.
+5.
+The plots for other configurations are provided in Fig.
+10 of Appendix D, while the actual numbers along with comparison with $\mathfrak{g}$-sim is provided in Table 4 and 5.
+The relative error for $\mathfrak{g}$-sim is significantly higher in the chosen examples, as compared to other methods.
 Hence, we ignore it in the plots, and only mention the values in Table 4.
-Later in IV.1.2, we also discuss how this method can be used as a starting point for finding ground states in more general Hamiltonians without the restriction of having a poly-DLA.
-Although, we are exploring a larger state space, often the ground state might infact lie close to the space searched by $\mathfrak{g}$-sim , rendering a quantum method unnecessary for the task.
-This infact happens in the odd qubit examples for XY Hamiltonian, where $\mathfrak{g}$-sim is able get very low errors without additional gates.
+In Sec.
+IV.1.2, we also discuss how this method can be used as a starting point for finding ground states in more general Hamiltonians without the restriction of having a poly-DLA.
+Although we are exploring a larger state space, the ground state might lie close to the space searched by $\mathfrak{g}$-sim, rendering a quantum method unnecessary for the task.
+This happens in the odd qubit examples for XY Hamiltonian, where $\mathfrak{g}$-sim is able to get very low relative errors without additional gates.
 We mention this here, but in the numerics we only focus on even qubit cases to demonstrate where our method can be effective and standard $\mathfrak{g}$-sim fails.
-Some of the odd qubit examples are demonstrated in Appendix.
-B.
-Experiment Details: We present the results for 4 different configurations for each qubit, with the number of YZ linear layers: 1,3,6 and 9, changing between the configurations.
-The results are given in Table 2 and 3 of Appendix A.
-The Full-PSR column refers to training the full ansatz using standard PSR, Alternate refers to training it by the Alternate training scheme till convergence and Alternate + Simultaneous trains the ansatz using Alternate Scheme for 500 iterations followed by Simultaneous training till convergence.
+Some of the odd qubit examples are demonstrated in Appendix G.
+Experiment Details: We present the results for 4 different configurations for each qubit, with the number of YZ linear layers in $U_{q}$: 1, 3, 6 and 9 changing between the configurations.
+The results are given in Table 3 and 4 of Appendix D, and visualized in Fig.
+5 for layer 1 and Fig.
+10 of Appendix D for layer 3, 6 and 9.
+Here, in the Alt+Sim method, we train the ansatz using Alternate for 500 iterations followed by Simultaneous till convergence.
 In order to keep the comparison fair, we use an Adam optimizer with initial learning rate $0.01$ and keep it consistent across all experiments.
 In Fig.
-6 and 7, we have plotted the mean and standard deviation (in black) over the successful trials when measuring the QPU calls reduction.
-For error, we find the data is often asymmetrically spread around the mean, owing to being lower-bounded by 00 (by definition Eq.(13)) and upper bounded by $10^{-3}$ (success threshold).
-Hence, for this we resort to the median, and show the spread from $25\%$ to $75\%$ quantiles in black instead of mean and standard deviation.
+5, we plot the mean and standard deviation (in black) over the successful trials when measuring the QPU calls reduction for the configuration with 1 YZ linear layer in $U_{q}$.
+We find the relative error is often asymmetrically spread around the mean, owing to being lower-bounded by 00 (by definition Eq.
+(7)) and upper bounded by $10^{-3}$ (success threshold).
+Hence, we resort to the median and show the spread from $25\%$ to $75\%$ quantiles in black instead of mean and standard deviation.
 This is due to the fact that median is a more robust metric when the data is distributed asymmetrically.
 This unequal distribution is also explicit from the $25\%$ to $75\%$ quantiles which extend unequally from the median.
+The corresponding plots for 3,6 and 9 YZ linear layers in $U_{q}$ are provided in Appendix D (Fig.
+10).
 Mitigating Barren Plateaus
-We also analyze the gradients of our ansatz for the given VQE task with varying number of qubits.
-In each case, we track the gradient of the first parameter from each of the Block Q (referred as PSR block) and Block G (referred as $\mathfrak{g}$-sim block) and plots its variance during the training (Fig.8).
+We analyze the parameter gradients for the given VQE task with varying number of qubits.
+We track the gradient of the first parameter from each of $U_{q}$ and $U_{\mathfrak{g}}$, and plots its variance during the training (Fig.6).
 Each datapoint is obtained after evaluating the variance from the full training of 64 independent trials.
 For comparison, we also perform the same VQE task using a 50 layer HEA (referred as HEA50), which is known to exhibit BP [95], and plot the variance of the first parameter of this circuit.
-In other words, while it is possible that our method will also run into the BP issue at a certain qubit count, atleast for the system sizes we have tested so far, the method holds up well against BP issues in previous methods.
-Secondly, we believe the reason the direct BP condition in [82, 83] is not applicable here because we choose gates from a poly-DLA and a shallow HEA circuit preventing us from forming a 2-design.
-However, we do not provide proof of this statement, and propose this as only a potential reason.
-HELIA shows a clear improvement in terms of slow decay in gradients in Fig 8.
-The gradients from $\mathfrak{g}$-sim block demonstrate an exponential decrease, but with much less slope as compared to the HEA50 circuit, while the PSR block shows a sub-exponential decay.
+HELIA shows a clear improvement in terms of slow decay in gradients in Fig.
+6 for the configuration with 1 YZ linear layer in $U_{q}$.
+The gradients for $U_{q}$ and $U_{\mathfrak{g}}$ demonstrate a much slower decay compared to the HEA50 circuit.
+Similar trend was observed for increasing YZ linear layers up to 9, as shown in Fig.
+11.
 Overall, this allows for larger qubit models to be trained efficiently without running into vanishingly small gradients.
-IV.1.2 Pre-training in LTFIM Hamiltonian (exp-DLA)
-The ability to combine $\mathfrak{g}$-sim and PSR efficiently in our protocol depends on the poly-DLA of the $\mathfrak{g}$-sim block as well as the associated measurement operator.
-This limits the kind of Hamiltonians on which we can perform VQE task.
-However, we can still use our method as a starting point for general Hamiltonians with exponential DLA as demonstrated in Algorithm 4.
+The BP condition in Refs.
+[96, 94] is not applicable here is because we choose gates from a poly-DLA and a shallow HEA circuit preventing us from forming a 2-design on $SU(2^{n})$.
+To understand this further, we consider the variance of parameters of $U_{\mathfrak{g}}$.
+Since we do not require shallow circuits for this block, we can focus on the case where the gates form a 2-design on $e^{i\mathfrak{g}}$ (not on full $SU(2^{n})$).
+Using the result from Ref.
+[96] and considering $\mathfrak{g}$ to be simple, we can write the following,
+$\text{Var}_{\phi}[C(\theta,\phi)]=\frac{\mathcal{P}_{\mathfrak{g}}(\rho_{q}(\theta))\mathcal{P}_{\mathfrak{g}}(\hat{O})}{dim(\mathfrak{g})},$ (11)
+where $\rho_{q}(\theta)=\tilde{U}_{q}(\theta)\rho_{0}\tilde{U}_{q}^{\dagger}(\theta)$.
+$\mathcal{P_{\mathfrak{g}}(\cdot)}$ denotes the $\mathfrak{g}$-Purity of an operator and $dim(\cdot)$ denotes the dimension of an algebra, as defined in [96].
+We elaborate on these functions in Appendix F.
+By design of HELIA, $\mathcal{P}_{\mathfrak{g}}(\hat{O})=1$ and $dim(\mathfrak{g})=poly(n)$.
+Hence we only need to ensure that $\mathcal{P}_{\mathfrak{g}}(\rho_{q}(\theta))$ is not decaying exponentially.
+By numerically testing at varying qubit counts using $\mathfrak{g}_{XY}$, we find that $\mathcal{P}_{\mathfrak{g}}(\rho_{q}(\theta))$ decays sub-exponentially for constant and logarithmic depth circuits, and exponentially for linear depth circuits.
+The details of the experiment are shown in Appendix F.
+Using Eq.(11), we can see that sub-exponentially decay of average purity implies that the variance of loss function with respect to $U_{\mathfrak{g}}$ parameters also decays sub-exponentially.
+Importantly, this does not guarantee substantial gradients for every $\theta$ value in $U_{q}$, but only on an average.
+As shown in Fig.
+6, numerical tests on the full circuit created by composing $U_{q}$ and $U_{\mathfrak{g}}$, shows polynomial decay of gradients in both.
+However, here we only provide an outline of proof for $U_{\mathfrak{g}}$, and show that polynomial gradients can be seen for sub-linear circuit depth of block $U_{q}$.
+We believe that the polynomial decay in $U_{q}$ also stems from the shallowness of the HEA circuit, and might be preserved for sub-linear circuit depths.
+We leave the mathematical proof for this intuition as future work.
+IV.1.2 Pre-training in LTFIM Hamiltonian (exponential-DLA)
+The ability to combine $\mathfrak{g}$-sim and PSR efficiently in our proposal depends on the poly-DLA of $U_{\mathfrak{g}}$ as well as the associated measurement operator.
+This limits the form of Hamiltonians on which we can perform VQE task.
+However, we can still use our method as a starting point for general Hamiltonians with exponential-DLA as demonstrated in Appendix I, Algorithm 4.
 Although similar to the pre-training approach used in Ref.
@@
 We expand this point further in the following paragraphs and highlight this important difference through a numerical example.
-To start the protocol, we need a step of choosing a subset of operators that form a poly-DLA.
+To start the pre-training method, we need to choose a subset of operators that form a poly-DLA.
 In poly-DLA phase, we perform Alternate, followed by Simultaneous training on the reduced Hamiltonian.
-During the Exponential DLA Phase of training in Alg.4, we train only the PSR block for some iterations before training the full block.
-This is done with the aim of keeping the overall quantum resources low, as the PSR block has fewer trainable parameters and hence needs less circuit evaluations to obtain gradients during training.
+During the exponential-DLA phase of training in Alg.
+4 (Appendix  I), we train only $U_{q}$ for some iterations before training the full block.
+This is done with the aim of keeping the overall quantum resources low, as $U_{q}$ has fewer trainable parameters and hence needs less circuit evaluations to obtain gradients during training.
 Finally, all of the parameters are trained together using standard PSR.
-We demonstrate the results of this protocol for longitudinal-transverse field Ising model (LTFIM),
+The experimental details are provided in Appendix I.
+We demonstrate the results of this proposal for longitudinal-transverse field Ising model (LTFIM),
 $H_{LTFIM}=\sum_{j}\alpha_{j}X_{j}X_{j+1}+\sum_{j}(\beta_{j}Z_{j}+\gamma_{j}X_{j})$ (12)
-which has an exponential DLA. A poly-DLA can be constructed by dropping the longitudinal terms
-$i\mathfrak{g}_{TFIM}=\langle\{iX_{j}X_{j+1},iZ_{j}\}\rangle_{Lie}$ (18)
-(a) LTFIM Hamiltonian with 10 qubits using PSR
-(b) LTFIM Hamiltonian with 10 qubits using our proposed method
-(c) LTFIM Hamiltonian with 12 qubits using PSR
-(d) LTFIM Hamiltonian with 12 qubits using proposed method
-Figure 9: Demonstration of PSR (9(a) and 9(c)) and our method (9(b) and 9(d)) to run VQE for LTFIM Hamiltonian (with an exponential DLA) in 10 & 12 qubit cases.
-The purple and red regions show Alternate and Simultaneous training respectively for the restricted Hamiltonian, followed by training the PSR block and full circuit in green and yellow regions respectively for the full LTFIM Hamiltonian.
-In both 10 & 12 qubit examples with 3 & 6 YZ linear configurations in the orange block, we notice that our training methods in the initial phase allows VQE to reach much closer to the target ground state during the final training, as compared to just using PSR for the full training, for similar number of of total training iterations.
-For the overall training, we also slightly reduce the QPU calls ($\sim 6-10\%$)
-Experiment Details: The PSR block is composed of varying layers of YZ linear ansatz (3 and 6 layers), while the $\mathfrak{g}$-sim block is generated out of gates from $\mathfrak{g}_{TFIM}$.
-In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of the PSR block and 1000 iterations of full circuit training.
-Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods.
-However, these choices are made ad-hoc and can be modified based on other stopping conditions, such as Early Stopping, that can further reduce QPU calls.
-Input: $T_{alt},T_{sim},T_{psr},T_{full},U_{q}(\mathbf{\theta_{0}}),H=\sum_{i}\beta_{i}P_{i}$
-Output: $\theta^{*},\phi^{*}$
-Initialize: $\vec{\theta}_{0}\sim\mathcal{N}(0,1);i=0$
-begin
-       poly-DLA Phase
-1       Obtain a subset of operators $\{P_{j}\}_{j\in\mathcal{A}}$ from $H$ such that it forms a poly-DLA $\mathfrak{g}$.
-2       Define $H_{poly}=\sum_{P_{i}\in\mathcal{A}}\beta_{i}P_{i}$, and unitary $U_{g}(\vec{\phi_{0}})=e^{-iP_{1}\phi_{1}^{0}}\ldots e^{-iP_{n}\phi_{n}^{0}}$
-3       Randomly initialize $\vec{\phi}_{0}\sim\mathcal{N}(0,1)$
-4       Define the cost function: $C_{poly}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})H_{poly}U_{g}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$
-5       Run Alternate optimization given in Alg.2 for $T_{alt}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
-6       Run Simultaneous optimization given in Alg.3 for $T_{sim}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
-7       Extract the optimized parameters $(\vec{\theta}^{*},\vec{\phi}^{*})$
-       Exponential DLA Phase
-8       Define the full cost function, $C_{full}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})HU_{g}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$ initialized at $(\vec{\theta}^{*},\vec{\phi}^{*})$
-9       Optimize only parameters $\vec{\theta}$ of $U_{q}(\vec{\theta})$ in $C_{full}(\vec{\theta},\vec{\phi})$ for $T_{psr}$ iterations using PSR
-10       Optimize full cost function $C_{full}(\vec{\theta},\vec{\phi})$ for $T_{full}$ using PSR
-11       Extract optimized parameters at the end of training $(\vec{\theta}^{*},\vec{\phi}^{*})$
-12
-Algorithm 4 VQE for General Hamiltonian
-We plot the error metric in Eq.(13) for the LTFIM Hamiltonian with 10 & 12 qubits in Fig.9.
+which has an exponential-DLA. One choice of poly-DLA can be constructed by dropping the longitudinal terms
+$i\mathfrak{g}_{TFIM}=\langle\{iX_{j}X_{j+1},iZ_{j}\}\rangle_{Lie}.$ (13)
+(a) Trained using standard PSR
+(b) Pre-training using our method followed by PSR
+Figure 7: Demonstration of PSR (a) and our method (b) to run VQE for 12 qubit LTFIM Hamiltonian (with an exponential-DLA) with 3 YZ linear layers in $U_{q}$.
+The orange and green regions show Alternate and Simultaneous training respectively for the restricted Hamiltonian, followed by training $U_{q}$ and full circuit in blue and pink regions respectively for the full LTFIM Hamiltonian.
+We notice that our training methods in the initial phase allows VQE to reach much closer to the target ground state during the final training, as compared to just using PSR for the full training, for similar number of of total training iterations.
+For the overall training, we also reduce the QPU calls by  $10.7\%$.
+We plot the relative error metric in Eq.
+(7) for the 12 qubit LTFIM Hamiltonian with 3 YZ linear layers in $U_{q}$ in Fig.
+7.
 Importantly, although the initial training phase is based on the reduced Hamiltonian, the plotted energy values correspond to expectation values of the full Hamiltonian.
-As shown in Fig.9, the algorithm is indeed able to converge to good solutions by using the our protocols in the first phase of the training.
-Compared with PSR, our protocol shows lower error value on an average in Fig.
-IV.1.2 for both 10 & 12 qubits.
-More importantly, the QPU calls required for the 10-qubit Hamiltonian VQE using PSR is $54.5e+$ 5 and $56.3e+$ 5 for Layer 3 and 5 respectively, while our proposed method requires $54.065e+$ 5 ($\sim 9.6\%$ reduction) and $55.925e+$ 5 ($\sim 6.5\%$ reduction) QPU calls respectively.
-Similarly, for 12 qubit, PSR uses $56.12e+$ 5 and $58.28e+$ 5 for Layer 3 and 5 respectively while our method requires $55.466e+$ 5 ($\sim 10.7\%$ reduction) and $57.698e+$ 5 ($\sim 7\%$ reduction) respectively.
-Overall, our method shows lower error while reducing the required quantum resources.
+As shown in Fig.
+7, the algorithm is indeed able to converge to good solutions by using our methods in the first phase of the training.
+Compared with PSR, our proposal shows lower relative error value on an average.
+The QPU calls required for PSR is $56.12e+$ 5 while our method requires $55.466e+$ 5 ($\sim 10.7\%$ reduction) respectively.
+Overall, our method shows lower relative error while reducing the required quantum resources for a variational task involving operators with exp-DLA.
+The corresponding plots for other configuration and qubit counts are shown in Appendix I (Fig.
+19).
+IV.1.3 Pre-training in LiH Hamiltonian (exponential-DLA)
+The pre-training method is useful for ground-state estimation of molecules as well, where the Hamiltonian can span an exponential-DLA. To demonstrate that, we consider the example of LiH molecule.
+The electronic structure of the LiH molecule was modeled using the Slater-type orbital (STO-3G) minimal basis set, resulting in a second-quantized fermionic Hamiltonian [97, 98].
+This Hamiltonian was then mapped to a qubit representation using the Jordan–Wigner transformation implemented using OpenFermion [99, 100].
+Without applying any orbital freezing or symmetry-based reductions, the full system comprises 12 spin orbitals, corresponding to a 12-qubit Hamiltonian.
+The resulting qubit Hamiltonian consists of a linear combination of 631 Pauli strings, each expressed as a tensor product over the Pauli operators ${I,X,Y,Z}$.
+An interatomic distance of $1.5~{}\text{\AA}$ is considered for the application of the VQE algorithm to this Hamiltonian.
+This uncompressed configuration serves as a benchmark for evaluating both the standard VQE approach and the proposed method.
+Since we are considering a chosen fixed Hamiltonian, we first need to decide what portion of the computation can be delegated to classical hardware.
+For this we choose the threshold of $n^{2}$ for an $n$-qubit Hamiltonian, to be the maximum allowed DLA size.
+This ensures we always use a reasonable amount of classical resource, and it does not become exponentially large in the number of qubits.
+Next, we arrange the Hamiltonian terms in descending order of magnitude and choose the first $k$ operators which have a DLA size below our mentioned threshold.
+In our case, this creates a DLA of size 78.
+Ignoring the identity, the operators which are considered within the DLA comprise of $99.47\%$ of the magnitude of the full Hamiltonian.
+These operators form the generators for $U_{\mathfrak{g}}$ in the pre-training method shown in the previous example.
+In practice, it might be possible to increase the DLA size further based on available classical hardware.
+Full PSR Our Proposal
+Relative Error 0.01090 0.0091
+QPU Reduction - 9.09%
+Table 1: Median Relative Error and QPU Reduction for Full-PSR vs our proposal
+We compare Full PSR and our proposal in Table 1 in terms of median relative error and total QPU calls required in the corresponding methods.
+In the above experiment, we are able to reduce relative error by 16.51% percent and QPU calls by 9.09% percent.
+The details of the experiment are provided in Appendix J.
 IV.2 Classification of Quantum Phases
-Figure 10: Classification Workflow: $\rho_{i}$ represents the initial state obtained from a training dataset, which is processed using $U_{q}(\theta)$ and $U_{g}(\phi)$ as described before.
-The measurement operator belongs to a chosen DLA, which also generates the gates for $U_{g}(\phi)$
-Our training protocols are useful for QML tasks such as classification. To illustrate this we choose the bond-alternating spin-$1/2$ Heisenberg chain [114],
-$H=J\sum_{i}\vec{S}_{2i-1}\cdot\vec{S}_{2i}+J^{\prime}\sum_{i}\vec{S}_{2i}\cdot\vec{S}_{2i+1}$ (19)
+Our training methods are also useful for QML tasks such as classification. For numerical examples, we consider the bond-alternating spin-$1/2$ Heisenberg chain [93]:
+$H=J\sum_{i=0}\vec{S}_{2i-1}\cdot\vec{S}_{2i}+J^{\prime}\sum_{i=0}\vec{S}_{2i}\cdot\vec{S}_{2i+1},$ (14)
 where $\vec{S}=(X,Y,Z)$ are the Pauli matrices. The model undergoes a second-order phase transition at $J/J^{\prime}=1$ [93].
-We pose the task of learning the quantum phase of the ground state as a classification problem for 10 & 12 qubit examples.
-Since, quantum phase transition only happens in the limit $N\rightarrow\infty$, it is not accurate to refer to these states as different phases.
+We consider the task of learning the quantum phase of the ground state as a classification problem for a 12 qubit case.
+Since quantum phase transition only happens in the limit $N\rightarrow\infty$, it is not accurate to refer to these states as different phases.
 Rather, we just assign two separate labels ±$\pm 1$ to the ground states for $J<J^{\prime}$ and $J>J^{\prime}$.
-Since the qubit counts are small enough, the training and test data (100 train + 100 test dataset) are generated by sampling $J,J^{\prime}$ uniformly from $[-1,1]$ numerically diagonalizing the Hamiltonian in Eq.(19) and obtaining the ground state.
-This ground state is used as an input into the PQC and the ansatz is trained to generate the correct labels after measuring a Pauli string, that is randomly chosen and fixed at the beginning of the experiment.
-For designing the PQC, we encounter the choice of gates for PSR and $\mathfrak{g}$-sim blocks.
-For the PSR block, we stick to the YZ linear ansatz choice with 9 layers.
-For $\mathfrak{g}$-sim block however, we do not have a unique choice like we had for previous examples.
-Hence, we explore 3 different DLA choices to construct the gates:
+As the qubit counts are small enough, the training and test data (100 train + 100 test dataset) are generated by sampling $J,J^{\prime}$ uniformly from $[-1,1]$, numerically diagonalizing the Hamiltonian in Eq.
+(14) and obtaining the ground state.
+This ground state is used as an initial state in the PQC and the ansatz is trained to generate the correct labels.
+For designing the PQC, we encounter the choice of gates for $U_{q}$ and $U_{\mathfrak{g}}$.
+For $U_{q}$, we stick to the YZ linear ansatz choice with 9 layers.
+For $U_{\mathfrak{g}}$ however, we do not have a unique choice like we had for previous examples.
+Hence, we explore three different DLA choices to construct the gates:
 $i\mathfrak{g}_{XY}$ =$=\langle\{iX_{j}X_{j+1},iY_{j}Y_{j+1}\}\rangle_{Lie}$ (15)
 $i\mathfrak{g}_{YZ}$ =$=\langle\{iZ_{j}Z_{j+1},iY_{j}Y_{j+1}\}\rangle_{Lie}$ (16)
-$i\mathfrak{g}_{ZX}$ =$=\langle\{iX_{j}X_{j+1},iZ_{j}Z_{j+1}\}\rangle_{Lie}$ (22)
-For each of these configurations of DLA, we run 10 independent trials and compare test accuracy. The results are plotted in Fig.1.
-We observe different peak accuracies for each of the DLA choices with $i\mathfrak{g}_{YZ}$ providing the highest accuracy in $\mathfrak{g}$-sim for 10 qubits ($0.854\pm 0.008$) and $i\mathfrak{g}_{ZX}$ providing the highest accuracy in $\mathfrak{g}$-sim for 12 qubits ($0.864\pm 0.013$) .
-In each case, the Alternate and Full-PSR method provide a higher peak accuracy on average, although only by $\sim 1-6\%$.
-However, the Alternate Method always requires much fewer QPU calls to reach peak accuracy as compared to Full-PSR as shown in Table 1.
-In other words, almost same accuracy is reached with our Alternate Method but using much lesser QPU calls compared to only the PSR method, and with better accuracy compared to the $\mathfrak{g}$-sim method.
-DLA Full PSR g-sim Alternate
-XY 0.898 $\pm$ 0.009 0.860 $\pm$ 0.004 0.897 $\pm$ 0.018
-YZ 0.911 $\pm$ 0.008 0.853 $\pm$ 0.005 0.907 $\pm$ 0.013
-ZX 0.905 $\pm$ 0.007 0.864 $\pm$ 0.013 0.908 $\pm$ 0.007
-(a) Test Accuracy across different methods
-DLA Full PSR g-sim Alternate
-XY 36736 $\pm$ 30481 - 13381 $\pm$ 6143
-YZ 32800 $\pm$ 28216 - 7080 $\pm$ 3540
-ZX 25256 $\pm$ 17172 - 11894 $\pm$ 9386
-(b) QPU Calls across different methods
-Table 1: Classification task using $\mathfrak{g}$-sim , Alternate and Full-PSR Training (with 9 layers in PSR block) for 12 qubits.
-Three different choices of poly-DLA (referred to as $\mathfrak{g}_{XY},\mathfrak{g}_{YZ}$ and $\mathfrak{g}_{ZX}$) are explored for $\mathfrak{g}$-sim , and also to construct the Block G for Alternate and Full-PSR Method.
-Average peak test accuracy and number of QPU calls required to reach peak accuracy is reported for each of the configuration.
-For all examples, the Alternate improves over the accuracy of $\mathfrak{g}$-sim , and is similar to the Full-PSR training with dramatically reduced QPU calls
+$i\mathfrak{g}_{ZX}$ =$=\langle\{iX_{j}X_{j+1},iZ_{j}Z_{j+1}\}\rangle_{Lie}.$ (17)
+Interestingly, the Hamiltonian is symmetric under the group action that maps $X_{i},Y_{i},Z_{i}$ Paulis cyclically into each other, which also maps the above mentioned DLA choices into each other.
+However, in practice we notice that choice of DLA still impacts the peak test accuracies.
+The details of the experiment are provided in Appendix J.
+The results are tabulated in Table 2.
+We observe different peak accuracies for each of the DLA choices with $i\mathfrak{g}_{ZX}$ providing the highest accuracy in $\mathfrak{g}$-sim ($0.873\pm 0.020$) .
+In each case, the Alternate and Full-PSR method provide a higher peak accuracy on average than $\mathfrak{g}$-sim .
+Furthermore, Alt+Sim is able to reach higher test accuracies than Full-PSR, although only by up to $2.8\%$.
+DLA g-sim Full PSR Alt+Sim
+XY 0.863 $\pm$ 0.005 0.884 $\pm$ 0.016 0.888 $\pm$ 0.021
+YZ 0.858 $\pm$ 0.006 0.887 $\pm$ 0.025 0.889 $\pm$ 0.020
+ZX 0.873 $\pm$ 0.020 0.883 $\pm$ 0.015 0.911 $\pm$ 0.008
+Table 2: Classification task using $\mathfrak{g}$-sim, Full-PSR and Alt+Sim training (with 9 layers in $U_{q}$) for 12 qubits.
+Three different choices of poly-DLA (referred to as $\mathfrak{g}_{XY},\mathfrak{g}_{YZ}$ and $\mathfrak{g}_{ZX}$) are explored for $\mathfrak{g}$-sim, and also to construct $U_{\mathfrak{g}}$ for Alternate and Full-PSR method.
+Average peak test accuracy is reported for each of the configuration.
+For all examples, the Alt+Sim method improves over the accuracy of $\mathfrak{g}$-sim and Full-PSR by up to 3.8% and 2.8% respectively.
 V Conclusion and Outlook
-In order to reach quantum advantage, we need to deploy larger and more complex algorithms on quantum computers while ensuring classical intractability.
-The reason this is critical for variational algorithms is that, since these are algorithms are heuristic in nature, identifying the thresholds for quantum utility or advantage is usually hard.
-Therefore, the rather seemingly pedestrian approach of deploying larger algorithms in practice and testing them rigorously, may actually turn out to produce valuable insights.
+Achieving quantum advantage requires deploying larger, more complex algorithms on quantum computers while ensuring classical intractability.
+For VQAs, which are inherently heuristic, determining thresholds for quantum utility or advantage is challenging.
+Consequently, a practical approach of implementing and rigorously testing larger-scale algorithms may yield critical insights into their potential for quantum advantage.
 In this work, we provide a step in that direction by an informed ansatz design and combining $\mathfrak{g}$-sim with PSR, thus delegating the gradient estimation task to both classical and quantum hardware carefully.
-We extensively test our novel protocols on several practical tasks like VQE and Classification for qubit counts ranging from 6 to 18.
-In our examples, we notice significant reduction in QPU calls (even upto 60% for the case of VQE in 16 qubit XY Hamiltonian with 1 YZ linear layer), especially at higher number of qubits, alongside improvement in error and success metrics for VQE and accuracy for Classification.
-Importantly, we also notice a slow decay in gradients as we scale up in qubits, providing a potential strategy in mitigating the ominous Barren Plateau phenomenon.
-These improvements highlight the benefit of our hybrid gradient-estimation protocol in bringing larger and more complex quantum models closer to us eventually the way towards quantum utility.
+We rigorously evaluate our proposals on practical tasks like VQE and quantum phase classification, for 6 to 18 qubits.
+Our results demonstrate significant reductions in QPU calls (even up to 60% for VQE on a 16 qubit XY Hamiltonian), particularly at higher number of qubits, alongside improved relative error and success metrics for VQE and enhanced accuracy for classification.
+Notably, we observe a slower gradient decay with increasing qubit counts, suggesting a potential strategy to mitigate the barren plateau phenomenon.
+These improvements highlight the benefit of our hybrid gradient-estimation method in bringing larger and more complex quantum models toward quantum utility.
 We now discuss several open questions for further research that can improve our proposed methods:
-  • In the numerical examples, we have restricted ourselves to only Pauli gates as the BCH formula takes a convenient expression in this case.
+  • In the numerical examples, we have restricted to only Pauli gates as the BCH formula takes a convenient expression in this case.
 Since in general, this step requires calculating an infinite series, efficient protocols for calculating or approximating it will allow for more interesting gate choices in $\mathfrak{g}$-sim.
@@
 Our proposed method is aimed at reducing the necessity of using PSR to only a limited number of parameters, but leaves open the possibility of improving PSR itself.
-  • True quantum utility lies at much higher qubit counts and in using actual quantum hardware.
+  • Practical quantum utility potentially lies at much higher qubit numbers than simulated in this article and would require a demonstration on actual quantum hardware.
 We believe our methods will continue to show improvements beyond 20 qubits, and have more noise-robustness owing to fewer quantum calls required.
@@
 [102].
-  • Another direction would be to move away from the standard PQC structure to include non-unitary or general Completely Positive Trace Preserving (CPTP) maps such as in [116].
-The authors thank quantum team at Fujitsu Research India Pvt.Ltd., especially Naipunnya Raj, Ruchira Bhat and Rajiv Sangle for their discussions and inputs.
+  • Another direction would be to move away from the standard PQC structure to include non-unitary or general Completely Positive Trace Preserving (CPTP) maps such as in Ref. [103].
+The authors thank quantum team at Fujitsu Research India, especially Naipunnya Raj, Ruchira Bhat and Rajiv Sangle for their discussions and inputs.
+The paper benefited greatly from discussions with Hannes Leipold and Bibhas Adhikari from Fujitsu Research America.
 The authors extend their immense gratitude to Yasuhiro Endo, Hirotaka Oshima and Shintaro Sato as well as the entire Robust Quantum Computing Department at Fujitsu Limited for their strategic and technical support.
@@
 The code used in this paper will be made available at a later date.
-A preliminary version of this paper has been presented at Fujitsu-IISc Quantum Workshop at Bangalore, India.
-Appendix A Numerical Experiments on XY Hamiltonian
-Here we tabulate the results of the numerical experiments comparing $\mathfrak{g}$-sim , standard PSR, Alternate and Alternate+Simultaneous method of training on the XY Hamiltonian for various qubit counts in Table 2 and 3, as described in Sec.
+A preliminary version of this paper has been presented at Fujitsu-IISc Quantum Workshop, Jan 23-24 (2025) at Bangalore, India and the Fujitsu Quantum Day (Mar 28, 2025) event at Kawasaki, Japan.
+The results will also be presented as a poster in TQC’25.
+Appendix A Parameter-shift Rule
+Parameter-shift Rule (PSR) is commonly used for accurately estimating gradients of a PQC.
+For a loss function of the form given in Eq.(2), where each unitary is defined as $U_{i}(\theta_{i})=e^{-i\theta_{i}G_{i}}$ for some hermitian operator $G_{i}$, the partial derivatives are given by
+$\frac{\partial{\ell}}{\partial{\theta_{k}}}$ $=i\bra{0}{U}^{\dagger}_{\geq k}G_{k}({U}^{\dagger}_{<k}{\hat{O}}U_{<k})U_{\geq k}\ket{0}$
+$-i\bra{0}U^{\dagger}_{\geq k}(U^{\dagger}_{<k}\hat{O}{U}_{<k})G_{k}{U}_{\geq k}\ket{0}$ (18)
+where we defined $U_{<k}=U_{1}W_{1}U_{2}W_{2}\ldots U_{k-1}W_{k-1}$ and $U_{\geq k}=U_{k}W_{k}U_{k+1}W_{k+1}\ldots U_{n}W_{n}$.
+PSR can be employed to obtain a closed-form expression of Eq.(A) when the eigenspectrum of the generator $G_{k}$ either has only two unique eigenvalues or is symmetric and evenly spaced [65, 28, 104, 68, 30].
+Considering the case where $G_{k}$ with only two unique eigenvalues ±$\pm r$ i.e., $G_{k}^{2}=r^{2}I$ the corresponding unitary can be rewritten as
+$U_{k}(\theta)=\cos{r\theta}I-ir^{-1}G_{k}\sin{r\theta}$ ) (19)
+Following [28], through some simple algebra the above equation can be turned into
+$\frac{\partial{\ell}}{\partial{\theta_{k}}}$ $=\frac{r}{2}\Big{(}\langle\psi|(I-ir^{-1}G)^{\dagger}Q(I-ir^{-1}G)|\psi\rangle$
+$\langle\psi|(I+ir^{-1}G)^{\dagger}Q(I+ir^{-1}G)|\psi\rangle\Big{)}$ (20)
+where we have defined $|\psi\rangle=U_{\geq k}|0\rangle$ and $Q={U}^{\dagger}_{<k}{\hat{O}}U_{<k}$.
+Finally using Eq.(19) and substituting $\theta=\pi/4r$, the partial derivative simplifies to a difference between the loss functions evaluated on either side of the parameter with a specific eigenvalue-related shift
+$\frac{\partial{\ell}}{\partial{\theta_{k}}}=r\Big{(}\ell(\boldsymbol{\theta}+\frac{\pi}{4r}\hat{e}_{k})-\ell(\boldsymbol{\theta}-\frac{\pi}{4r}\hat{e}_{k})\Big{)}.$ (21)
+It is important to note that, despite its similarity to the finite difference approximation, the above formula is infact an exact estimation of the partial derivative upto shot noise.
+Appendix B $\mathfrak{g}$-sim Method
+In order to understand, how $\mathfrak{g}$-sim method works, we need to know how the elements of the DLA $\mathfrak{g}$ transform under commutation with elements of the DLA $\mathfrak{g}$, and under conjugation with elements of the dynamical Lie group $\mathcal{G}$.
+As $\mathfrak{g}$ is closed under commutation, the commutator between any two elements in $\mathfrak{g}$ is an element of $\mathfrak{g}$ and can be decomposed as a linear combination of a Schmidt-orthonormal basis $\{iG_{\alpha}:\alpha\in\{1,2,\ldots,\text{dim}(g)\}\}$ of $\mathfrak{g}$.
+In particular,
+[$\commutator{iG_{\alpha}}{iG_{\beta}}=\sum_{\gamma=1}^{\text{dim}(\mathfrak{g})}f_{\alpha\beta}^{\gamma}iG_{\gamma}$ (22)
+where $f_{\alpha\beta}^{\gamma}=\text{Tr}[iG_{\gamma}[iG_{\alpha},iG_{\beta}]]\in\mathbb{R}$ is computed via the standard projection in vector space $\mathfrak{g}$.
+The coefficients $f_{\alpha\beta}^{\gamma}$ are called the structure constants of the DLA $\mathfrak{g}.$
+Input: $T_{max},\eta,H=\sum_{i}^{g_{max}}a_{i}G_{i},\linebreak U(\vec{\theta})=e^{-iG_{1}\theta_{1}}\ldots e^{-iG_{n}\theta_{n}}$
+Output: Optimized $\vec{\theta^{*}}$
+Initialize: $\vec{\theta}\sim Norm(0,1)$
+Define: $C(\vec{\theta})=\langle 0^{\otimes q}|U^{\dagger}(\vec{\theta})HU(\vec{\theta})|0^{\otimes q}\rangle$
+Set: $H^{n}=H;a^{n}_{k}=a_{k},H^{n}_{k}=a_{k}^{n}G_{k}$
+begin
+       $i=0$
+       while $i\leq T_{max}$ do
+             $m=n$
+             while $m\geq 1$ do
+                   $k=0$
+                   for $k\leq g_{max}$ do
+1                         $H_{k}^{m-1}=e^{iG_{m}\theta_{m}}(a_{k}^{m}G_{k})e^{-iG_{m}\theta_{m}}=\sum_{j}b^{m-1}_{j}G_{j}$
+                         $k\leftarrow k+1$
+2
+3                  $H^{m-1}=\sum_{i}^{g_{max}}H_{i}^{m-1}=\sum_{i}^{g_{max}}a_{i}^{m-1}G_{i}$
+4
+                  $m\leftarrow m-1$
+            $l=0$
+             while $l\leq g_{max}$ do
+5                   $g_{l}=\langle 0^{\otimes q}|G_{l}|0^{\otimes q}\rangle$
+                   $l\leftarrow l+1$
+6            $C\leftarrow\sum_{i}^{g_{max}}a_{i}^{0}g_{i}$
+7             Obtain gradients $\gradient_{\vec{\theta}}C$ classically using an automatic differentiation framework
+8             $\vec{\theta}\leftarrow\vec{\theta}-\eta\gradient_{\vec{\theta}}C$
+9
+10            $i\leftarrow i+1$
+11
+12
+13
+Algorithm 3 $\mathfrak{g}$-sim Implementation for VQE
+Using the above DLA structure we demonstrate an implementation of the full $\mathfrak{g}$-sim sub-routine in action for VQE in Alg.3.
+We choose a Hamiltonian with a DLA of polynomial dimension.
+The ansatz is chosen as $U(\boldsymbol{\theta})=e^{-i\theta_{1}G_{1}}\ldots e^{-i\theta_{n}G_{n}}$, where the generators are from the same DLA, allowing us to efficiently use the DLA structure in our computation.
+The crucial factor facilitating the use of DLA structure is the Baker-Campbell-Hausdorff (BCH) formula [105] in Step. 3, which is given by
+[…]
+For operators $G_{i},G_{j}$ belonging to a chosen DLA, the expression for $e^{i\theta G_{i}}G_{j}e^{-iG_{i}}$ consists of only nested commutators.
+Since the DLA is closed under commutation operation, all of the terms in the infinite series in Eq.(23) belong to the DLA itself.
+By choosing a DLA scaling polynomially in the number of qubits, we ensure the expansion only has polynomially many unique operators, whose co-efficients are functions of the structure constants $f_{\alpha\beta}^{\gamma}$ as shown in Eq.(22).
+However, it might still be difficult to obtain a closed form expression for Eq.(23).
+Hence, we further simplify it to the scenario when all the operators are n-qubit Pauli strings.
+This gives a very simple expression
+[…]
+where $P_{i},P_{j}$ are n-qubit Pauli strings.
+We further make our numerics computationally efficient by leveraging the $2n\times 2n$ binary symplectic matrix representation of Pauli, rather than their $2^{n}\times 2^{n}$ unitary matrix representation (Section 2.1 of [106]).
+As a result, we can replace matrix multiplication by bit-wise addition, which is much faster.
+Fast application of the BCH formula due to choosing Pauli strings as well a DLA of polynomial dimension allows to calculate the cost function efficiently with polynomial compute resources.
+The gradients can then be efficiently calculated using an automatic-differentiation framework such as Pytorch [66], Tensorflow [67].
+Appendix C Overview of Barren Plateaus (BPs) and proposed mitigation techniques
+BPs are characterized by regions in the parameter space where the gradient of the loss function becomes exponentially small, severely impeding the efficient optimization of quantum circuits [95, 96, 94].
+The presence of a BP landscape implies that, for most parameter configurations, an exponentially large number of measurement samples is required to reliably estimate the loss gradient, which is essential for effective optimization.
+VQA is believed to exhibit a BP when variations in the model parameters $\theta$ lead to only exponentially small changes in the loss function $l_{\theta}(\rho,\hat{O})$, or magnitude of gradients $\partial l_{\theta}(\rho,\hat{O})/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\theta$[31, 96, 94].
+Cerezo et al., in their investigation of cost function locality, established that the inclusion of global observables $(\hat{O})$ in the cost function results in the emergence of exponentially vanishing gradients, irrespective of the circuit’s shallow depth [107].
+Additional factors contributing to this phenomenon include the degree of entanglement [108, 109, 110] and the presence of noise [111].
+A unified theoretical framework, employing group theoretical principles, covering above factors were presented in Refs.
+[96, 94].
+Beyond theoretical insights, substantial efforts have been devoted to developing practical strategies for mitigating BPs [112].
+These include advanced initialization techniques, such as layer-wise training [113], Bayesian-inspired parameter settings [114], meta-learning for initializing parameters [115], initializing circuits with sequences of identity blocks [116], and transfer-learning-based initialization approaches [117].
+Alternative optimization frameworks have also been introduced, including adaptive learning rate strategies [118] and novel line search methods designed to facilitate navigation through flat loss landscapes [119].
+Mhiri et.al.
+provided an unified analysis on warm-starting approaches in BP-affected loss landscape [120].
+Classical neural network-based methodologies have also been explored, such as generating parameters for PQCs using neural networks [121, 122].
+Empirical studies, such as those conducted in Ref.
+[123], have validated the effectiveness of these approaches in reducing the prevalence of BPs in practical applications, including QNNs.
+Enrique et al.
+have shown that the optimization landscapes of certain quantum circuit architectures inspired by classical tensor network structures such as, tree tensor networks (qTTN) and the multiscale entanglement renormalization ansatz (qMERA) are free of BPs [110].
+Furthermore, the HVA and parameterized matchgate circuits have also been shown to effectively mitigate or entirely circumvent BPs [101, 124].
+Recent research has shed light on the classical simulability of BP-free models, offering valuable insights into their underlying structure.
+Cerezo et al.
+suggested that BP-free models might be classically simulable when classical data is gathered during an initial data acquisition phase from the quantum device [35].
+This hypothesis stems from the observation that BPs arise from the curse of dimensionality, with mitigation strategies effectively reducing the problem to subspaces that are classically tractable.
+Appendix D Numerical Experiments on XY Hamiltonian
+The success, relative error and QPU calls reduction have been plotted for the HELIA configuration with 1 layer of YZ linear ansatz in Block Q in Fig.
+5.
+For the configurations with 3,6 and 9 YZ linear layers we plot the corresponding metrics in Fig.10.
+(a) Success, Relative Error and QPU Call Reduction for 3 YZ linear layers in Block Q
+(a) Success, Relative Error and QPU Call Reduction for 6 YZ linear layers in Block Q
+(a) Success, Relative Error and QPU Call Reduction for 9 YZ linear layers in Block Q
+Figure 10: XY Hamiltonian VQE (3,6 and 9 YZ linear layer): The success (Row 1), relative error (Row 2) and QPU calls reduction (Row 3) are plotted for configurations with 3 (a), 6 (b) and 9(c) layer of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G for 6 to 18 qubits.
+For each qubit count and metric, PSR (orange), Alternate (blue) and Alternate+Simultaneous(purple) training methods are compared.
+On the first row, we show that Alternate and Alternate+Simultaneous protocol generally has better success rates than PSR.
+For relative error (Row 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black).
+The Alternate method struggles in these examples, but Alternate+Simultaneous gives lower relative error than PSR for all qubits.
+Most importantly, on the lowermost row QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction).
+The spread in standard deviation is shown in black.
+Except for 6 qubit case in 9 layer configuration, we consistently see reduction in QPU calls using our methods, with lower relative error values and higher success metrics as compared to PSR.
+We also tabulate the results of the numerical experiments comparing $\mathfrak{g}$-sim , standard PSR, Alternate and Alternate+Simultaneous method of training on the XY Hamiltonian for various qubit counts in Table 3 and 4, as described in Sec.
 IV.1.1 and visualized in Fig.
-6 and 7.
+5 and 10.
 We run VQE on XY Hamiltonian for qubits ranging from 6 to 18, for the configurations with 1, 3, 6 and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G.
-Overall, our method consistently shows reduction in error of estimating the ground state (upto an order of magnitude reduction), improved success rates of reaching correct solution (upto 39% increase) and reduction in QPU calls (upto $\sim 60\%$ reduction), hence reducing the overall quantum resources compared to standard PSR training.
+Overall, our method consistently shows reduction in relative error of estimating the ground state (upto an order of magnitude reduction), improved success rates of reaching correct solution (upto 39% increase) and reduction in QPU calls (upto $\sim 60\%$ reduction), hence reducing the overall quantum resources compared to standard PSR training.
 Qubits Success(%) Full-PSR Alternate Alt+Sim 6 81.25 70.31 90.62 8 70.31 59.38 75.00 10 29.69 48.44 50.00 12 65.62 70.31 90.62 14 51.56 51.56 51.56 16 21.88 43.75 48.44 18 35.94 50.00 54.69 (a) Success rate for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits Success(%) Full-PSR Alternate Alt+Sim 6 75.00 81.25 75.00 8 68.75 65.62 68.75 10 37.50 40.62 48.44 12 59.38 70.31 73.44 14 40.62 40.62 43.75 16 29.69 34.38 34.38 18 25.00 34.38 42.19 (b) Success rate for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits Success(%) Full-PSR Alternate Alt+Sim 6 82.81 79.69 81.25 8 79.69 81.25 79.69 10 43.75 45.31 54.69 12 64.06 75.00 81.25 14 40.62 35.94 46.88 16 23.44 32.81 37.50 18 28.12 37.50 37.50 (c) Success rate for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits Success(%) Full-PSR Alternate Alt+Sim 6 82.81 81.25 98.44 8 54.69 57.81 60.94 10 43.75 43.75 54.69 12 71.88 85.94 90.62 14 57.81 48.44 53.12 16 42.19 43.75 32.81 18 28.12 29.69 54.69 (d) Success rate for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 4.27 $\pm$ 2.98 -33.27 $\pm$ 17.07 8 16.98 $\pm$ 10.62 27.53 $\pm$ 13.20 10 18.73 $\pm$ 10.87 41.88 $\pm$ 21.81 12 24.19 $\pm$ 12.65 45.25 $\pm$ 21.95 14 23.66 $\pm$ 13.55 32.07 $\pm$ 15.24 16 34.59 $\pm$ 16.09 60.00 $\pm$ 25.11 18 25.95 $\pm$ 12.14 46.81 $\pm$ 26.52 (e) QPU Reduction for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 7.71 $\pm$ 3.98 22.62 $\pm$ 12.21 8 20.36 $\pm$ 7.64 36.58 $\pm$ 15.15 10 22.57 $\pm$ 5.90 49.32 $\pm$ 18.98 12 25.48 $\pm$ 8.67 43.24 $\pm$ 16.37 14 39.11 $\pm$ 14.60 47.99 $\pm$ 20.32 16 39.50 $\pm$ 12.94 56.12 $\pm$ 21.31 18 39.02 $\pm$ 13.58 52.63 $\pm$ 22.83 (f) QPU Reduction for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 19.51 $\pm$ 6.45 7.93 $\pm$ 3.25 8 16.61 $\pm$ 4.52 22.87 $\pm$ 8.77 10 25.22 $\pm$ 7.24 32.82 $\pm$ 14.19 12 31.20 $\pm$ 10.25 22.63 $\pm$ 8.37 14 31.66 $\pm$ 11.57 46.14 $\pm$ 17.34 16 46.86 $\pm$ 20.55 50.88 $\pm$ 19.61 18 40.39 $\pm$ 16.54 45.45 $\pm$ 20.12 (g) QPU Reduction for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G Qubits QPU Reduction (%) Alternate Alt+Sim 6 -4.23 $\pm$ 1.25 -4.70 $\pm$ 2.13 8 14.00 $\pm$ 3.55 17.64 $\pm$ 7.73 10 18.27 $\pm$ 4.36 22.70 $\pm$ 10.14 12 27.63 $\pm$ 8.91 13.74 $\pm$ 3.93 14 25.34 $\pm$ 8.07 38.17 $\pm$ 14.61 16 30.14 $\pm$ 10.76 44.47 $\pm$ 19.66 18 27.78 $\pm$ 9.02 39.07 $\pm$ 15.71 (h) QPU Reduction for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
@@
 The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR.
-The metrics show that our methods are generally better at reaching higher success rates and lower error(in Table 3), while the QPU calls are reduced significantly
-Qubits Error (1e-5)
+The metrics show that our methods are generally better at reaching higher success rates and lower relative error(in Table 4), while the QPU calls are reduced significantly
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
@@
 18 118.38 117.59 119.3 0.0758 0.0758 0.1137 0.0716 0.0590 0.0842 0.0674 0.0632 0.0758
-(a) Error for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
-Qubits Error (1e-5)
+(a) Relative Error for 1 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
@@
 18 118.38 117.59 119.3 0.1390 0.0821 0.3032 3.7105 0.7981 9.5521 0.0674 0.0674 0.0758
-(b) Error for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
-Qubits Error (1e-5)
+(b) Relative Error for 3 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
@@
 18 118.38 117.59 119.3 0.1137 0.0927 0.1853 4.9487 2.4343 15.1262 0.0842 0.0674 0.1032
-(c) Error for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
-Qubits Error (1e-5)
+(c) Relative Error for 6 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt+Sim
@@
 18 118.38 117.59 119.3 0.1390 0.0863 0.2843 6.6966 4.5486 13.9533 0.1011 0.0842 0.1137
-(d) Error for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
-Table 3: XY Hamiltonian VQE (Error) : The error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G.
-The error for $\mathfrak{g}$-sim with only XY Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods.
-Alternate and Alternate + Simultaneous method shows generally lower errors compared to the Full-PSR training
-Appendix B Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
+(d) Relative Error for 9 layer YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G
+Table 4: XY Hamiltonian VQE (Relative Error) : The relative error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and XY Hamiltonian DLA gates in Block G.
+The relative error for $\mathfrak{g}$-sim with only XY Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods.
+Alternate and Alternate + Simultaneous method shows generally lower relative error compared to the Full-PSR training
+Appendix E Barren Plateau Mitigation for VQE on XY Hamiltonian
+Here we plot the variance of the gradients for configurations with 3,6 and 9 YZ linear layers in Block Q of HELIA, when used for VQE on XY Hamiltonian in Fig.
+11.
+In each case, we track the gradient of the first parameter from each of the Block Q (referred as $U_{q}$) and Block G (referred as $U_{g}$) and plots its variance during the training (Fig.11).
+Each datapoint is obtained after evaluating the variance from the full training of 64 independent trials.
+(a) 3 YZ linear layers in Block Q
+(b) 6 YZ linear layers in Block Q
+(c) 9 YZ linear layers in Block Q
+Figure 11: Gradients for XY Hamiltonian VQE: Variance of the first parameter of Block Q ($U_{q}$, red) and G ($U_{g}$, blue) are plotted with increasing qubits for both Full-PSR and our Alternate method for finding ground state of XY Hamiltonian using VQE.
+The diagrams show the plots for 3(a), 6(b) and 9(c) YZ linear layer in Block Q.
+For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green.
+In both Block Q and G, we notice a slower decay of gradients as compared to HEA50.
+Overall the plot shows that our chosen ansatz is able to preserve higher magnitude of gradients even at high qubits in both of the blocks.
+For comparison we plot a deep hardware efficient ansatz (labelled HEA50) in green in Fig.11.
+HELIA shows a clear improvement in terms of slow decay in gradients for all the configurations of Block Q.
+The gradients for Block Q and G demonstrate a much slower decay compared to the HEA50 circuit.
+Overall, this allows for larger qubit models to be trained efficiently without running into vanishingly small gradients.
+Appendix F $\mathfrak{g}$-Purity of Hardware Efficient Ansatz at various circuit depths
+Using results from Ref. [96] and considering $\mathfrak{g}$ to be simple we obtain Eq.(11). The $\mathfrak{g}$-Purity is defined as,
+[…]
+[…]
+Based on Eq.
+(11), we see that in order to have sub-exponential decay of gradients, $\mathcal{P}_{\mathfrak{g}}(\rho_{q}(\theta))$ or $g$-Purity plays an important role.
+In Fig.
+12, we numerically evaluate it for varying circuit depths considering the $\mathfrak{g}_{XY}$ DLA.
+Both constant and logarithmic depth circuits show sub-exponential decay while linear depth circuit has exponential decay.
+This indicates that HELIA with logarithmic depth HEA should also be able to avoid BP.
+However, we keep that direction open for future exploration.
+Figure 12: Average $\mathfrak{g}$-Purity of HEA for $\mathfrak{g}_{XY}$ DLA (defined in Eq.
+(25)) with increasing qubits and varying circuit depths.
+For constant ($O(1)$) and logarithmic depth in qubits ($O(log(n))$), the decay is less than exponential (shown in blue and purple respectively).
+For linear depth ($O(n)$ in green), $\mathfrak{g}$-Purity decreases exponentially.
+Each point in the plot corresponds to an average of 10000 independent parameter samples.
+Appendix G Example instances of $\mathfrak{g}$-sim reaching accurate solution without quantum hardware
 Although our aim has been to elevate $\mathfrak{g}$-sim to a hybrid scheme with a more expressive circuit, often it might not be necessary for a given task.
@@
 We use $\mathfrak{g}$-sim for finding the ground state of the XY Hamiltonian (Eq.
-14) for 13 and 17 qubit cases, and the results are shown in Fig.11.
-Clearly, $\mathfrak{g}$-sim is able to converge to good solutions (Error $\sim 10^{-3}-10^{-5}$) without adding any additional gates.
+8) for 13 and 17 qubit cases, and the results are shown in Fig.13.
+Clearly, $\mathfrak{g}$-sim is able to converge to good solutions (Relative Error $\sim 10^{-3}-10^{-5}$) without adding any additional gates.
 (a) XY Hamiltonian with 13 qubits
 (b) XY Hamiltonian with 17 qubits
-Figure 11: Error vs Iteration curve for finding ground state of an XY Hamiltonian using only $\mathfrak{g}$-sim .
-For 13 (11(a)) and 17 (11(b)) qubits, $\mathfrak{g}$-sim is enough to obtain the ground state and does not require our protocol.
-Appendix C VQE for TFIM Hamiltonian
+Figure 13: Relative Error vs Iteration curve for finding ground state of an XY Hamiltonian using only $\mathfrak{g}$-sim .
+For 13 (13(a)) and 17 (13(b)) qubits, $\mathfrak{g}$-sim is enough to obtain the ground state and does not require our protocol.
+Appendix H VQE for TFIM Hamiltonian
 We also implement our training schemes for VQE using a different Hamiltonian example, in order to demonstrate that the improvements are not restricted to a specific choice.
 For this, we use the Transverse Field Ising Models (TFIM) Hamiltonian,
-$H_{TFIM}=\sum_{i}\alpha_{i}X_{i}X_{i+1}+\sum_{j}\beta_{j}Z_{j}$ (23)
-which has a poly-DLA scaling as $2n^{2}-n$ for $n$ qubits. We measure the improvement in terms of Error, Success and QPU calls reduction as defined in IV.
-We tabulate the numerical results comparing $\mathfrak{g}$-sim , Full-PSR, Alternate and Alternate+Simultaneous in Table 4 and 5 and provide plots for visualization in Fig.12.
-Except the 12 qubit case, where $\mathfrak{g}$-sim is able to produce accurate solutions (similar to the scenario mentioned in App.B), our proposed Alternate+Simultaneous method generally achieves a lower error and higher success rate compared to the other protocols, while almost always reducing the QPU calls necessary compared to Full-PSR.
-Figure 12: TFIM Hamiltonian VQE: The success (Column 1), error (Column 2) and QPU calls reduction (Column 3) are plotted for configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and Hamiltonian DLA gates in Block G, for various qubit counts in each row.
-For each configuration and metric, PSR (orange), Alternate(blue) and Alternate+Simultaneous(purple) training methods are compared.
-For error (Column 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black).
-For QPU calls reduction only Alternate and Alternate+Simultaneous is plotted since we compare it relative to PSR (which will coincide with 0% QPU calls reduction).
+$H_{TFIM}=\sum_{i=0}^{N}\alpha_{i}X_{i}X_{i+1}+\sum_{j}\beta_{j}Z_{j}$ (26)
+which has a poly-DLA scaling as $2n^{2}-n$ for $n$ qubits. We measure the improvement in terms of Relative Error, Success and QPU calls reduction as defined in IV.
+We tabulate the numerical results comparing $\mathfrak{g}$-sim , Full-PSR, Alternate and Alternate+Simultaneous in Table 5 and 6 and provide plots for visualization in Fig.
+17.
+Except the 12 qubit case, where $\mathfrak{g}$-sim is able to produce accurate solutions (similar to the scenario mentioned in App.G), our proposed Alternate+Simultaneous method generally achieves a lower relative error and higher success rate compared to the other protocols, while almost always reducing the QPU calls necessary compared to Full-PSR.
+(a) Success, Relative Error and QPU Call Reduction for 1 YZ linear layers in Block Q
+(a) Success, Relative Error and QPU Call Reduction for 3 YZ linear layers in Block Q
+(a) Success, Relative Error and QPU Call Reduction for 6 YZ linear layers in Block Q
+(a) Success, Relative Error and QPU Call Reduction for 9 YZ linear layers in Block Q
+Figure 17: TFIM Hamiltonian VQE (1,3,6 and 9 YZ linear layer): The success (Row 1), relative error (Row 2) and QPU calls reduction (Row 3) are plotted for configurations with 1 (a), 3 (b) and 6 (c) and 9 (d) layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G for 8 to 14 qubits.
+For each qubit count and metric, PSR (orange), Alternate (blue) and Alternate+Simultaneous(purple) training methods are compared.
+On the first row, we compare the success rates for both Alternate and Alternate+Simultaneous protocol with PSR.
+For relative error (Row 2), we plot the median and the $25\%$ and $75\%$ quantiles (in black).
+The Alt+Sim method gives comparable or lower relative error than PSR for all qubits.
+Most importantly, on the lowermost row QPU calls reduction is plotted for only Alternate and Alternate+Simultaneous since we compare it relative to PSR (which will coincide with 0% QPU calls reduction).
 The spread in standard deviation is shown in black.
-We consistently see reduction in QPU calls using our methods, with frequently lower error values and higher success metrics as compared to PSR
-Figure 13: Gradients for TFIM Hamiltonian VQE: Variance of the first parameter from PSR and $\mathfrak{g}$-sim block are plotted with increasing qubits for both Full-PSR (blue) and our Alternate method (orange) for finding ground state of TFIM Hamiltonian using VQE.
-The diagrams also show the plots for varying number of PSR block layers (1,3,6,9 from top to bottom).
+Except for 14 qubit case in 9 layer configuration, we consistently see reduction in QPU calls using our methods, with lower relative error values and higher success metrics as compared to PSR.
+(a) 1 YZ linear layer in Block Q
+(b) 3 YZ linear layer in Block Q
+(c) 6 YZ linear layer in Block Q
+(d) 9 YZ linear layer in Block Q
+Figure 18: Gradients for TFIM Hamiltonian VQE: Variance of the first parameter from Block Q ($U_{q}$,red) and G ($U_{g}$,blue) are plotted with increasing qubits for both Full-PSR and our Alternate method for finding ground state of TFIM Hamiltonian using VQE.
+The diagrams also show the plots for 1 (a), 3 (b), 6(c), 9 (d) YZ linear layers in Block Q .
 For comparison, the same task is performed using a deep HEA (referred as HEA50) with standard PSR and variance of its first partial derivative is plotted in green.
@@
 The QPU call reduction is measured relative to the Full-PSR method where all parameters are trained by standard PSR.
-The metrics show that our methods are generally better at reaching higher success rates and lower error(in Table 5), while the QPU calls are reduced significantly
-Qubits Error (1e-5)
+The metrics show that our methods are generally better at reaching higher success rates and lower relative error(in Table 6), while the QPU calls are reduced significantly
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
@@
 14 24.8 24.8 24.8 12.3975 12.3975 12.4076 12.3975 12.3975 12.4304 12.4178 12.4178 12.4178
-(a) Error for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
-Qubits Error (1e-5)
+(a) Relative Error for 1 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
@@
 14 24.8 24.8 24.8 12.4178 12.3874 12.5901 12.4989 12.3975 12.8790 12.4279 12.3975 12.4381
-(b) Error for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
-Qubits Error (1e-5)
+(b) Relative Error for 3 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
@@
 14 24.8 24.8 24.8 12.4279 11.3585 12.5420 13.3352 11.9059 15.0483 12.4381 11.2799 12.4786
-(c) Error for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
-Qubits Error (1e-5)
+(c) Relative Error for 6 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
+Qubits Relative Error (1e-5)
 $\mathfrak{g}$-sim Full-PSR Alternate Alt + Sim
@@
 14 24.8 24.8 24.8 12.4685 7.7801 12.6256 17.4254 13.2718 23.5684 12.4381 9.5845 12.4887
-(d) Error for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
-Table 5: TFIM Hamiltonian (Error): The error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G.
-The error for $\mathfrak{g}$-sim with only TFIM Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods.
-Alternate and Alternate + Simultaneous method shows generally lower errors compared to the Full-PSR training
-We also provide plots of variance of gradients for the first parameter from each of $\mathfrak{g}$-sim and PSR block in Fig.
-13.
+(d) Relative Error for 9 layer YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G
+Table 6: TFIM Hamiltonian (Relative Error): The relative error median, 25% quantile (Q25) and 75% quantile (Q75) for Full-PSR, Alternate and Alternate + Simultaneous (referred as Alt+Sim) for the configurations with 1, 3, 6, and 9 layers of YZ linear ansatz in Block Q and TFIM Hamiltonian DLA gates in Block G.
+The relative error for $\mathfrak{g}$-sim with only TFIM Hamiltonian DLA gates for each qubit count is shown for comparison, which is several orders of magnitude higher than the other methods.
+Alternate and Alternate + Simultaneous method shows generally lower error compared to the Full-PSR training
+We also provide plots of variance of gradients for the first parameter from each of Block Q and G in Fig.
+18.
 Consistent with our previous example, the gradients decay slower using our choice of ansatz as compared to a deep Hardware Efficient Ansatz with 50 layers.
+Input: $T_{alt},T_{sim},T_{psr},T_{full},U_{q}(\mathbf{\theta_{0}}),H=\sum_{i}\beta_{i}P_{i}$
+Output: $\theta^{*},\phi^{*}$
+Initialize: $\vec{\theta}_{0}\sim\mathcal{N}(0,1);i=0$
+begin
+       poly-DLA Phase
+1       Obtain a subset of operators $\{P_{j}\}_{j\in\mathcal{A}}$ from $H$ such that it forms a poly-DLA $\mathfrak{g}$.
+2       Define $H_{poly}=\sum_{P_{i}\in\mathcal{A}}\beta_{i}P_{i}$, and unitary $U_{g}(\vec{\phi_{0}})=e^{-iP_{1}\phi_{1}^{0}}\ldots e^{-iP_{n}\phi_{n}^{0}}$
+3       Randomly initialize $\vec{\phi}_{0}\sim\mathcal{N}(0,1)$
+4       Define the cost function: $C_{poly}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})H_{poly}U_{\mathfrak{g}}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$
+5       Run Alternate optimization given in Alg.1 for $T_{alt}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
+6       Run Simultaneous optimization given in Alg.2 for $T_{sim}$ iterations on cost function $C_{poly}(\vec{\theta},\vec{\phi})$
+7       Extract the optimized parameters $(\vec{\theta}^{*},\vec{\phi}^{*})$
+       Exponential-DLA Phase
+8       Define the full cost function, $C_{full}(\vec{\theta},\vec{\phi})=\langle 0^{\otimes n}|U_{q}^{\dagger}(\vec{\theta})U_{g}^{\dagger}(\vec{\phi})HU_{\mathfrak{g}}(\vec{\phi})U_{q}(\vec{\theta})|0^{\otimes n}\rangle$ initialized at $(\vec{\theta}^{*},\vec{\phi}^{*})$
+9       Optimize only parameters $\vec{\theta}$ of $U_{q}(\vec{\theta})$ in $C_{full}(\vec{\theta},\vec{\phi})$ for $T_{psr}$ iterations using PSR
+10       Optimize full cost function $C_{full}(\vec{\theta},\vec{\phi})$ for $T_{full}$ using PSR
+11       Extract optimized parameters at the end of training $(\vec{\theta}^{*},\vec{\phi}^{*})$
+12
+Algorithm 4 VQE for General Hamiltonian
+Appendix I VQE for LTFIM Hamiltonian (exponential-DLA)
+In this section, we provide the algorithm (Algo.
+4) and plots for VQE on 10 and 12 qubit LTFIM Hamiltonian using PSR and our proposed method in Fig.
+19.
+The relative error metric in Eq.(7) is shown for 3 and 6 YZ linear layers in Block Q.
+Although the initial training phase is based on the reduced Hamiltonian, the plotted energy values correspond to expectation values of the full Hamiltonian.
+Experiment Details: $U_{q}$ is composed of 3 layers of YZ linear ansatz, while $U_{\mathfrak{g}}$ is generated out of gates from $\mathfrak{g}_{TFIM}$.
+In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of $U_{q}$ and 1000 iterations of full circuit training.
+Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods.
+However, these choices are made ad-hoc and can be modified based on other stopping conditions, such as early stopping to further reduce QPU calls.
+(a) Trained using standard PSR
+(b) Pre-training using our method followed by PSR
+(c) Trained using standard PSR
+(d) Pre-training using our method followed by PSR
+(e) Trained using standard PSR
+(f) Pre-training using our method followed by PSR
+Figure 19: Demonstration of PSR (a),(c), (e) and our method (b),(d),(f) to run VQE for 10 and 12 qubit LTFIM Hamiltonian (with an exponential DLA) with 3 and 6 YZ linear layers in Block Q.
+The orange and green regions show Alternate and Simultaneous training respectively for the restricted Hamiltonian, followed by training Block Q and full circuit in blue and pink regions respectively for the full LTFIM Hamiltonian.
+We notice that our training methods in the initial phase allows VQE to reach much closer to the target ground state during the final training, as compared to just using PSR for the full training, for similar number of of total training iterations.
+For the overall training, we also reduce the QPU calls ($\sim 6-10\%$).
+As shown in Fig.19, the algorithm is indeed able to converge to good solutions by using the our protocols in the first phase of the training.
+Compared to PSR, our protocol shows lower relative error value on an average.
+The QPU calls required for the 10-qubit Hamiltonian VQE using PSR is $54.5e+$ 5 and $56.3e+$ 5 for Layer 3 and 6 respectively, while our proposed method requires $54.065e+$ 5 ($\sim 9.6\%$ reduction) and $55.925e+$ 5 ($\sim 6.5\%$ reduction) QPU calls respectively.
+Similarly, for 12 qubit, PSR uses $58.28e+$ 5 for Layer 6 while our method requires $57.698e+$ 5 ($\sim 7\%$ reduction) respectively.
+Appendix J Experimental details of VQE on LiH and Classification task
+LiH Experimental details: $U_{q}$ is composed of 6 layers of YZ linear ansatz, while $U_{\mathfrak{g}}$ is generated out of gates from the above DLA.
+In the shown examples, we run Alternate optimization for 250 iterations and Simultaneous optimization for 100 iterations with the reduced Hamiltonian, followed by 200 iterations of $U_{q}$ and 1000 iterations of full circuit training.
+Similarly for Full-PSR training, we run 1450 iterations of PSR to match the number of total iterations in both methods.
+These ad-hoc choices can potentially be further optimized to reduce resources and improve accuracy.
+Classification Experiment Details: We construct HELIA with 9 YZ linear layers in $U_{q}$.
+$U_{\mathfrak{g}}$ is constructed from one of the 3 DLA choices mentioned above, and we repeat the experiment for each of the choices.
+The measurement operator is chosen randomly from the corresponding DLA and fixed at the beginning of the experiment.
+For each of these configurations of DLA, we run 20 independent trials and compare test accuracy.